### v1

# A Unified Framework for Inference with General Missingness Patterns and Machine Learning Imputation

Abstract
Pre-trained machine learning (ML) predictions have been increasingly used to complement incomplete data to enable downstream scientific inquiries, but their naive integration risks biased inferences. Recently, multiple methods have been developed to provide valid inference with ML imputations regardless of prediction quality and to enhance efficiency relative to complete-case analyses. However, existing approaches are often limited to missing outcomes under a missing-completely-at-random (MCAR) assumption, failing to handle general missingness patterns under the more realistic missing-at-random (MAR) assumption. This paper develops a novel method which delivers valid statistical inference framework for general Z-estimation problems using ML imputations under the MAR assumption and for general missingness patterns. The core technical idea is to stratify observations by distinct missingness patterns and construct an estimator by appropriately weighting and aggregating pattern-specific information through a masking-and-imputation procedure on the complete cases. We provide theoretical guarantees of asymptotic normality of the proposed estimator and efficiency dominance over weighted complete-case analyses. Practically, the method affords simple implementations by leveraging existing weighted complete-case analysis software. Extensive simulations are carried out to validate theoretical results. The paper concludes with a brief discussion on practical implications, limitations, and potential future directions.
1 Introduction
With the rapid development of machine learning (ML) and artificial intelligence (AI) techniques, ML models, particularly those trained on large-scale datasets, have become increasingly capable of making highly accurate predictions. These predictions can be treated as imputations for missing data (in both dependent $Y$ and independent variables $\mathbf{X}$) due to nonresponse, expensive data collection costs, or other factors. Because ML-imputed datasets can be seamlessly applied in downstream statistical analyses, such appealing impute-then-analyze procedures is a double-edged sword. On the one hand, the ease of obtaining ML predictions for imputation catalyzes wide adoptions across various disciplines, including finance (Bryzgalova et al.,, 2025), social science (Egami et al.,, 2023), clinical psychology (Little,, 2024), biomedicine (Barrabés et al.,, 2024), and public health (Perkins et al.,, 2018) studies, where missing data appear commonly during the data collection process. On the other hand, imputing missing values with ML predictions of arbitrary quality may produce datasets that have different conditional and marginal distributions from the underlying true distribution. Proceeding with analysis without careful consideration may result in biased estimation and invalid inference.
To address this issue, recent research has explored prediction-based methods to incorporate ML predictions for single imputation while guaranteeing valid statistical inference on datasets with partially missing outcomes (see, for example, Wang et al., (2020); Angelopoulos et al., 2023a ; Egami et al., (2023); Salerno et al., 2025a , or Salerno et al., 2025b ). Specifically, prediction-powered inference (PPI, Angelopoulos et al., 2023a, ) is a framework where inference is conducted using ML-imputed outcomes, with model parameter estimates calibrated, or “rectified”, using a complete-case subset. Since this approach does not impose assumptions on the imputation model itself, it allows for a broad selection of ML-based imputation methods, the quality of which impacts the efficiency but not the validity of inference (e.g., consistency, achieving nominal coverage rate). Several extensions of this framework have adapted the idea of PPI to important areas including causal inference (Demirel et al.,, 2024), Bayesian inference (Hofer et al.,, 2024; Ročková and O’Hagan,, 2025), and genome-wide association studies (GWAS, Miao et al.,, 2024). Furthermore, recognizing that the PPI approach does not guarantee a smaller variance relative to complete-case analysis (CCA), a few authors have proposed weighting the rectifier with tuning parameters to achieve better efficiency than CCA (Angelopoulos et al., 2023b, ; Miao et al.,, 2023; Gan et al.,, 2024; Gronsbell et al.,, 2024; Ji et al.,, 2025).
Despite the tremendous progress and development in prediction-based inference, two key challenges remain unaddressed. First, the existing literature considers only the assumption of missing completely at random (MCAR), leaving open the question of the applicability of prediction-based inference under the more common and complicated assumption, that is, the assumption missing at random (MAR) (Rubin,, 1976; Little and Rubin,, 2019). Except for several specific cases such as linear regression analysis and odds ratio estimation under certain missing probability assumptions  (Example 3.3 and 3.4, Little and Rubin,, 2019), failure to account for the MAR assumption can result in biased estimates. Second, the current literature focuses exclusively on partially missing outcomes or specific patterns in covariates missingness (e.g., blockwise, Kluger et al.,, 2025), but does not account for general patterns of non-monotone missingness, where missing values can occur in any cell within the tabular (“rectangular”) dataset. For example, in regression analysis, missing values are common in both outcomes and covariates, and we may observe different patterns of missingness in different subjects, e.g., missing $Y$ and $X_{1}$ for one subject and missing $Y$ and $X_{2}$ for the other.
In this paper, we address these challenges by proposing a unified framework called Pattern-Stratified PPI (PS-PPI) [^1] By Corollary 3.2, the proposed procedure enjoys efficiency dominance over WCCA (under MAR, general missing patterns). This resembles the property of PPI++ being no less efficient than CCA (under MCAR, outcome missing only). We chose to term our method “PS-PPI” rather than “PS-PPI++” to avoid redundancy. to account for both general missingness patterns and missing-at-random mechanism under the Z-estimation framework. We make a few theoretical and methodological contributions with useful practical implications. First, we generalize the existing PPI-type estimators that only deal with missing outcomes under MCAR. We develop an inferential procedure that integrates the concept of missingness patterns that are widely used in semiparametetric inference to deal with non-monotone missing data (Tsiatis,, 2006). We will show the proposed estimator belongs to the class of Augmented Inverse Probability Weighted (AIPW) estimators, building on a prior work under MCAR with only outcomes subject to missingness (Gronsbell et al.,, 2024). Second, we provide theoretical guarantee that the resulting PS-PPI estimator is at least as efficient as the weighted complete-case analysis (WCCA) even when the imputation models are of low quality. Third, our proposed PS-PPI procedure is straightforward to implement. It works by breaking the analysis into several weighted complete-case (WCC) Z-estimation problems, each of which can be solved with existing statistical packages designed for WCC analysis.
The remainder of the paper is organized as follows. In Section 2, we introduce necessary notation, the concepts of missing data mechanisms and missingness patterns, and the inferential target under the Z-estimation framework; we review the core idea used in existing PPI literature that deals with missing outcome or blockwise missing covariates, along with their limitations. In Section 3, we introduce the details of the proposed PS-PPI procedure and present theoretical results on its asymptotic properties; we provide examples to show how our proposed framework subsumes previous works as special cases. In Section 4, we perform data simulation to verify the validity and superior performance of our proposed PS-PPI relative to alternatives that have limited ability to extract and integrate information from incomplete datasets with multiple different missingness patterns. We discuss the limitations of this work and potential future directions in Section 5.
2 Setup
2.1 Notation
In this subsection we introduce the notation; a comprehensive glossary of the notation is presented in the Supplementary Materials. Let $\mathcal{D}=\{\bm{Y}_{1},\ldots,\bm{Y}_{N}\}$ denote the full data for $N$ observations, subject to potential missingness, where $\bm{Y}_{i}=(Y_{i1},\ldots,Y_{ip})^{\top}\in{\mathcal{Y}}={\mathcal{Y}}_{1}\times\cdots\times{\mathcal{Y}}_{p}$ represents the $p$-dimensional data for subject $i$, ${\mathcal{Y}}_{j}\subseteq\mathbb{R}$ is the marginal sample space of the $j$-th component in $\bm{Y}_{i}$. Each $\bm{Y}_{i}$ is assumed to be independently and identically sampled from a population distribution $\mathbb{P}_{\bm{Y}}$. In practice, three types of variables may be included in the vector $\bm{Y}_{i}$: (a) a variable treated as the scalar outcome in a specific scientific inquiry; (b) q1q-1 variables treated as covariates; (c) pqp-q auxiliary variables not included in the scientific analysis but potentially useful to predict variables in type (a) and (b). In this paper, unless otherwise noted, we do not explicitly distinguish among these types in our notation to preserve the flexibility of our proposed method in large-scale analyses such as LabWAS (Goldstein et al.,, 2020), where a variable may serve as an outcome in one analysis and as a covariate in another. However, for readers who require this distinction in their scientific inquiries (e.g., via a regression analysis), one may assume that $Y_{i1}$ and $Y_{i2},\ldots,Y_{iq}$} represent the outcome and the covariates, respectively. Note that the partition of $\bm{Y}_{i}$ into three types of variables mentioned above is purely for the ease of understanding and has no impact on the theoretical results in this paper.
We will assume that the analyst may not observe the full data $\mathcal{D}$}. Let $\bm{R}_{i}=(R_{i1},\ldots,R_{ip})^{\top}\\ \in\{0,1\}^{p}$ be a binary vector of $p$ dimension, where $R_{ij}=1$ indicates that $Y_{ij}$ is observed and $R_{ij}=0$ otherwise. Consequently, we denote the observed components $\bm{Y}_{i,\text{obs}}=\{Y_{ij}|R_{ij}=1,j=1,\ldots,p\}\in\prod_{\{j:R_{ij}=1\}}{\mathcal{Y}}_{j}$, and similarly the missing components $\bm{Y}_{i,\text{mis}}=\{Y_{ij}|R_{ij}=0,j=1,\ldots,p\}\in\prod_{\{j:R_{ij}=0\}}{\mathcal{Y}}_{j}$, for subject $i=1,\ldots,N$. Let ${\mathcal{D}}_{\text{obs}}=\{\bm{Y}_{1,\text{obs}},...,\bm{Y}_{N,\text{obs}}\}$ denote the observed data.
2.1.1 Missingness patterns and mechanisms
In the existing literature on prediction-based inference, typically only the outcome variable is subject to potential missingness (e.g., Angelopoulos et al., 2023a, ; Angelopoulos et al., 2023b, ; Miao et al.,, 2023; Gan and Liang,, 2023; Gronsbell et al.,, 2024), with the exception of  Kluger et al., (2025) which focused on blockwise covariate missingness patterns. However, in practice, missingness can occur anywhere within the dataset, e.g., for both outcomes and covariates under the regression setting.
To address a more general missingness structure, we introduce the concept of missingness patterns and formalize the general missingness pattern problem by treating each missing pattern as a special case of coarsening (Heitjan and Rubin,, 1991), following Tsiatis, (2006). A missingness pattern is defined as an element in the sample space $\{0,1\}^{p}$ of the missing indicator vector $\bm{R}_{i}$. Excluding the case where all variables are missing, there are at most $2^{p}-1$ different missingness patterns appearing in the data. In the context of prediction-based inference, it is generally assumed that there exist fully observed, low-cost variables that serve as inputs for predicting the missing values (i.e., type (c) variables). Under this assumption, the number of distinct missingness patterns is reduced to at most $2^{q}$. Each of the missingness patterns can be indexed with $k\in\{1,...,K\}$; we define $k=\infty$ to indicate that the subject is fully observed. The $k$-th missingness pattern is denoted as $\bm{R}^{(k)}=(R^{(k)}_{1},\ldots,R^{(k)}_{p})\in\{0,1\}^{p}$. We call the $i$-th subject $\bm{Y}_{i}$ to be in the $k$-th stratum of missingness pattern if $\bm{R}_{i}=\bm{R}^{(k)}$. Following Tsiatis, (2006), for the $k$-th missingness pattern, we introduce a deterministic and many-to-one coarsening function $G_{k}:{\mathcal{Y}}\mapsto\prod_{\{j:R^{(k)}_{j}=1\}}{\mathcal{Y}}_{j}$, such that $G_{k}(\bm{Y}_{i})=\bm{Y}_{i,\text{obs}}$ if $\bm{Y}_{i}$ is under the $k$-th missingness pattern. For fully observed data, $G_{\infty}(\bm{Y}_{i})$ is the identity mapping. Note that the coarsening notation is borrowed only to emphasize the constructive nature of our proposed method, which combines pattern-stratified estimates (see Equation (2)); the prediction-based inference for general coarsened data where $G_{k}$ is broader than entry subset selection is beyond the scope of this work. With the notation of the coarsening functions and missingness patterns, we can represent the observed dataset as ${\mathcal{D}}_{\text{obs}}=\{(G_{{\mathcal{C}}_{i}}(\bm{Y}_{i}),{\mathcal{C}}_{i})\}^{N}_{i=1}$, where ${\mathcal{C}}_{i}\in\{1,...,K,\infty\}$ is the coarsening index for subject $i$.
Missing data mechanisms play an important role in dictating how researchers should deal with the missingness in data analyses (Rubin,, 1976). Here we provide a brief overview for completeness and refer readers to Little and Rubin, (2019) for a comprehensive, book-length treatment of this topic. Based on how the conditional distribution of the coarsening index ${\mathcal{C}}_{i}$ given $\bm{Y}_{i}$ can be simplified, a few general categories with distinct analytic implications can be defined (Tsiatis,, 2006). In the simplest case with the most restrictive assumption, the probability distribution of the coarsening index ${\mathcal{C}}_{i}$ does not depend on any observed or unobserved variables, namely, $\pi({\mathcal{C}}_{i}=k|\bm{Y}_{i};{\bm{\omega}})=\pi({\mathcal{C}}_{i};{\bm{\omega}})$ where $\bm{\omega}$} is a vector of parameters of the conditional distribution. In this scenario, we consider the data to be missing completely at random (MCAR). A more common and realistic case compared to MCAR is that the missingness pattern depends on variables that are observed: $\pi({\mathcal{C}}_{i}|\bm{Y}_{i};{\bm{\omega}})=\pi({\mathcal{C}}_{i}|\bm{Y}_{i,\text{obs}};{\bm{\omega}})$, referred to as missing at random (MAR). If the missingness depends on the missing values themselves: $\pi({\mathcal{C}}_{i}|\bm{Y}_{i};{\bm{\omega}})=\pi({\mathcal{C}}_{i}|\bm{Y}_{i,\text{obs}},\bm{Y}_{i,\text{mis}};{\bm{\omega}}),$ it is referred to as missing not at random (MNAR). In practice, the MAR assumption is more commonly considered because it does not suffer from the identifiability problems that exists in the MNAR assumption, yet it is weaker than the MCAR assumption (Tsiatis,, 2006). In this paper, we will introduce and provide theoretical analyses of our methods under the MAR assumption.
2.1.2 Prediction of missing or masked values
For each missingness pattern $k=1,\ldots,K$, suppose we have access to a prediction function $h_{k}$ trained on data external to ${\mathcal{D}}_{\text{obs}}$ where $h_{k}:\prod_{\{j:\bm{R}^{(k)}_{j}=1\}}{\mathcal{Y}}_{j}\mapsto\prod_{\{j:\bm{R}^{(k)}_{j}=0\}}{\mathcal{Y}}_{j}$. A key ingredient in our proposed method will use the following operation: in addition to filling the missing values for the subjects in the stratum of the $k$-th missingness pattern, we can also mask the entries for any fully-observed subject according to the missingness pattern $k$, i.e., $\{Y_{ij}:j\in\{j^{\prime}:R_{j^{\prime}}^{(k)}=0\}\}$ for subject $i$ with ${\mathcal{C}}_{i}=\infty$ and then use $h_{k}$ to obtain their predictions to impute these “pseudo” missing entries.
In fact, to study prediction-based inference, we will show in our theoretical analyses that it is central to generalize this notion of data masking-and-imputation by missingness pattern $k$ from the fully observed subjects (${\mathcal{C}}_{i}=\infty$) to the entire population. For each missingness pattern $k=1,\ldots,K$, let the random vector $\bm{Y}^{(k)}_{\text{imp}}=\left\{G_{k}(\bm{Y}),h_{k}\left(G_{k}(\bm{Y})\right)\right\}$ be comprised of observed components $G_{k}(\bm{Y})$ and masked-and-imputed components $h_{k}\left(G_{k}(\bm{Y})\right)$, where $\bm{Y}$ follows the population full-data distribution $\mathbb{P}_{\bm{Y}}$. We assume that $\bm{Y}^{(k)}_{\text{imp}}$ is distributed according to a population-level distribution $\mathbb{P}^{(k)}_{\bm{Y}_{\text{imp}}}$. We note two key points. First, $h_{k}$ can be deterministic or stochastic as long as there exists a fixed population distribution $\mathbb{P}^{(k)}_{\bm{Y}_{\text{imp}}}$ which we will assume in theoretical analyses (Section 3). This assumption is less tenable if the prediction mappings $\{h_{k}\}$ shift over time, e.g., due to model updates that are continually pushed into online repositories to progressively impute different subsets of data. Second, the current setup does not require distinct prediction models to be trained for each of the $K$ missingness patterns. When assuming type (c) variables exist, the predictions can be accomplished by using at most $q$ univariate base prediction models whenever a missing value is encountered. In this sense, $h_{k}$ can be viewed as invoking a subset of $\sum_{j}(1-R^{(k)}_{j})$ of the $q$ base prediction models for missingness pattern $k$. For example, the model for predicting diabetes status ($Y_{i1}$) using BMI only ($Y_{ip}$) can be used for any missingness pattern $k$ with $R^{(k)}_{1}=0$ and $R^{(k)}_{p}=1$. As a result, although the number of missingness patterns may grow exponentially with the number of variables $p$, the number of base prediction models required increases only linearly.
2.2 Inferential target
In this paper, we focus on obtaining valid inference on a $d$-dimensional parameter $\bm{\theta}$} in a Z-estimation framework based on a dataset with general missingness patterns under the MAR assumption. Z-estimation framework is a widely used framework for establishing an estimator in statistical literature. Several most commonly used statistical model such as generalized linear model and generalized estimating equations belong to the Z-estimation framework. Formally, define the true population parameter $\bm{\theta}$$^{\ast}$ as the unique solution of
$\mathbb{E}_{\mathbb{P}_{\bm{Y}}}[{\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}})]=\bm{0},$
where $\bm{\psi}$} are known vector-valued maps to a typically $d$-dimensional space, as defined in van der Vaart, (2000). For independent and identically distributed, fully observed data $\mathcal{D}$} of size $N$, a Z-estimator $\widehat{{\bm{\theta}}}_{\text{Full}}$ is obtained by solving the following empirical estimating equations:
$\frac{1}{N}\sum^{N}_{i=1}{\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}})=\bm{0}.$ (1)
Under certain regularity conditions introduced in van der Vaart, (2000) and detailed in Section S1.2 in the Supplementary Materials, $\widehat{{\bm{\theta}}}_{\text{Full}}$ is asymptotically normal with mean $\bm{\theta}$$^{\ast}$ and variance $\dot{{\bm{\psi}}_{Full}}^{-1}\mathbb{E}[{\bm{\psi}}{\bm{\psi}}^{\top}]\dot{{\bm{\psi}}_{Full}}^{-1\top}$, where ${\bm{\psi}}={\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}}^{\ast})$ and $\dot{{\bm{\psi}}_{Full}}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}}{\partial{\bm{\theta}}}\Bigr{|}_{{\bm{\theta}}={\bm{\theta}}^{\ast}}\right]$.
2.3 Rationale of prediction-based inference
In the presence of missing values, we will propose a general approach and provide the theoretical guarantee of valid inference in Section 3. The proposed approach couples the insights from the original outcome-based prediction-powered inference (PPI) framework (Angelopoulos et al., 2023a, ; Angelopoulos et al., 2023b, ) and the classical augmented inverse probability weighted (AIPW) estimators, of which outcome-based PPI is a special case (Gronsbell et al.,, 2024). It is useful to overview the rationale behind the original outcome-based PPI approach, which is designed to incorporate additional information from observations with missing outcomes and the predicted outcomes. Technically, they can be written as a difference between a base estimate (CCA estimate) $\widehat{{\bm{\theta}}}_{\text{CC}}$ and an estimate of zero $\widehat{\mathbf{\Delta}}$: [^2] The notation of $\widehat{\mathbf{\Delta}}$ in this paper differs from the one used in Angelopoulos et al., 2023a to denote a rectifier to correct the bias introduced by ML predictions. See Example 1. The current form is more convenient for constructing more efficient estimators and their theoretical analyses.
$\widehat{{\bm{\theta}}}_{\text{PP}}=\widehat{{\bm{\theta}}}_{\text{CC}}-\widehat{\mathbf{\Delta}}$,
[…]
Example 1 (PPI for mean outcome estimation).
We illustrate with a simple example based on population mean estimation of a scalar outcome $Y$ in the presence of outcome missingness introduced in Angelopoulos et al., 2023a . Suppose the data are independently and identically distributed according to a population distribution comprised of two partitions of size $m$ and $M$ with marginal mean outcome $\mu=\mathbb{E}[Y]$. For the first partition, we observe the outcome $Y^{\tt lab}_{i}$ and their corresponding ML-predicted alternatives $\widehat{Y}^{\tt lab}_{i}$: ${\mathcal{D}}_{\tt lab}=\{(Y^{\tt lab}_{i},\widehat{Y}^{\tt lab}_{i})\}^{m}_{i=1}$, which is referred to as the “labeled subset”. For the second, often larger, partition of data, we do not get to observe the outcome but can obtain their ML-predicted surrogates, i.e., ${\mathcal{D}}_{\tt unlab}=\{\widehat{Y}^{\tt unlab}_{i}\}^{M}_{i=1}$, referred to as “unlabeled subset”. The full data distribution of $Y$ between the two partitions are identical, i.e., the outcome is MCAR. For both partitions, the predictions are obtained from other observed covariates and prediction models that have been externally trained to predict $Y$ using these covariates. As such, the pair (Y,Y^)($\widehat{Y}$) has the same distribution between the labeled and unlabled partitions, although $Y$ is missing for the unlabled subset.
The CCA estimator for $\mu$ is the empirical mean based on the complete subset of data: $\widehat{\theta}_{\text{CC}}=\bar{Y}^{\tt lab}$, and the corresponding estimation of zero is computed as $\widehat{\Delta}=\bar{\widehat{Y}}^{\tt lab}-\bar{\widehat{Y}}^{\tt unlab}$ because the predicted outcomes in the labeled and unlabeled partitions share the same distribution under MCAR. By rewriting $\widehat{\theta}_{\text{PP}}=\widehat{\theta}_{\text{CC}}-\widehat{\Delta}=\bar{Y}^{\tt lab}-(\bar{\widehat{Y}}^{\tt lab}-\bar{\widehat{Y}}^{\tt unlab})=\bar{\widehat{Y}}^{\tt unlab}-\underbrace{(\bar{\widehat{Y}}^{\tt lab}-\bar{Y}^{\tt lab})}_{\tt``rectifier"}$, we have
$\mathrm{Var}(\widehat{\theta}_{\text{PP}})=\frac{1}{M}\mathrm{Var}(\widehat{Y}^{\tt unlab}_{i})+\frac{1}{m}\mathrm{Var}(Y^{\tt lab}_{i}-\widehat{Y}^{\tt lab}_{i}),$
where the numerator of the first term is determined by the variance of prediction; the numerator of the second term is determined by the accuracy of the outcome prediction. $\widehat{\theta}_{\text{PP}}$ will have a small variance when $M\gg m$ to let the second term dominate and the machine learning prediction is precise enough such that the variance in the second term is small. To see this, when $M\gg m$, we will have $\mathrm{Var}(\widehat{\theta}_{\text{PP}})\approx\frac{1}{m}\mathrm{Var}(Y^{\tt lab}_{i}-\widehat{Y}^{\tt lab}_{i})<\mathrm{Var}(\widehat{\theta}_{\text{CC}})=\frac{1}{m}\mathrm{Var}(Y_{i}^{\tt lab})$ when the predictions are of good quality, or more precisely when the correlation satisfies $\text{Corr}(Y,\widehat{Y})>\frac{1}{2}\sqrt{\mathrm{Var}(Y)/\mathrm{Var}(\widehat{Y})}$.
However, it is critical to understand when $\widehat{{\bm{\theta}}}_{\text{PP}}$ is more efficient than the complete-case analysis in settings other than the mean estimation problem. Gronsbell et al., (2024) analyzed the efficiency of the $\widehat{{\bm{\theta}}}_{\text{PP}}$ by deriving its influence function under the linear regression setting, and connect it to the class of AIPW estimators. Specifically, Gronsbell et al., (2024) shows that $\widehat{{\bm{\theta}}}_{\text{PP}}$ is an AIPW estimator and is not the most efficient in its class. Although the optimal estimator is difficult to obtain, a weighted version of the PPI estimator was proposed to guarantee the efficiency improvement compared to the complete-case analysis (Gronsbell et al.,, 2024). The estimator is of the form $\widehat{{\bm{\theta}}}_{\text{Chen}}=\widehat{{\bm{\theta}}}_{\text{CC}}-\widehat{\mathbf{W}}_{\text{Chen}}\widehat{\mathbf{\Delta}}$, where $\widehat{\mathbf{W}}_{\text{Chen}}$ is a tuning parameter estimated from the data, originally proposed in  Chen and Chen, (2000) under a different double-sampling application context but the same setup. Furthermore, recent modifications to the $\widehat{{\bm{\theta}}}_{\text{PP}}$ all consider multiplying $\widehat{\mathbf{\Delta}}$ with a data-driven weight to seek efficiency improvement over $\widehat{{\bm{\theta}}}_{\text{CC}}$ (Angelopoulos et al., 2023b, ; Miao et al.,, 2023; Gan et al.,, 2024).
2.4 Limitations of existing prediction-based inference methods
We now discuss the limitations of current methods, which we will address. First, existing PPI methods do not account for the missing-at-random mechanism. For example, in mean estimation, suppose $Y$ is the annual income response in a questionnaire. An auxiliary variable, the level of education, affects the missing probability of $Y$, such that respondents with lower level of education tend to skip the response to the income question, leaving their $Y$ missing. In such a case, the CCA estimator $\widehat{\theta}_{\text{CC}}$ itself will become biased, as it can only reflect the income levels of the population with higher education levels. However, although being designed as an estimate of zero, $\widehat{\Delta}$ in this circumstance can only capture the mean differences of the predicted income between the two education groups, rather than the differences in true income, providing no guaranteed consistency of the result $\widehat{\theta}_{\text{PP}}$ to the average income level of the entire population. Second, existing methods only consider missing outcome or blockwise missingness scenarios, although in practice general missing data patterns commonly appear. For example, in responding to a questionnaire, respondents may randomly skip questions, leading to non-monotone missingness in the resulting dataset; in a longitudinal study, participants may skip a visit and return in the next visit (Sun and Tchetgen Tchetgen,, 2018). In such cases, existing methods can only be carried out by giving up the incomplete covariates, or truncating the data at the point where first missingness occurred.
3 Pattern-Stratified Prediction-Powered Inference
3.1 Method
Figure 1: Proposed statistical procedure for ensuring valid prediction-based inference in datasets with general missingness patterns under the missing at random (MAR) assumption. (a) we observe a set of samples with missing values. We assign each missingness pattern with an index. (b) for the $k$-th missingness pattern, we collect the samples with ${\mathcal{C}}_{i}=k$, and impute the missing values with ML predictions. Weighted complete-case analysis are conducted over the imputed datasets to obtain $\widehat{{\bm{\gamma}}}_{2,k}$; (c) for completely observed samples, we obtain the weighted analysis estimates $\widehat{{\bm{\theta}}}_{\text{WCC}}$. For each $k$-th missingness pattern, we mask the observed values from the fully observed samples to construct each missingness pattern. We then impute the masked values with machine learning prediction and conduct weighted complete-case analysis to obtain $\widehat{{\bm{\gamma}}}_{1,k}$.
The PS-PPI estimator is defined as the difference between the weighted CCA estimator $\widehat{{\bm{\theta}}}_{\text{WCC}}$, computed from the fully observed subset, and a weighted summation of $K$ estimators of zero $\widehat{\mathbf{\Delta}}_{k}$ derived from each missingness pattern:
$\widehat{{\bm{\theta}}}_{\text{PS-PPI}}=\widehat{{\bm{\theta}}}_{\text{WCC}}-\sum^{K}_{k=1}\widehat{\mathbf{W}}_{k}\widehat{\mathbf{\Delta}}_{k}.$ (2)
Here we introduce these notation in turn. First, $\widehat{{\bm{\theta}}}_{\text{WCC}}$ is the weighted complete-case analysis estimator obtained from the following weighted estimating equations:
$\sum_{i=1}^{N}{\bm{\psi}}^{(i)}_{\text{WCC}}=\sum_{i=1}^{N}\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}_{\infty}(\bm{Y}_{i})}{\bm{\psi}}\left(\bm{Y}_{i};{\bm{\theta}}\right)=\bm{0}.$ (3)
[…]
$\sum^{N}_{i=1}{\bm{m}}^{(i)}=\sum^{N}_{i=1}{\bm{m}}(G_{{\mathcal{C}}_{i}}(\bm{Y}_{i});{\bm{\omega}})=\bm{0},$ (4)
where $\bm{m}$} is a vector-valued function that has the same dimension as $\bm{\omega}$} and satisfies regularity conditions detailed in Appendix S1.2.
Second, $\widehat{\mathbf{\Delta}}_{k}$ is the estimate of zero from the $k$-th missingness pattern. Specifically, we define $\widehat{\mathbf{\Delta}}_{k}=\widehat{{\bm{\gamma}}}_{1,k}-\widehat{{\bm{\gamma}}}_{2,k}$, where $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{2,k}$ are the solutions of the weighted estimating equations:
$\sum_{i=1}^{N}{\bm{\psi}}^{(i)}_{1,k}=\sum^{N}_{i=1}\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}_{\infty}(\bm{Y}_{i})}{\bm{\psi}}\left(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}}_{1,k}\right)=\bm{0},$ (5)
and
$\sum_{i=1}^{N}{\bm{\psi}}^{(i)}_{2,k}=\sum^{N}_{i=1}\frac{I({\mathcal{C}}_{i}=k)}{\widehat{\pi}_{k}(G_{k}(\bm{Y}_{i}))}{\bm{\psi}}\left(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}}_{2,k}\right)=\bm{0},$ (6)
respectively. That is, we first compute $\widehat{{\bm{\gamma}}}_{1,k}$ by substituting the values in the missing entries with predictions generated from $h_{k}$, and solve the weighted estimating equations based on complete-case data. Second, we compute $\widehat{{\bm{\gamma}}}_{2,k}$ by imputing the actual missing entries with the predictions from $h_{k}$ using data with $k$-th missingness pattern. We will show that such construction leads to an consistent and asymptotically normal estimation of zero; as shown in Proposition 1, the population parameter of the estimating equations ${\bm{\psi}}_{1,k}$ and ${\bm{\psi}}_{2,k}$ are both ${\bm{\gamma}}^{\ast}_{k}$. To have such proposition, we need the following assumption:
Assumption 1 (Z-estimation target parameter under population-level mask-and-impute distributions).
There exists a ${\bm{\gamma}}^{\ast}_{k}$ such that it is the unique solution of
$\mathbb{E}_{\bm{Y}}[{\bm{\psi}}(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}})]=\bm{0}.$
Proposition 1 (Estimation of zero).
Under Assumption 1, $({\bm{\omega}}^{\ast},{\bm{\gamma}}^{\ast}_{k})$ will lead to both $\mathbb{E}{\bm{\psi}}_{1,k}=\bm{0}$ and $\mathbb{E}{\bm{\psi}}_{2,k}=\bm{0}$, where the expectations are taken over the full data distribution $\mathbb{P}_{\bm{Y}}$.
The proof is detailed in S1.1. The intuition behind Proposition 1 is that, after accounting for the missing mechanism, the “pseudo” imputed dataset derived from the complete-case data used in Equation (5) is distributed as $\mathbb{P}^{(k)}_{\bm{Y}_{\text{imp}}}$, which is the same distribution as the imputed dataset from data with $k$-th missingness pattern used in Equation (6).
Finally, $\{\widehat{\mathbf{W}}_{k}\}$ are a set of tuning parameters in the form of weight matrices with dimension $d\times d$. $\{\widehat{\mathbf{W}}_{k}\}$ may depend on the observed data, and in Corollary 3.2 we provide a specific construction that guarantee efficiency gains over WCCA.
Asymptotic properties of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$.
We provide Theorem 1 to show that $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is asymptotically normal.
Theorem 1 (Asymptotic normality of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$).
[…]
The proof is detailed in Section S1.2 in the Supplementary Materials. The asymptotic variance-covariance matrix in the proof of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ accounts for the uncertainty of the estimated propensity score models. Following Sun and Tchetgen Tchetgen, (2018) and Robins et al., (1994), we provide a convenient robust sandwich estimators that ignores the uncertainty from the propensity score model. Such estimators results in a slightly more conservative yet valid and implementable sandwich variance for $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ when the propensity scores are estimated; the form is shown below. Our numerical experiments in Section 4 validate efficiency advantages of the PS-PPI estimator based on this strategy.
Corollary 3.1 (Asymptotic variance-covariance matrix of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$).
[…]
$\Sigma_{\text{PS-PPI}}=\Sigma_{{\bm{\theta}}}-\sum^{K}_{k=1}\mathbf{W}_{k}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}-\sum^{K}_{k=1}(\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}})^{\top}\mathbf{W}_{k}^{\top}+\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\mathbf{W}_{k}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\mathbf{W}_{k^{\prime}}^{\top}+\sum^{K}_{k=1}\mathbf{W}_{k}\Sigma_{{\bm{\gamma}}_{2,k}}\mathbf{W}_{k}^{\top}.$ (7)
Here, $\Sigma_{{\bm{\theta}}}$ is the asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{WCC}}$. $\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$ is the asymptotic covariance between $\widehat{{\bm{\theta}}}_{\text{WCC}}$ and $\widehat{{\bm{\gamma}}}_{1,k}$. $\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$ is the asymptotic covariance between $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{1,k^{\prime}}$; when $k=k^{\prime}$, it is the asymptotic variance of $\widehat{{\bm{\gamma}}}_{1,k}$. $\Sigma_{{\bm{\gamma}}_{2,k}}$ is the asymptotic variance of $\widehat{{\bm{\gamma}}}_{2,k}$.
From Equation (7), $\Sigma_{\text{PS-PPI}}$ is a function of $\{\mathbf{W}_{k}\}=(\mathbf{W}_{1},...,\mathbf{W}_{K})$ based on which we may seek the optimal $\{\mathbf{W}^{\ast}_{k}\}$ such that $\Sigma_{\text{PS-PPI}}(\{\mathbf{W}^{\ast}_{k}\})\preceq\Sigma_{\text{PS-PPI}}(\{\mathbf{W}_{k}\})$ for any $\{\mathbf{W}_{k}\}$, where, for two square matrices $\mathbf{A}$ and $\mathbf{B}$, $\mathbf{A}\preceq\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive semidefinite; $\mathbf{A}\prec\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive definite. Such optimal $\{\mathbf{W}_{k}\}$ has no closed-form and may be computationally expensive to solve. However, when assuming identical $\mathbf{W}_{k}$ across $k$, we have a closed-form solution of $\{\mathbf{W}^{\ast}_{k}\}$, and the resulting $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is guaranteed to be more efficient than $\widehat{{\bm{\theta}}}_{\text{WCC}}$, as we state in Corollary 3.2 and prove in Appendix S1.3.
Corollary 3.2.
Suppose $\mathbf{W}_{k}=\mathbf{W}$ are the same across $k$, then $\mathbf{W}^{\ast}=\left(\sum^{K}_{j=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,j}}\right)\times\left[\left(\sum^{K}_{j=1}\sum^{K}_{j^{\prime}=1}\Sigma_{{\bm{\gamma}}_{1,j},{\bm{\gamma}}_{1,j^{\prime}}}\right)+\left(\sum^{K}_{j=1}\Sigma_{{\bm{\gamma}}_{2,j}}\right)\right]^{-1}$ leads to an optimal $\Sigma_{\text{PS-PPI}}$. Accordingly, the asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is
$\Sigma_{\text{PS-PPI}}=\Sigma_{{\bm{\theta}}}-\left(\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}\right)\times\left[\left(\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\right)+\left(\sum^{K}_{k=1}\Sigma_{{\bm{\gamma}}_{2,k}}\right)\right]^{-1}\times\left(\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}\right)^{\top}.$ (8)
Furthermore, we have $\Sigma_{\text{PS-PPI}}\preceq\Sigma_{{\bm{\theta}}}$, where $\Sigma_{{\bm{\theta}}}$ is the asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{WCC}}$. $\Sigma_{\text{PS-PPI}}=\Sigma_{{\bm{\theta}}}$ if and only if $\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}=\bm{0}_{d\times d}$.
The $\mathbf{W}^{\ast}$ defined in Corollary 3.2 is controlled by three components: 1) $\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$: the covariance between $\widehat{{\bm{\theta}}}_{\text{WCC}}$ and each of the $\widehat{{\bm{\gamma}}}_{1,k}$; 2) $\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$: the covariance between any two of the $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{1,k^{\prime}}$; 3) $\Sigma_{{\bm{\gamma}}_{1,k}}$ and $\Sigma_{{\bm{\gamma}}_{2,k}}$: the variance of $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{2,k}$ under each missingness pattern. Specifically, a large weight matrix is assigned to all the estimation of zero $\widehat{\mathbf{\Delta}}_{k}$ when: 1) we have large covariance between $\widehat{{\bm{\theta}}}_{\text{WCC}}$ and $\widehat{{\bm{\gamma}}}_{1,k}$, which suggests that the imputations are faithfully reflect the missing true values; 2) we have smaller variance in $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{2,k}$, which suggests that we can construct a precise estimation of zero with small variance; 3) we have small covariances between any pair of the $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{1,k^{\prime}}$, which suggests that we can efficiently aggregate across different missingness patterns without introducing significantly additional uncertainty.
Because the optimal $\mathbf{W}^{*}$ depends on the variance-covariance matrices, it can be estimated by letting $\widehat{\mathbf{W}}=\left(\sum^{K}_{j=1}\widehat{\Sigma}_{{\bm{\theta}},{\bm{\gamma}}_{1,j}}\right)\times\left[\left(\sum^{K}_{j=1}\sum^{K}_{j^{\prime}=1}\widehat{\Sigma}_{{\bm{\gamma}}_{1,j},{\bm{\gamma}}_{1,j^{\prime}}}\right)+\left(\sum^{K}_{j=1}\widehat{\Sigma}_{{\bm{\gamma}}_{2,j}}\right)\right]^{-1}$.
The theoretical results from Theorem 1 and Corollary 3.2 allow us to compute the PS-PPI estimator and the corresponding uncertainty quantification for $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ in a Z-estimation framework. In particular, $\widehat{{\bm{\theta}}}_{WCC}$, $\widehat{{\bm{\gamma}}}_{1,k}$, $\widehat{{\bm{\gamma}}}_{2,k}$ can be obtained by solving the estimating equations in Equation (3), (5), (6). For uncertainty quantification, $\widehat{\Sigma}_{{\bm{\theta}}}$, $\widehat{\Sigma}_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$, $\widehat{\Sigma}_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$, and $\widehat{\Sigma}_{{\bm{\gamma}}_{2,k}}$ can be calculated by deriving the specific closed-form given a Z-estimation framework based on results in Equation (S1.2), (S1.3), (S1.4), (S1.5), and compute their empirical estimates. We provide a more detailed instruction along with the results for linear regression and logistic regression as examples for reference, see Section S1.4 in the Supplementary Materials for details. For those Z-estimation framework without easily-computable empirical estimates of the variance-covariance matrices, one may apply the resampling strategies to address such challenge, for example, Bootstrap and delete-d Jackknife (Efron,, 1982). Under certain technical conditions, both Bootstrap and delete-$d$ Jackknife can produce a consistent variance-covariance estimation (Shao and Wu,, 1989; Shao,, 1989; Cheng and Huang,, 2010).
It is worth noting that the PS-PPI method could be implemented in a more convenient way when there exists a well-developed package for the target Z-estimation framework used for weighted complete-case analysis. Specifically, Equation (3), (5), (6) can be viewed as separate weighted complete-case analysis on different data using the same Z-estimation framework. Hence, no significant additional efforts need to be made to compute $\widehat{{\bm{\theta}}}_{WCC}$, $\widehat{{\bm{\gamma}}}_{1,k}$, $\widehat{{\bm{\gamma}}}_{2,k}$, as well as variance-covariance matrices $\widehat{\Sigma}_{{\bm{\theta}}}$ and $\widehat{\Sigma}_{{\bm{\gamma}}_{2,k}}$ once a weighted estimation software is available. The major effort for the implementation will be, therefore, to compute $\widehat{\Sigma}_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$ and $\widehat{\Sigma}_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$. As discussed above, this could be implemented by either deriving the specific closed-form based on Equation (S1.3) and (S1.4), or applying resampling strategies. Since both $\widehat{\Sigma}_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$ and $\widehat{\Sigma}_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$ are variance-covariance matrices only relevant to fully-observed data, the resampling will be only conducted among fully-observed data. In this paper, we take the delete-$1$ Jackknife as our strategy to estimate the covariances. See Algorithm 1 for the full algorithmic details to obtain $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ and the corresponding uncertainty estimates.
3.2 Connections to the existing literature
Connections to the setting when only outcome is missing.
We show that the PS-PPI procedure subsumes the outcome-based PPI framework in the existing literature. Specifically, assume that in addition to the fully observed data subset, there is only one missingness pattern with only the outcome, say $Y_{i1}$, missing. Furthermore, assume that the underlying missing mechanism is MCAR, i.e., the propensity score model $\bm{\pi}(\cdot)$ is a constant 2-dimensional vector $(\pi_{1},\pi_{\infty})$ across all samples, where $\pi_{1}$ is the probability of missing outcome, and $\pi_{\infty}$ is the probability of being fully observed. The PS-PPI estimator under this setting become $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}=\widehat{{\bm{\theta}}}_{\text{CC}}-\widehat{\mathbf{W}}(\widehat{{\bm{\gamma}}}_{1,1}-\widehat{{\bm{\gamma}}}_{2,1})$. Note that here we have the CCA estimate $\widehat{{\bm{\theta}}}_{\text{CC}}$ instead of $\widehat{{\bm{\theta}}}_{\text{WCC}}$, because the weights are the same across all samples. In  Gronsbell et al., (2024) and Chen and Chen, (2000), their proposed estimator is $\widehat{{\bm{\theta}}}_{\text{Chen}}=\widehat{{\bm{\theta}}}_{\text{CC}}-\widehat{\mathbf{W}}_{\text{Chen}}(\widehat{{\bm{\gamma}}}_{\text{lab}}-\widehat{{\bm{\gamma}}}_{\text{all}})$. Here, $\widehat{{\bm{\theta}}}_{\text{CC}}$ follows the same definition in this paper and $\widehat{{\bm{\gamma}}}_{\text{lab}}=\widehat{{\bm{\gamma}}}_{1,1}$ with $K=1$ missingness pattern (missing outcome only); $\widehat{{\bm{\gamma}}}_{\text{all}}$ is a variant of $\widehat{{\bm{\gamma}}}_{2,1}$ in this paper, which is estimated by all the data in ${\mathcal{D}}_{\text{obs}}$, instead of only data with ${\mathcal{C}}_{i}=1$; $\widehat{\mathbf{W}}_{\text{Chen}}$ is a tuning parameter matrix estimated from the data. We show the asymptotic equivalence between $\widehat{{\bm{\theta}}}_{\text{Chen}}$ and $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ Proposition 2.
Proposition 2.
Assume that only the outcome, say $Y_{i1}$, is possibly missing, we have that $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ and $\widehat{{\bm{\theta}}}_{\text{Chen}}$ are asymptotically equivalent in the linear regression and MCAR setting considered in Gronsbell et al., (2024).
Proposition 2 is proved by showing that their influence functions are the same. Note that in $\widehat{{\bm{\theta}}}_{\text{Chen}}$, $\widehat{{\bm{\gamma}}}_{\text{all}}$ is estimated using both partially and fully-observed data. This is different from $\widehat{{\bm{\gamma}}}_{2,1}$ in $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ which is estimated using only the missing-outcome subset of data.  Gronsbell et al., (2024) argue that this modification guarantees a better efficiency when $\mathbf{W}$ is the identity matrix than the original PPI procedure (Angelopoulos et al., 2023a, ). However, when an optimal $\mathbf{W}^{\ast}$ is selected, Proposition 2 indicates that such efficiency improvement is absorbed into the $\mathbf{W}^{\ast}$, thereby the two estimators become asymptotically equivalent. We elaborate this equivalence using the example of perfect predictions, i.e., the predictions are exactly the true values. Specifically, we have the following remark:
Remark 3.1.
Suppose perfect predictions are available, then for $\widehat{{\bm{\theta}}}_{\text{Chen}}$, $\widehat{\mathbf{W}}_{\text{Chen}}={\bm{I}}$; for $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$, $\widehat{\mathbf{W}}=\widehat{\pi}_{1}{\bm{I}}$.
Remark 3.1 can be proved by deriving the closed-form of variance-covariance matrices under linear regression and obtain the closed form of $\Sigma_{\text{PS-PPI}}$. From Remark 3.1, when perfect predictions are available, $\widehat{{\bm{\theta}}}_{\text{Chen}}=\widehat{{\bm{\gamma}}}_{\text{all}}$, which is equivalent to the results from the underlying full data since the predictions are perfect. On the other hand, $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}=\widehat{{\bm{\theta}}}_{\text{CC}}-\widehat{\pi}_{1}(\widehat{{\bm{\gamma}}}_{1,1}-\widehat{{\bm{\gamma}}}_{2,1})=\widehat{\pi}_{\infty}\widehat{{\bm{\theta}}}^{(\infty)}_{\text{CC}}+\widehat{\pi}_{1}\widehat{{\bm{\theta}}}^{(1)}_{\text{CC}}$. Here $\widehat{{\bm{\theta}}}^{(\infty)}_{\text{CC}}$ is the CCA estimate obtained from the complete-case subset; $\widehat{{\bm{\theta}}}^{(1)}_{\text{CC}}$ is the CCA estimate obtained from the underlying full data of the missing-outcome subset of data. We are able to replace $\widehat{{\bm{\gamma}}}_{2,1}$ with $\widehat{{\bm{\theta}}}^{(1)}_{\text{CC}}$ because the perfect predictions make the imputed dataset exactly the same as the underlying full data. As a result, $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ can be viewed as a weighted sum of the full data estimator from complete-case and missing-outcome subset of data, respectively; the weights are the proportion of the sample size of each subsets $\widehat{\pi}_{\infty}$ and $\widehat{\pi}_{1}$. Therefore, $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ can be seen as an inverse variance weighted version of $\widehat{{\bm{\theta}}}_{\text{Chen}}$.
$\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ also shares core insights with other existing outcome-based PPI literature (Angelopoulos et al., 2023b, ; Miao et al.,, 2023; Gan et al.,, 2024), although they may not be asymptotically equivalent. In particular, Angelopoulos et al., 2023b ; Miao et al., (2023); Gan et al., (2024) introduce tunable weights within the estimating equations, and derive similar estimators with guarantees of being more efficient than CCA. All of these methods are based on the outcome-missing setting. The proposed PS-PPI procedure introduces the tunable weights when combining the estimators of zero obtained from each missingness pattern. Each estimator of zeros can be obtained by completing two weighted CCA steps (for obtaining $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{2,k}$), each of which is easy-to-implement by leveraging existing weighted CCA software, often available, for the same Z-estimation problem upon careful construction of imputed datasets following Algorithm 1.
Connections to the blockwise covariates missing setting.
In addition, the PS-PPI procedure also subsumes the blockwise covariates missing scenario considered in Kluger et al., (2025). In the framework of PS-PPI, the missing blockwise covariates setting is similar to the missing outcome settings, where we only have two missingness patterns: 1) fully observed; 2) $s$ different variables in $\bm{Y}_{i}$ are missing simultaneously. Furthermore, the missing mechanism is assumed to be MAR while the propensity score model is assumed to be known. In this setting, we show the asymptotic equivalence between $\widehat{{\bm{\theta}}}_{\text{PTD}}$ proposed in Kluger et al., (2025) and $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ in Proposition 3:
Proposition 3.
$\widehat{{\bm{\theta}}}_{\text{PTD}}$ and $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ are asymptotically equivalent.
Connections to semiparametric inference literature.
In the context of semiparametric inference theory, $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ belongs to the class of AIPW estimators, which can be shown by its influence functions derived from the proof of Theorem 1 (detailed in Section S1.5). In the context of semiparametric theory, one may be interested in searching for the most efficient estimators within the AIPW class estimators under specific settings (Robins et al.,, 1994, 1995; Chaudhuri and Guilkey,, 2016). However, under general missingness patterns and with an assumption of MAR, the optimal AIPW Z-estimators, despite being the most efficient estimator in integrating the predicted surrogates, suffer from difficulties in implementation as it generally does not have closed-form and would require iterative computation to solve. We refer readers to Robins et al., (1994); Tsiatis, (2006); Chaudhuri and Guilkey, (2016) for a more in-depth discussion. However, $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ does not aim to guarantee optimal influence functions within the AIPW class estimators. Rather, the major advantage of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is its ability to be easily applied in analysis without significant additional efforts in the implementation. With the rapid development of deep learning and machine learning techniques, highly precise ML predictions become increasingly available. Enabling analysts to easily combine these ML-based prediction surrogates with their statistical models to perform the analysis without additional significant implementation efforts is necessary. As discussed above, our proposed PS-PPI procedure circumvents the implementation challenges by allowing researchers to apply existing statistical software designed for weighted complete-case analysis to obtain the PS-PPI estimator: each component of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ (i.e., $\widehat{{\bm{\theta}}}_{\text{WCC}},\{\widehat{{\bm{\gamma}}}_{1,k}\},\{\widehat{{\bm{\gamma}}}_{2,k}\}$) can be obtained from weighted complete-case analysis algorithms and the weights can be obtained by resampling techniques (see Algorithm 1).
Connections to multiple imputation.
Multiple imputation based on Rubin’s rule is a well-established methodology that has been widely applied to account for the uncertainty introduced by imputation models to guarantee valid downstream statistical inference (Little and Rubin,, 2019). Multiple imputation often assumes a scenario where the observed data are the only resources, and imputation models are fitted based on these observed data to generate repeated imputations. To guarantee valid inference from these multiply-imputed datasets, multiple imputation imposes assumptions on the imputation models (Rubin,, 1987). Specifically, multiple imputation requires a proper imputation procedure—the unbiasedness of the point estimates and the estimated variance averaged over multiple imputed dataset, as well as the between-imputation variability—to derive a valid statistical inference (Murray,, 2018; Rubin,, 1996). Such scenario is different from the scenario assumed in PS-PPI, as well as the broader prediction-based inference literature. In additional to the observed data, current prediction-based inference literature assumes that analysts also have access to some potentially precise predictions generated from models that are trained by external data. Commonly, these external models are ML/DL-based, and therefore are either challenging to manipulate to satisfy any assumptions, or beyond analysts’ control. This is the core reason why prediction-based inference literature, including this work, do not impose any assumptions towards the prediction models. However, under the MAR missing mechanism, this weak assumption on the quality of the prediction models comes at the expense of requiring the correct specification of the propensity score model. We present a numerical experiment to investigate PS-PPI’s robustness against propensity score misspecification in Section 4 and leave detailed discussion towards propensity score model specification in Section 5.
Algorithm 1 Pseudocode for obtaining $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ and uncertainty estimates.
1:An observed dataset ${\mathcal{D}}_{\text{obs}}=\{(G_{{\mathcal{C}}_{i}}(\bm{Y}_{i}),{\mathcal{C}}_{i})\}^{N}_{i=1}$ with $K$ missingness patterns (${\mathcal{C}}=1,\ldots,K$) and a subset of complete data ($\mathcal{C}$$\infty$); a set of external machine learning prediction models $h_{k}$ imputing missing entries under each missingness pattern; a propensity score models (either known or estimated from the data) $\widehat{\bm{\pi}}(\cdot)\in\mathbb{R}^{K+1}$, with each dimension $\pi_{k}(\cdot)$ representing the probability of ${\mathcal{C}}=k$. A Z-estimation procedure $f({\tt complete\leavevmode\nobreak\ data};{\tt weighting\leavevmode\nobreak\ function})$ that supports weighted complete-case analysis and returns the estimates and the variance estimates for the parameter of interest $\bm{\theta}$.
2:for k=1,,Kk=1,...,K do
3:  For subjects with ${\mathcal{C}}_{i}=k$, impute the missing values with ML predictions and obtain the ML-imputed dataset ${\mathcal{D}}_{2,k}=\{\bm{Y}^{(k)}_{\text{imp},i}:{\mathcal{C}}_{i}=k\}$.
4:  Compute
$\widehat{{\bm{\gamma}}}_{2,k},\widehat{\Sigma}_{{\bm{\gamma}}_{2,k}}\leftarrow f\left({\mathcal{D}}_{2,k},\pi^{-1}_{k}(\cdot)\right).$
end for
5:Denote the set of completely observed data ${\mathcal{D}}_{\text{CC}}=\{\bm{Y}_{i}:{\mathcal{C}}_{i}=\infty\}$.
6:for $j=1,...,|{\mathcal{D}}_{\text{CC}}|$ do $\triangleright$ Jackknife
7:  Drop the $j$-th completely observed data and obtain ${\mathcal{D}}^{(j)}_{\text{CC}}$.
8:  Compute
$\widehat{{\bm{\theta}}}^{(j)}_{\text{WCC}}\leftarrow f\left({\mathcal{D}}^{(j)}_{\text{CC}},\pi^{-1}_{\infty}(\cdot)\right).$
9:  for k=1,,Kk=1,...,K do
10:   Mask observed values based on $k$-th missingness patterns and impute predictions to obtain the pseudo ML-imputed dataset ${\mathcal{D}}^{(j)}_{1,k}=\{\bm{Y}^{(k)}_{\text{imp},i}:\bm{Y}_{i}\in{\mathcal{D}}^{(j)}_{\text{CC}}\}$.
11:   Compute
$\widehat{{\bm{\gamma}}}^{(j)}_{1,k}\leftarrow f\left({\mathcal{D}}^{(j)}_{1,k},\pi^{-1}_{\infty}(\cdot)\right).$
  end for end for
12:Compute the jackknife estimates of $\widehat{{\bm{\theta}}}_{\text{WCC}},\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{\Sigma}_{{\bm{\theta}}},\widehat{\Sigma}_{{\bm{\gamma}}_{1,k}},\widehat{\Sigma}_{{\bm{\theta}},{\bm{\gamma}}_{1,k}},\widehat{\Sigma}_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$.
[…]
15:Compute $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ from Equation (2): $\widehat{\bm{\theta}}\leftarrow\widehat{{\bm{\theta}}}_{\text{WCC}}-\sum_{k=1}^{K}\widehat{\mathbf{W}}\left(\widehat{{\bm{\gamma}}}_{1,k}-\widehat{{\bm{\gamma}}}_{2,k}\right).$
16:Compute the estimated variance $\widehat{\Sigma}_{\text{PS-PPI}}$ based on Equation (8) by plugging in estimated quantities.
[…]
4 Numerical Experiments
In this section, we conduct data simulation based on synthetic data to show the operating characteristics of the proposed method to corroborate the theoretical prediction that the proposed approach produces consistent estimation for parameters of interest and yields valid uncertainty quantification. We compare our proposed approach with the following baseline methods: 1) unweighted complete-case analysis, where parameter estimation is performed using only fully observed data (“CCA”); 2) complete-case analysis of inverse probability weighted (IPW), where estimation is carried out using fully observed data weighted by true or estimated propensity scores (“WCCA”); 3) multiple imputation (MI) (Rubin,, 1976), a classical and popular method to deal with dataset with non-monotone missingness pattern under the MAR assumption. In this work, we use the implementation of MI provided by the MICE 3.16.0333https://github.com/amices/mice package in R; 4) PPI++ (Angelopoulos et al., 2023b, ), a recent extension of the prediction-powered inference framework that guarantees better estimation efficiency compared to the CCA approach. For PPI++, both fully observed cases and cases with missing outcomes are used for parameter estimation; cases with missing covariates will be dropped.
4.1 Experimental setup
To demonstrate the operating characteristics of the proposed method under various ML prediction qualities, here we focus on a typical analytic model specified by a multivariate linear regression as follows:
$Y_{i}=\beta_{0}+\beta_{1}X_{i1}+\beta_{2}X_{i2}+\epsilon_{i},$
where the covariates are generated from $X_{i1}=0.1\exp(Z_{i1})+\tau_{i}$, $X_{i2}=\sin(Z_{i2})+\nu_{i}$, with $(Z_{i1},Z_{i2})$ randomly and independently drawn from a bivariate normal distribution with mean zero and a homogeneous marginal variance $\sigma_{z}^{2}$ and Pearson correlation $\rho$. The error terms are independently distributed as $\epsilon_{i}\sim{\mathcal{N}}(0,\sigma^{2})$, $\tau_{i}\sim{\mathcal{N}}(0,\sigma_{\tau}^{2})$, $\nu_{i}\sim{\tt Exponential}(\lambda)$, respectively. To connect with the notation $\bm{Y}_{i}$ introduced in Section 3, we have $\bm{Y}_{i}=(Y_{i},X_{i1},X_{i2},Z_{i1},Z_{i2})$ in this simulation setting. Both $Z_{i1}$ and $Z_{i2}$ are assumed to be fully observed but are assumed not intended to be adjusted in the analysis model. They are part of the simulation to mimic the scenario in which complete and low-cost external information can be used to produce ML predictions of possibly missing $(Y_{i},X_{i1},X_{i2})$ that are required in an analysis model of scientific interest. We set ${\mathcal{C}}_{i}=\infty$ to indicate that all $Y_{i},X_{i1},X_{i2}$ are observed. We introduce three distinct missingness patterns in addition to the fully observed data as follows: ${\mathcal{C}}_{i}=1$ indicates that $Y_{i}$ is missing; ${\mathcal{C}}_{i}=2$ indicates that only $X_{i2}$ is missing; ${\mathcal{C}}_{i}=3$ indicates that both $Y_{i}$ and $X_{i1}$ are missing. We assume a MAR mechanism. We define a propensity score model in the form proposed by Sun and Tchetgen Tchetgen, (2018) as follows:
$\textrm{logit}\{\pi({\mathcal{C}}_{i}|Z_{i},Y_{i,\text{obs}})\}=\begin{cases}-1+0.1X_{i2}+0.1Z_{i1}+0.1X_{i1}X_{i2},&{\mathcal{C}}_{i}=1\\ -1.8-0.2Y_{i}+0.1X_{i1}+0.1Z_{i1}+0.3X_{i1}Y_{i},&{\mathcal{C}}_{i}=2\\ -1.0+0.1X_{i2}+0.2Z_{i1},&{\mathcal{C}}_{i}=3\\ \end{cases},$ (9)
$\pi({\mathcal{C}}_{i}=\infty|Z_{i},Y_{i,\text{obs}})=1-\sum^{3}_{k=1}\pi({\mathcal{C}}_{i}=k|Z_{i},Y_{i,\text{obs}}).$ (10)
The average number of subjects under each missingness pattern in 500 simulations is $1,469.9$ for complete cases, and $1,350.7$, $829.5$, $1,349.9$ for ${\mathcal{C}}=1,2,3$, respectively.
To generate simulation data, we first sample $(Z_{i1},Z_{i2})$, then generate $(Y_{i},X_{i1},X_{i2})$ accordingly to obtain the full data. For each subject, we then sample ${\mathcal{C}}_{i}$ from the propensity score model and mask the variables based on the missingness pattern assigned to construct the observed dataset. To enable nimble control of bias and variance of the ML predictions for revealing the performance of the different methods under distinct qualities of ML predictions, instead of an explicit ML prediction, we design the ML predictions by adding additive noise and bias at a pre-specified level onto the unknown real values. More specifically,
$\widehat{X}_{i1}=X_{i1}+b_{i1}+\epsilon_{i1},$
$\widehat{X}_{i2}=X_{i2}+b_{i2}+\epsilon_{i2},$
$\widehat{Y}_{i}=Y_{i}+b_{i3}+\epsilon_{i3}.$
Here both $\epsilon_{i1}$, $\epsilon_{i2}$, and $\epsilon_{i3}$ follow Normal distribution ${\mathcal{N}}(0,\sigma^{2}_{\text{pred}})$, while $b_{i1}$, $b_{i2}$, and $b_{i3}$ follow Exponential distribution ${\tt Exponential}(\lambda_{\text{pred}})$. The two types of error terms represent the noise and bias imposed on the ML predictions, respectively. When these noises and bias are fixed constants 0, we have the perfect ML predictions.
In our simulation, we fix $\sigma=0.5$, $\sigma_{\tau}=0.3$, $\lambda=0.02$, $\sigma_{z}=0.2$, and $\rho=0.4$. To investigate the impact of the quality of the ML predictions on the performance of our proposed methods, we iterate both $\sigma^{2}_{\text{pred}}$ and $\lambda_{\text{pred}}$ from $\{0,0.2,0.4,0.6,0.8,1.0,1.5,2.0\}$. For propensity scores, we adopt three settings. In the first setting, we assume that the propensity score model is unknown yet correctly specified in terms of both the functional forms and covariates. Under this assumption, we estimate the propensity score by maximizing the likelihood function proposed in Sun and Tchetgen Tchetgen, (2018). In the second setting, we again assume an unknown propensity score model in the form proposed by Sun and Tchetgen Tchetgen, (2018), while we only include $Z_{1}$ in the propensity score model to make it misspecified. This setting is used to investigate the robustness of the PS-PPI procedure against propensity score model misspecification. In the final setting, we assume that the true propensity score is known and can be directly applied to estimate the parameters of interest $\bm{\beta}$}. We set the sample size $N=5,000$ and repeat the simulation for $500$ times.
4.2 Evaluation metrics
To evaluate the performance of included methods in synthetic data simulation, we compute the point estimates and report three key metrics: 1) the average estimated bias; 2) the coverage rates of the 95% confidence intervals (CIs); 3) the average width of these intervals. Specifically, suppose that $M$ replicates are generated in data simulation, and for the $j$-th replicate, let $\widehat{\beta}^{(j)}$ be the point estimate for a scalar parameter, and $(L^{(j)},U^{(j)})$ its 95% CI. We compute the average estimated bias as follows:
$\texttt{Avg. Estimated Bias}=\frac{1}{M}\sum^{M}_{j=1}\left(\widehat{\beta}^{(j)}-\beta_{\text{true}}\right).$
We compute the coverage rate as follows:
$\texttt{Coverage Rate}=\frac{1}{M}\sum^{M}_{j=1}I\left(\beta_{\text{true}}\in(L^{(j)},U^{(j)})\right).$
Finally, we compute the average width of the 95% CIs as follows:
$\texttt{Avg. 95\% CI Width}=\frac{1}{M}\sum^{M}_{j=1}\left(U^{(j)}-L^{(j)}\right).$
4.3 Results
Figure 2: Widths of 95% confidence intervals, coverage rates, and biases for different methods under the assumption of unknown propensity scores while with estimated propensity scores from correctly specified propensity score models. The X-axis in all panels represents the prediction noise level used in the simulations. The bias level is fixed at $\lambda_{\text{pred}}=0$. For Y-axis: in (a), shows the $\log_{10}$ of the widths of the 95% confidence intervals; in (b), indicates the coverage rate of the 95% confidence intervals; in (c), shows the bias of the estimated parameters. In (b), the red dashed line represents the predefined confidence level 0.95.
We present the results with varying noise levels and fixed bias level ($\lambda_{\text{pred}}=0$), and estimated propensity scores in Figure 2. Additional experimental results on varying prediction bias, as well as different settings of the propensity score models are provided in Figure S1S2S3S4, and S5, in Section S2.1 of the Supplementary Materials. Note that the coverage rate, the width of the 95% CIs, and the bias remain unchanged for CCA, WCCA, and MI when varying the quality of the ML predictions, as they do not apply these predictions in their estimating procedures.
Coverage rate
In terms of the coverage rate of the 95% CIs, the CCA approach achieves a coverage rate of 51.0% and 81.6% for $\beta_{0}$ and $\beta_{1}$, respectively, which is below the nominal coverage rate of 95%. This is expected because it does not incorporate propensity scores to adjust for the missing-at-random mechanism that renders the complete-cases not representative of the entire population. WCCA achieves conservative coverage rates of 98.2%, 98.0%, and 96.2% on the three coefficients, respectively. The conservativeness is due to the use of the robust sandwich variance that does not account for the uncertainty from the estimated propensity score model. From Figure S2, where all the settings are the same except that the propensity score model is known, WCCA exhibits valid coverages around 95% for all the three coefficients. MI achieves permissive coverage rates of 93.2%, 91.6% and 56.6%, respectively for the three parameters, falling short of the nominal 95% level. Under perfect predictions, PPI++ achieves coverage rates of 71.6% and 89.2% for $\beta_{0}$ and $\beta_{1}$ due to the fact that it does not use the propensity score, similar to the CCA approach. When the quality of the ML predictions degrades, the coverage rates of these two coefficients decrease and converge to levels similar to the CCA approach. PS-PPI, on the other hand, achieves a coverage rate of around 95% when the perfect predictions are available. When the quality of the ML predictions decreases, it tends to provide a conservative confidence intervals, similar to its base estimator, i.e., the WCCA estimator.
Widths
The widths of the 95% CIs for all the three coefficients obtained from CCA are 0.027, 0.085, and 0.129, respectively; the widths of the 95% CIs for all the three coefficients obtained from WCCA are 0.028, 0.087, and 0.130, respectively. For both CCA and WCCA, the width of the 95% CIs are similar for all the three coefficients. MI has 95% CIs of width 0.025, 0.077, and 0.130 on all the three coefficients (however the real coverage rates for $\beta_{1}$ and $\beta_{2}$ are unsatisfactory). PS-PPI yields uniformly narrower 95% CIs compared to WCCA, the only method with valid coverage rates for all the three coefficients. Furthermore, the widths of 95% CIs obtained from PS-PPI is strongly impacted by the quality of the ML predictions. Specifically, PS-PPI tends to produce narrower 95% CIs when ML predictions have better quality. As a comparison, the PPI method shows similar patterns in terms of the relationship of ML predictions quality and the width of 95% CIs. Furthermore, PS-PPI produces narrower 95% CIs than PPI in most of the cases, except the 95% CI of $\beta_{2}$ when the quality of ML predictions becomes mediocre.
Bias
From the bias evaluation, it can be seen that CCA, PPI++, and MI’s failure in achieving 95% nominal coverage are primarily due to its biased estimation. For the bias in the estimation of $\beta_{0}$ and $\beta_{1}$, WCCA results in a simulated average bias of 0.0005 and 0.0006, respectively. On the other hand, PS-PPI results in 0.0007 and 0.0008 with the worst quality of ML predictions. Nevertheless, for the parameter where CCA, PPI++ and MI fails to achieve 95% nominal coverage, they exhibit greater simulated bias. Specifically, MI exhibits a simulated average bias of 0.115 for the estimation of $\beta_{2}$, while CCA results in around 0.043 of simulated bias for $\beta_{1}$. In terms of PPI, the bias increases from 0.023 to 0.043 when the quality of ML predictions degrades.
Similar results are also observed from Figure S1S2S3. The simulation results demonstrate that under non-monotone missingness patterns with MAR assumption for the missing mechanism, the proposed PS-PPI approach can produce consistent estimation and valid inference with efficiency no worse than WCCA when the propensity score model is correctly estimated or known. Such consistency and validity in inference are irrelevant to the quality of the ML predictions. Furthermore, when precise ML predictions are available, PS-PPI achieves greater estimation efficiency compared to WCCA. This corroborates our theoretical results.
Robustness against propensity score model misspecification
The simulation results under misspecified propensity score models are shown in Figure S4 and S5. Note that WCCA and PS-PPI are the only two methods that are impacted by the propensity score model specification. When propensity score model is misspecified in terms of the covariates, all the methods show a degree of performance in the width of the 95% CIs similar to the one shown in Figure 2. In terms of the coverage rate, Figure S4 and S5 show that WCCA fails to achieve the nominal 95% coverage rate for $\beta_{0}$ and $\beta_{1}$. PS-PPI achieve the 95% coverage rate when the ML predictions are sufficiently precise, while these coverage rates decline when the quality of ML predictions degrades. However, it can be seen that the coverage rates from PS-PPI still consistently outperform the WCCA baseline. Furthermore, PS-PPI also maintains a higher coverage rate compared to the PPI method, which only utilizes the fully-observed and missing-outcome subsets of data.
5 Conclusion and Discussion
This paper introduces a unified framework for conducting valid statistical inference using machine learning imputations in datasets characterized by general, multiple different missing data patterns under the MAR assumption. By leveraging the concept of missingness patterns and building upon the theoretical foundations of AIPW estimators, our work significantly extends the applicability of prediction-based inference beyond the more restrictive MCAR assumption and simple missing data structures, such as missing outcomes or blockwise missing covariates. A key practical advantage of our proposed methods is their adaptability to any Z-estimation framework, allowing researchers to utilize existing statistical software for weighted complete-case analysis with minimal additional implementation effort.
The significance of this development lies in the theoretical foundation it provides to potentially unlock more robust and reliable insights from the vast and often messy datasets prevalent in modern scientific inquiry. As machine learning models become increasingly sophisticated in their predictive capabilities, their use for imputation is a natural and appealing step. However, without rigorous inferential correction, analyses based on such imputed data risk biased conclusions. PS-PPI provides a principled approach to harness the power of ML predictions while maintaining statistical validity, which is crucial in fields ranging from biomedicine and public health to economics and social sciences, where missing data are ubiquitous and decisions based on data analysis can have profound consequences. We provide a theoretical guarantee of improved efficiency over weighted complete-case analysis, making it an attractive option especially when ML predictions are of reasonable quality.
Despite these important contributions, our work has several limitations that inform avenues for future research. First, a central assumption of our methodology is the correct specification of the propensity score model, which governs the probability of observing different missingness patterns. This assumption is the price we pay in the challenging scenario where prediction-based inference is most needed: analysts typically have minimal or no control of the prediction models. As such, assumptions on the prediction models need to be avoided, and the correct specification of the propensity score model comes as the price. Although we demonstrate valid inference when propensity scores are known or consistently estimated from a correctly specified model, results from Figure S4 and S5 show that propensity score model misspecification will lead to biased estimation when the ML predictions are not sufficiently precise. In practice, however, correctly specifying the propensity score model, especially when the number of missingness patterns $K$ is large, is challenging if not impossible. As a result, although this paper provide a theoretical foundation for an arbitrary $K$, it is more practically feasible when $K$ is small. Future work should focus on robust propensity score model against misspecification. Furthermore, it is worth noting that results from Figure S4 and S5 exhibits empirical evidence of the double robustness of the proposed PS-PPI method, where PS-PPI achieve the 95% nominal coverage rate when the quality of the ML predictions are precise enough (or perfect). Recalling that the PS-PPI estimator is an AIPW estimator, a thorough theoretical investigation to the potential double robustness property of the PS-PPI estimator is needed as a future work to clarify the quality requirement of the ML predictions to bypass the challenging modeling process of the propensity score models.
Second, another practical limitation of PS-PPI is that it imposes assumptions to the minimum size of data within each missingness patterns to guarantee the learnability of the parameter of interest. For example, to fit a linear regression analysis model with $p$ parameters (including the intercept), at least $p+1$ data points are needed. In this case, PS-PPI requires each missingness pattern within the data to contain at least $p+1$ data points, and if such requirement is not satisfied, then data with that missingness pattern have to be dropped from the dataset.
Third, while PS-PPI provides a data-driven approach for selecting tuning parameters ($\mathbf{W}_{k}$) that guarantees efficiency gains, the assumption of identical $\mathbf{W}_{k}$ across strata was made for simplicity and ease of implementation. Future work could explore more adaptive stratum-specific weight selection to potentially achieve even greater efficiency, although this may involve more complex optimization procedures. Furthermore, the number of possible missingness patterns can grow exponentially with the number of variables. Although in many practical applications, the number of observed patterns is manageable, scenarios with a very large number of distinct patterns could pose computational challenges for the stratification and subsequent analyses. Research into methods for adaptively grouping or regularizing coarsening patterns could be beneficial.
Fourth, while our framework is versatile within Z-estimation problems, its extension to other inferential paradigms or specific complex data types, such as longitudinal or spatial data where the temporal and spatial dependency of missingness is critical, or high-dimensional settings where variable selection and inference are performed simultaneously, presents fertile ground for future research. The current work also focuses on the MAR assumption. Extending these prediction-powered inference ideas to handle Missing Not At Random (MNAR) scenarios, while notoriously challenging due to identifiability issues, would be a significant theoretical and practical advancement.
The broader implications of this research are multifold. By providing a more general and statistically sound framework for using ML-imputed data, PS-PPI can encourage the adoption of advanced predictive modeling in substantive research areas without sacrificing inferential rigor. This can lead to more powerful and nuanced analyses of existing datasets, potentially uncovering new scientific insights that were previously obscured by missing information or by the limitations of traditional missing data techniques. Moreover, the ability to validly incorporate ML predictions could influence how researchers approach data collection, potentially allowing for more targeted and cost-effective acquisition of complete data on subsets of variables, knowing that robust inferential methods exist to handle the resulting missingness.
Supplementary Materials
We first provide a glossary for key notation used in the Main Paper.
Symbol Description
$\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ The PS-PPI estimator proposed in this work.
$\widehat{{\bm{\theta}}}_{\text{WCC}}$ An estimator from complete-case analysis with inverse probability weighting (IPW).
$\widehat{{\bm{\theta}}}_{\text{PP}}$ An estimator from outcome-based prediction-powered inference (Angelopoulos et al., 2023a, ).
$\widehat{{\bm{\theta}}}_{\text{CC}}$ An estimator from unweighted complete-case analysis.
$\widehat{{\bm{\theta}}}_{\text{Chen}}$ An outcome-based PPI++ style estimator from Gronsbell et al., (2024); Chen and Chen, (2000) that guarantees to be more efficient than $\widehat{{\bm{\theta}}}_{\text{CC}}$.
$\widehat{{\bm{\theta}}}_{\text{PTD}}$ An estimator from Kluger et al., (2025) for blockwise covariates missing with nonuniform sampling weights.
$\widehat{\mathbf{\Delta}}_{k}$ Estimates of zero obtained from the $k$-th missingness patterns; $\widehat{\mathbf{\Delta}}=\widehat{{\bm{\gamma}}}_{1,k}-\widehat{{\bm{\gamma}}}_{2,k}$.
$\widehat{{\bm{\gamma}}}_{1,k}$ The solution of Equation (5), which use complete-case data, with all entries supposed to be missing in the $k$-th missingness patterns being substituted with predictions generated from $h_{k}$.
$\widehat{{\bm{\gamma}}}_{2,k}$ The solution of Equation (6), which use data with $k$-th missingness patterns, with all missing values imputed by predictions generated from $h_{k}$.
$\widehat{\mathbf{W}}_{k}$ Weight matrices of dimension $d\times d$ that may depend on data. In this work, Corollary 3.2 is used to determine $\widehat{\mathbf{W}}_{k}$.
$\widehat{\Sigma}_{{\bm{\theta}}}$ The asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{WCC}}$.
$\widehat{\Sigma}_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$ The asymptotic covariance between $\widehat{{\bm{\theta}}}_{\text{WCC}}$ and $\widehat{{\bm{\gamma}}}_{1,k}$.
$\widehat{\Sigma}_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$ The asymptotic covariance between $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{1,k^{\prime}}$, $k,k^{\prime}\in\{1,...,K\}$. When $k=k^{\prime}$, it becomes the asymptotic variance of $\widehat{{\bm{\gamma}}}_{1,k}$.
$\widehat{\Sigma}_{{\bm{\gamma}}_{2,k}}$ The asymptotic variance of $\widehat{{\bm{\gamma}}}_{2,k}$.
$\widehat{\Sigma}_{\text{PS-PPI}}$ The asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$.
Appendix S1 Proofs
S1.1 Proof of Proposition 1
Proof.
Under the $k$-th missingness pattern, from Equation (5) and Assumption 1 we have:
$\;\mathbb{E}_{(\bm{Y},{\mathcal{C}})}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}\left(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}}^{\ast}_{k}\right)\right]$
$=\mathbb{E}_{\bm{Y}}\left[{\bm{\psi}}\left(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}}^{\ast}_{k}\right)\mathbb{E}_{({\mathcal{C}}|\bm{Y})}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}\right]\right]$
$=\mathbb{E}_{\bm{Y}}\left[{\bm{\psi}}\left(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}}^{\ast}_{k}\right)\right]=\bm{0}.$
Similarly, from Equation (6) we have:
$\quad\;\mathbb{E}_{(\bm{Y},{\mathcal{C}})}\left[\frac{I({\mathcal{C}}_{i}=k)}{\pi_{j}(G_{k}(\bm{Y}_{i}))}{\bm{\psi}}\left(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}}^{\ast}_{k}\right)\right]$
$=\mathbb{E}_{\bm{Y}}\left[{\bm{\psi}}\left(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}}^{\ast}_{k}\right)\mathbb{E}_{({\mathcal{C}}|\bm{Y})}\left[\frac{I({\mathcal{C}}_{i}=k)}{\pi_{j}(G_{k}(\bm{Y}_{i}))}\right]\right]$
$=\mathbb{E}_{\bm{Y}}\left[{\bm{\psi}}\left(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}}^{\ast}_{k}\right)\right]=\bm{0}.$
S1.2 Proof of Theorem 1
We first stack the estimating equations and obtain:
$\bm{0}=\sum^{N}_{i=1}\bm{\Psi}^{(i)}=\begin{bmatrix}\sum^{N}_{i=1}{\bm{\psi}}_{\text{WCC}}^{(i)}\\[6.0pt] \sum^{N}_{i=1}{\bm{\psi}}_{1,1}^{(i)}\\[2.0pt] \vdots\\[2.0pt] \sum^{N}_{i=1}{\bm{\psi}}^{(i)}_{1,K}\\[6.0pt] \sum^{N}_{i=1}{\bm{\psi}}^{(i)}_{2,1}\\[2.0pt] \vdots\\[2.0pt] \sum^{N}_{i=1}{\bm{\psi}}^{(i)}_{2,K}\\[6.0pt] \sum^{N}_{i=1}{\bm{m}}^{(i)}\end{bmatrix}$.
[…]
Assumption 2.
We assume that $\bm{\Psi}$ is measurable vector-valued function and for every $\bm{\vartheta}_{1}$ and $\bm{\vartheta}_{2}$ in a neighborhood of $\bm{\vartheta}^{\ast}$,
[…]
Proof.
By working on the stacked estimating equations we have the following asymptotic linear expansion:
$\sqrt{N}(\widehat{\bm{\vartheta}}-\bm{\vartheta}^{\ast})=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}-\dot{\bm{\Psi}}^{-1}\bm{\Psi}^{(i)}+o_{p}(1),$
where
$\dot{\bm{\Psi}}=\mathbb{E}\Bigl{[}\tfrac{\partial\bm{\Psi}}{\partial\bm{\vartheta}}\Bigr{|}_{\bm{\vartheta}=\bm{\vartheta}^{\ast}}\Bigr{]}=\mathbb{E}\!\left[\,\begin{bmatrix}\frac{\partial{\bm{\psi}}_{\text{WCC}}}{\partial{\bm{\theta}}}&\frac{\partial{\bm{\psi}}_{\text{WCC}}}{\partial{\bm{\gamma}}_{1,1}}&\cdots&\frac{\partial{\bm{\psi}}_{\text{WCC}}}{\partial{\bm{\gamma}}_{1,K}}&\frac{\partial{\bm{\psi}}_{\text{WCC}}}{\partial{\bm{\gamma}}_{2,1}}&\cdots&\frac{\partial{\bm{\psi}}_{\text{WCC}}}{\partial{\bm{\gamma}}_{2,K}}&\frac{\partial{\bm{\psi}}_{\text{WCC}}}{\partial{\bm{\omega}}}\\[6.0pt] \frac{\partial{\bm{\psi}}_{1,1}}{\partial{\bm{\theta}}}&\frac{\partial{\bm{\psi}}_{1,1}}{\partial{\bm{\gamma}}_{1,1}}&\cdots&\frac{\partial{\bm{\psi}}_{1,1}}{\partial{\bm{\gamma}}_{1,K}}&\frac{\partial{\bm{\psi}}_{1,1}}{\partial{\bm{\gamma}}_{2,1}}&\cdots&\frac{\partial{\bm{\psi}}_{1,1}}{\partial{\bm{\gamma}}_{2,K}}&\frac{\partial{\bm{\psi}}_{1,1}}{\partial{\bm{\omega}}}\\[2.0pt] \vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots\\[2.0pt] \frac{\partial{\bm{m}}}{\partial{\bm{\theta}}}&\frac{\partial{\bm{m}}}{\partial{\bm{\gamma}}_{1,1}}&\cdots&\frac{\partial{\bm{m}}}{\partial{\bm{\gamma}}_{1,K}}&\frac{\partial{\bm{m}}}{\partial{\bm{\gamma}}_{2,1}}&\cdots&\frac{\partial{\bm{m}}}{\partial{\bm{\gamma}}_{2,K}}&\frac{\partial{\bm{m}}}{\partial{\bm{\omega}}}\end{bmatrix}_{\bm{\vartheta}=\bm{\vartheta}^{\ast}}\,\right].$
From van der Vaart, (Theorem 5.21 2000) and Assumption 2, $\sqrt{N}(\widehat{\bm{\vartheta}}-\bm{\vartheta}^{\ast})$ is asymptotically normal. To show that $\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})$ is asymptotically normal, we first introduce the following notation:
$\dot{{\bm{\psi}}}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}_{\text{WCC}}}{\partial{\bm{\theta}}}\Bigr{|}_{{\bm{\theta}}={\bm{\theta}}^{\ast}}\right],\leavevmode\nobreak\ \dot{{\bm{\psi}}}_{k}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}_{1,k}}{\partial{\bm{\gamma}}_{1,k}}\Bigr{|}_{{\bm{\gamma}}_{1,k}={\bm{\gamma}}^{\ast}_{k}}\right]=\mathbb{E}\left[\frac{\partial{\bm{\psi}}_{2,k}}{\partial{\bm{\gamma}}_{2,k}}\Bigr{|}_{{\bm{\gamma}}_{2,k}={\bm{\gamma}}^{\ast}_{k}}\right],$
$\dot{{\bm{\psi}}}^{\bm{\omega}}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}_{\text{WCC}}}{\partial{\bm{\omega}}}\Bigr{|}_{{\bm{\omega}}={\bm{\omega}}^{\ast}}\right],\leavevmode\nobreak\ \dot{{\bm{\psi}}}^{\bm{\omega}}_{1,k}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}_{1,k}}{\partial{\bm{\omega}}}\Bigr{|}_{{\bm{\omega}}={\bm{\omega}}^{\ast}}\right],\leavevmode\nobreak\ \dot{{\bm{\psi}}}^{\bm{\omega}}_{2,k}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}_{2,k}}{\partial{\bm{\omega}}}\Bigr{|}_{{\bm{\omega}}={\bm{\omega}}^{\ast}}\right],\leavevmode\nobreak\ \mathbf{M}=\mathbb{E}\left[\frac{\partial{\bm{m}}}{\partial{\bm{\omega}}}\Bigr{|}_{{\bm{\omega}}={\bm{\omega}}^{\ast}}\right].$
Then, by noticing that all the non-diagonal elements except those in the last row are $\bm{0}$, we have:
$\dot{\bm{\Psi}}=\begin{bmatrix}\dot{{\bm{\psi}}}&\bm{0}&\cdots&\dot{{\bm{\psi}}}^{{\bm{\omega}}}\\ \bm{0}&\dot{{\bm{\psi}}}_{1}&\cdots&\dot{{\bm{\psi}}}^{{\bm{\omega}}}_{1,1}\\ \vdots&\vdots&\ddots&\vdots\\ \bm{0}&\bm{0}&\cdots&\mathbf{M}\end{bmatrix}=\begin{bmatrix}\mathbf{J}_{1}&\mathbf{J}_{2}\\ \bm{0}&\mathbf{M}\end{bmatrix}$
where $\mathbf{J}_{1}=\text{diag}(\dot{{\bm{\psi}}},\dot{{\bm{\psi}}}_{1},\cdots,\dot{{\bm{\psi}}}_{K},\dot{{\bm{\psi}}}_{1},\cdots\dot{{\bm{\psi}}}_{K})$, $\mathbf{J}_{2}=(\dot{{\bm{\psi}}}^{\bm{\omega}},\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,1},\cdots,\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,K},\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,1},\cdots\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,K})^{\top}$. Hence, from Lu and Shiou, (2002) we can obtain the inversion of the upper triangle $2\times 2$ block matrix $\dot{\bm{\Psi}}$:
$\dot{\bm{\Psi}}^{-1}=\begin{bmatrix}\dot{{\bm{\psi}}}^{-1}&\bm{0}&\cdots&-\dot{{\bm{\psi}}}^{-1}\dot{{\bm{\psi}}}^{{\bm{\omega}}}\mathbf{M}^{-1}\\[6.0pt] \bm{0}&\dot{{\bm{\psi}}}_{1}^{-1}&\cdots&-\dot{{\bm{\psi}}}_{1}^{-1}\dot{{\bm{\psi}}}_{1,1}^{{\bm{\omega}}}\mathbf{M}^{-1}\\[2.0pt] \vdots&\vdots&\ddots&\vdots\\[2.0pt] \bm{0}&\bm{0}&\cdots&\mathbf{M}^{-1}\end{bmatrix}$
On the other hand, we know that $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}=\widehat{{\bm{\theta}}}_{\text{WCC}}-\sum^{K}_{k=1}\widehat{\mathbf{W}}_{k}(\widehat{{\bm{\gamma}}}_{1,k}-\widehat{{\bm{\gamma}}}_{2,k})=\widehat{{\bm{c}}}^{\top}\widehat{\bm{\vartheta}}$, where $\widehat{{\bm{c}}}=(\mathbf{I},-\widehat{\mathbf{W}}_{1},\cdots,-\widehat{\mathbf{W}}_{K},\widehat{\mathbf{W}}_{1},\cdots,\widehat{\mathbf{W}}_{K},\bm{0})^{\top}$. Since $\widehat{\mathbf{W}}_{k}\overset{p}{\rightarrow}\mathbf{W}_{k}$, we have $\widehat{{\bm{c}}}\overset{p}{\rightarrow}{\bm{c}}$, where ${\bm{c}}=(\mathbf{I},-\mathbf{W}_{1},\cdots,-\mathbf{W}_{K},\mathbf{W}_{1},\cdots,\mathbf{W}_{K},\bm{0})^{\top}$. Based on Proposition 1, we have both ${\bm{\theta}}^{\ast}=\widehat{{\bm{c}}}^{\top}\bm{\vartheta}^{\ast}$ and ${\bm{\theta}}^{\ast}={\bm{c}}^{\top}\bm{\vartheta}^{\ast}$. Therefore, we can obtain the asymptotic linear expansion of $\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})$:
$\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})=\widehat{{\bm{c}}}^{\top}\sqrt{N}(\widehat{\bm{\vartheta}}-\bm{\vartheta}^{\ast})=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}-{\bm{c}}^{\top}\dot{\bm{\Psi}}^{-1}\bm{\Psi}^{(i)}+o_{p}(1),$
where
${\bm{c}}^{\top}\dot{\bm{\Psi}}^{-1}=\begin{bmatrix}\dot{{\bm{\psi}}}^{-1}\\[6.0pt] -\mathbf{W}_{1}\dot{{\bm{\psi}}}^{-1}_{1}\\[6.0pt] \vdots\\ -\mathbf{W}_{1}\dot{{\bm{\psi}}}^{-1}_{K}\\[6.0pt] \mathbf{W}_{1}\dot{{\bm{\psi}}}^{-1}_{1}\\[6.0pt] \cdots\\ \mathbf{W}_{K}\dot{{\bm{\psi}}}^{-1}_{K}\\[6.0pt] \left(\dot{{\bm{\psi}}}^{-1}\dot{{\bm{\psi}}}^{\bm{\omega}}-\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}(\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,k}-\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,k})\right)\mathbf{M}^{-1}\end{bmatrix}^{\top}$
As a linear transformation of $\sqrt{N}(\widehat{\bm{\vartheta}}-\bm{\vartheta}^{\ast})$, we know that $\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})$ is asymptotically normal. We now derive the form of its asymptotic variance. Denote ${\bm{\psi}}^{(i)}={\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}})$, and ${\bm{\psi}}^{(i)}_{k}={\bm{\psi}}(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}})$; since $\widehat{\mathbf{W}}_{k}\overset{p}{\rightarrow}\mathbf{W}_{k}$, we have:
$\;\;\;\;\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})$
$=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}\left[-\dot{{\bm{\psi}}}^{-1}{\bm{\psi}}^{(i)}_{\text{WCC}}+\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}({\bm{\psi}}^{(i)}_{1,k}-{\bm{\psi}}^{(i)}_{2,k})\right$.
$\;\;\;+\left.\left(\dot{{\bm{\psi}}}^{-1}\dot{{\bm{\psi}}}^{\bm{\omega}}-\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}(\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,k}-\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,k})\right)\mathbf{M}^{-1}{\bm{m}}^{(i)}\right]+o_{p}(1)$
$=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}\left[\underbrace{-\dot{{\bm{\psi}}}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}_{\infty}(\bm{Y}_{i})}{\bm{\psi}}^{(i)}+\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}_{\infty}(\bm{Y}_{i})}-\frac{I({\mathcal{C}}_{i}=k)}{\widehat{\pi}_{k}(G_{k}(\bm{Y}_{i}))}\right){\bm{\psi}}^{(i)}_{k}}_{{\bm{\Gamma}}_{i}}\right$.
$\;\;\;+\left.\left(\dot{{\bm{\psi}}}^{-1}\dot{{\bm{\psi}}}^{\bm{\omega}}-\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}(\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,k}-\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,k})\right)\mathbf{M}^{-1}{\bm{m}}^{(i)}\right]+o_{p}(1).$
Notice that,
$\dot{{\bm{\Gamma}}}_{\bm{\omega}}:=\mathbb{E}\left[\frac{\partial{\bm{\Gamma}}_{i}}{\partial{\bm{\omega}}}\Bigr{|}_{{\bm{\omega}}={\bm{\omega}}^{\ast}}\right]=\dot{{\bm{\psi}}}^{-1}\dot{{\bm{\psi}}}^{\bm{\omega}}-\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}(\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,k}-\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,k}).$
As a result,
$\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}\left({\bm{\Gamma}}_{i}+\dot{{\bm{\Gamma}}}_{\bm{\omega}}\mathbf{M}^{-1}{\bm{m}}_{i}\right)+o_{p}(1),$
and therefore the $i$-th influence function of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is
$\varphi_{i}={\bm{\Gamma}}_{i}+\dot{{\bm{\Gamma}}}_{\bm{\omega}}\mathbf{M}^{-1}{\bm{m}}_{i}.$
Accordingly, $\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}})\overset{d}{\rightarrow}{\mathcal{N}}(0,\text{Var}(\varphi_{i}))$.
$\text{Var}(\varphi_{i})=\mathbb{E}\left[\left({\bm{\Gamma}}_{i}+\dot{{\bm{\Gamma}}}_{\bm{\omega}}\mathbf{M}^{-1}{\bm{m}}_{i}\right)\left({\bm{\Gamma}}_{i}+\dot{{\bm{\Gamma}}}_{\bm{\omega}}\mathbf{M}^{-1}{\bm{m}}_{i}\right)^{\top}\right]$ (S1.1)
S1.3 Proof of Corollary 3.1
Proof.
[…]
${\bm{\Gamma}}_{i}=-\underbrace{\dot{{\bm{\psi}}}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}^{(i)}}_{\mathbf{A}}+\underbrace{\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}-\frac{I({\mathcal{C}}_{i}=k)}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\right){\bm{\psi}}^{(i)}_{k}}_{\mathbf{B}}.$
Correspondingly, the asymptotic variance becomes $\text{Var}({\bm{\Gamma}}_{i})=\mathbb{E}\left[{\bm{\Gamma}}_{i}{\bm{\Gamma}}^{\top}_{i}\right]=\mathbb{E}[\mathbf{A}\mathbf{A}^{\top}-\mathbf{A}\mathbf{B}^{\top}-\mathbf{B}\mathbf{A}^{\top}+\mathbf{B}\mathbf{B}^{\top}]$. Here,
$\mathbf{A}\mathbf{A}^{\top}$ $=\dot{{\bm{\psi}}}^{-1}{\bm{\psi}}{\bm{\psi}}^{\top}\dot{{\bm{\psi}}}^{-1\top}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi^{2}_{\infty}(\bm{Y}_{i})},$
$\mathbf{A}\mathbf{B}^{\top}$ $=\sum^{K}_{k=1}\dot{{\bm{\psi}}}^{-1}{\bm{\psi}}{\bm{\psi}}^{\top}_{k}\dot{{\bm{\psi}}}^{-1\top}_{k}\mathbf{W}^{\top}_{k}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}-\frac{I({\mathcal{C}}_{i}=k)}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\right)$
$=\sum^{K}_{k=1}\dot{{\bm{\psi}}}^{-1}{\bm{\psi}}{\bm{\psi}}^{\top}_{k}\dot{{\bm{\psi}}}^{-1\top}_{k}\mathbf{W}^{\top}_{k}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi^{2}_{\infty}(\bm{Y}_{i})},$
$\mathbf{B}\mathbf{B}^{\top}$ $=\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}{\bm{\psi}}_{k}{\bm{\psi}}^{\top}_{k^{\prime}}\dot{{\bm{\psi}}}^{-1\top}_{k^{\prime}}\mathbf{W}^{\top}_{k^{\prime}}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi^{2}_{\infty}(\bm{Y}_{i})}+\frac{I({\mathcal{C}}_{i}=k)I({\mathcal{C}}_{i}=k^{\prime})}{\pi_{k}(G_{k}(\bm{Y}_{i}))\pi_{k^{\prime}}(G_{k^{\prime}}(\bm{Y}_{i}))}\right)$
Therefore,
$\mathbb{E}\left[\mathbf{A}\mathbf{A}^{\top}\right]$ $=\underbrace{\dot{{\bm{\psi}}}^{-1}\mathbb{E}\left[\frac{{\bm{\psi}}{\bm{\psi}}^{\top}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}^{-1\top}}_{\Sigma_{{\bm{\theta}}}},$ (S1.2)
$\mathbb{E}\left[\mathbf{A}\mathbf{B}^{\top}\right]$ $=\sum^{K}_{k=1}\underbrace{\dot{{\bm{\psi}}}^{-1}\mathbb{E}\left[\frac{{\bm{\psi}}{\bm{\psi}}^{\top}_{k}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}^{-1\top}_{k}}_{\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}}\mathbf{W}^{\top}_{k},$ (S1.3)
$\mathbb{E}\left[\mathbf{B}\mathbf{B}^{\top}\right]$ $=\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\mathbf{W}_{k}\underbrace{\dot{{\bm{\psi}}}^{-1}_{k}\mathbb{E}\left[\frac{{\bm{\psi}}_{k}{\bm{\psi}}^{\top}_{k^{\prime}}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}^{-1\top}_{k^{\prime}}}_{\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}}\mathbf{W}^{\top}_{k^{\prime}}$ (S1.4)
$+\sum^{K}_{k=1}\mathbf{W}_{k}\underbrace{\dot{{\bm{\psi}}}^{-1}_{k}\mathbb{E}\left[\frac{{\bm{\psi}}_{k}{\bm{\psi}}^{\top}_{k}}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\right]\dot{{\bm{\psi}}}^{-1\top}_{k}}_{\Sigma_{{\bm{\gamma}}_{2,k}}}\mathbf{W}^{\top}_{k}.$ (S1.5)
From the asymptotic linear expansion of $\widehat{{\bm{\theta}}}_{\text{WCC}}$, $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\theta}}}_{2,k}$, we know that $\Sigma_{{\bm{\theta}}}$ is the asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{WCC}}$; $\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$ is the asymptotic covariance between $\widehat{{\bm{\theta}}}_{\text{WCC}}$ and $\widehat{{\bm{\gamma}}}_{1,k}$; $\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$ is the asymptotic covariance between $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{1,k^{\prime}}$; $\Sigma_{{\bm{\gamma}}_{2,k}}$ is the asymptotic variance of $\widehat{{\bm{\theta}}}_{2,k}$. As a results, the asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is:
$\Sigma_{\text{PS-PPI}}=\Sigma_{{\bm{\theta}}}-\sum^{K}_{k=1}\mathbf{W}_{k}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}-\sum^{K}_{k=1}(\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}})^{\top}\mathbf{W}_{k}^{\top}+\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\mathbf{W}_{k}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\mathbf{W}_{k^{\prime}}^{\top}+\sum^{K}_{k=1}\mathbf{W}_{k}\Sigma_{{\bm{\gamma}}_{2,k}}\mathbf{W}^{\top}_{k}.$ (S1.6)
S1.4 Instruction of estimating $\Sigma_{\text{PS-PPI}}$.
To obtain a consistent estimate of $\Sigma_{\text{PS-PPI}}$, we may want to obtain consistent estimates for both the meat (e.g., $\mathbb{E}\left[\frac{{\bm{\psi}}{\bm{\psi}}^{\top}}{\pi_{\infty}(\bm{Y}_{i})}\right]$ for $\Sigma_{\bm{\theta}}$) and the bread (e.g., $\dot{{\bm{\psi}}}^{-1}$ and $\dot{{\bm{\psi}}}^{-1\top}$ for $\Sigma_{\bm{\theta}}$) parts for each of the variance-covariance matrices $\Sigma_{{\bm{\theta}}}$, $\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$, $\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$, and $\Sigma_{{\bm{\gamma}}_{2,k}}$. A straightforward approach is to use the sample average over the fully-observed data to estimate the population average shown in both the meat and the bread parts. This can be done once the closed-form of $\bm{\psi}$${\bm{\psi}}_{k}$ and $\dot{{\bm{\psi}}}$$\dot{{\bm{\psi}}}_{k}$ can be derived. In the following we provide two examples, linear regression and logistic regression, for reference.
S1.4.1 Linear Regression
Under OLS, the $\bm{\psi}$} of the linear regression for each subject:
${\bm{\psi}}^{(i)}=X_{i}(y_{i}-X_{i}^{\top}{\bm{\theta}}),$
and the expected first-order derivative of ${\bm{\psi}}^{(i)}$:
$\dot{{\bm{\psi}}}=\mathbb{E}\left[X_{i}X_{i}^{\top}\right].$
Similarly we can obtain ${\bm{\psi}}^{(i)}_{k}$ and $\dot{{\bm{\psi}}}_{k}$ by substituting the $y_{i}$ and $X_{i}$ with their imputed counterpart under the $k$-th missingness pattern. As such, taking $\Sigma_{\bm{\theta}}$ for an example, denoting $N_{\textrm{CC}}=\sum^{N}_{i=1}I({\mathcal{C}}_{i}=\infty)$, we can compute the consistent estimate of the meat:
$\widehat{\mathbb{E}}\left[\frac{{\bm{\psi}}{\bm{\psi}}^{\top}}{\pi_{\infty}(\bm{Y}_{i})}\right]=\frac{1}{N_{\textrm{CC}}}\sum^{N}_{i=1}I({\mathcal{C}}_{i}=\infty)\frac{X_{i}(y_{i}-X_{i}^{\top}\widehat{{\bm{\theta}}}_{\text{WCC}})^{2}X^{\top}_{i}}{\widehat{\pi}_{\infty}(X_{i},y_{i})}.$
For the bread:
$\widehat{\dot{{\bm{\psi}}}}=\frac{1}{N_{\textrm{CC}}}\sum^{N}_{i=1}I({\mathcal{C}}_{i}=\infty)X_{i}X_{i}^{\top}.$
Note that $\widehat{{\bm{\theta}}}_{\text{WCC}}$ is plugged in to compute ${\bm{\psi}}^{(i)}$. For the corresponding meat and bread parts in all the other variance-covariance matrices, we could apply the similar procedure to obtain their consistent estimates. When computing ${\bm{\psi}}_{k}$ and $\dot{{\bm{\psi}}}_{k}$, $\widehat{{\bm{\gamma}}}_{1,k}$ will be plugged in into the estimating equations.
S1.4.2 Logistic Regression
In the logistic regression, the procedure for obtaining the consistent estimates of the variance-covariance matrices are similar; the only differences are the specific form of ${\bm{\psi}}^{(i)}$, ${\bm{\psi}}^{(i)}_{k}$, $\dot{{\bm{\psi}}}$, and $\dot{{\bm{\psi}}}_{k}$. Specifically, under the logistic regression setting, the log-likelihood for each subject is:
$\ell_{i}({\bm{\theta}})=y_{i}X_{i}^{\top}{\bm{\theta}}-\log\left(1+e^{X_{i}^{\top}{\bm{\theta}}}\right).$
Therefore ${\bm{\psi}}^{(i)}$ is in the form of its first-order derivative:
${\bm{\psi}}^{(i)}=X_{i}\left(y_{i}-\frac{e^{X_{i}^{\top}{\bm{\theta}}}}{1+e^{X_{i}^{\top}{\bm{\theta}}}}\right),$
and
$\dot{{\bm{\psi}}}=\mathbb{E}\left[X_{i}X^{\top}_{i}\mu_{i}(1-\mu_{i})\right],$
where $\mu_{i}=\frac{e^{X_{i}^{\top}{\bm{\theta}}}}{1+e^{X_{i}^{\top}{\bm{\theta}}}}$. We can similarly obtain ${\bm{\psi}}^{(i)}_{k}$ and $\dot{{\bm{\psi}}}_{k}$ by substituting the $y_{i}$ and $X_{i}$ with their imputed counterpart under the $k$-th missingness pattern.
S1.5 Comment on the PS-PPI estimator as an AIPW estimator
In this subsection we make comment on why $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is an AIPW estimator. Specifically, when the propensity score model is known, the $i$-th influence function of the AIPW estimators, as defined in Tsiatis, (2006), is:
$\varphi_{\text{Tsiatis},i}$ $=-\mathbb{E}\left\{\frac{\partial{\bm{\psi}}}{\partial{\bm{\theta}}}\Bigr{|}_{{\bm{\theta}}={\bm{\theta}}^{\ast}}\right\}^{-1}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}}^{\ast})+\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}\left(\sum^{K}_{k=1}\pi_{k}(G_{k}(\bm{Y}_{i}))L_{2k}(G_{k}(\bm{Y}_{i}))\right)\right$.
$\;\;\;\left.-\sum^{K}_{k=1}I({\mathcal{C}}=k)L_{2k}(G_{k}(\bm{Y}_{i}))\right]$
$=-\dot{{\bm{\psi}}}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}^{(i)}-\sum^{K}_{k=1}\left(\frac{I({\mathcal{C}}_{i}=\infty)\pi_{k}(G_{k}(\bm{Y}_{i}))}{\pi_{\infty}(\bm{Y}_{i})}-I({\mathcal{C}}=k)\right)\dot{{\bm{\psi}}}^{-1}L_{2k}(G_{k}(\bm{Y}_{i})).$
where $L_{2k}(G_{k}(\bm{Y}_{i}))$ are a series of arbitrarily chosen functions. By carefully designing $L_{2k}(G_{k}(\bm{Y}_{i}))$ as:
$L_{2k}(G_{k}(\bm{Y}_{i}))=-\frac{1}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\dot{{\bm{\psi}}}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}{\bm{\psi}}^{(i)}_{k}$,
we have $\varphi_{\text{Tsiatis},i}={\bm{\Gamma}}_{i}$, which suggests that the $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ estimator is an AIPW estimator when the propensity score model is known. Similar conclusion could be derived when the propensity score model is unknown but correctly specified, as one only need to further expand $\widehat{{\bm{\omega}}}$ in ${\bm{\Gamma}}_{i}$ (and $\varphi_{\text{Tsiatis},i}$) to obtain $\varphi_{i}$.
S1.6 Proof of Corollary 3.2
Proof.
If $\mathbf{W}_{k}$ are the same across $k$, then
$\Sigma_{{\bm{\theta}}_{\text{PS-PPI}}}=\Sigma_{{\bm{\theta}}}-\sum^{K}_{k=1}\mathbf{W}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}-\sum^{K}_{k=1}(\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}})^{\top}\mathbf{W}^{\top}+\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\mathbf{W}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\mathbf{W}^{\top}+\sum^{K}_{k=1}\mathbf{W}\Sigma_{{\bm{\gamma}}_{2,k}}\mathbf{W}^{\top}.$
Equating its first-order derivative to $\bm{0}$:
$-\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}-\sum^{K}_{k=1}(\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}})^{\top}+2\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\mathbf{W}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}+2\sum^{K}_{k=1}\mathbf{W}\Sigma_{{\bm{\gamma}}_{2,k}}=\bm{0}.$
Solving the equation we know that
$\mathbf{W}^{\ast}=\left(\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}\right)\times\left[\left(\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\right)+\left(\sum^{K}_{k=1}\Sigma_{{\bm{\gamma}}_{2,k}}\right)\right]^{-1}$
leads to the optimal efficiency of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$. Plugging in $\mathbf{W}^{\ast}$ into Equation (7), we have
$\Sigma_{\text{PS-PPI}}=\Sigma_{{\bm{\theta}}}-\left(\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}\right)\times\left[\left(\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\right)+\left(\sum^{K}_{k=1}\Sigma_{{\bm{\gamma}}_{2,k}}\right)\right]^{-1}\times\left(\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}\right)^{\top}.$
Furthermore, we know that
$\left(\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\right)+\left(\sum^{K}_{k=1}\Sigma_{{\bm{\gamma}}_{2,k}}\right)=\textrm{Var}\left(\sum^{K}_{k=1}\widehat{{\bm{\gamma}}}_{1,k}\right)+\sum^{K}_{k=1}\textrm{Var}(\widehat{{\bm{\gamma}}}_{2,k}).$
This suggests that $\left(\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\right)+\left(\sum^{K}_{k=1}\Sigma_{{\bm{\gamma}}_{2,k}}\right)$ is a positive definite matrix, therefore its inverse is a positive definite matrix. Therefore,
$\left(\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}\right)\times\left[\left(\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\right)+\left(\sum^{K}_{k=1}\Sigma_{{\bm{\gamma}}_{2,k}}\right)\right]^{-1}\times\left(\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}\right)^{\top}$
is positive definite matrix. As a result, we have $\Sigma_{\text{PS-PPI}}\preceq\Sigma_{{\bm{\theta}}}$.
Now we prove the if-and-only-if condition. If $\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}=\bm{0}$, then obviously $\Sigma_{\text{PS-PPI}}=\Sigma_{{\bm{\theta}}}$. If $\Sigma_{\text{PS-PPI}}=\Sigma_{{\bm{\theta}}}$, since $\left[\left(\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\right)+\left(\sum^{K}_{k=1}\Sigma_{{\bm{\gamma}}_{2,k}}\right)\right]^{-1}$ is a positive definite matrix, it cannot be $\bm{0}$. Hence, $\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}=\bm{0}$. ∎
S1.7 Proof of Proposition 2
Proof.
From Gronsbell et al., (2024), we know that
$\textbf{IF}_{i,\text{Chen}}=-\dot{{\bm{\psi}}}^{-1}\left\{{\bm{\psi}}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}+\text{Cov}({\bm{\psi}},{\bm{\psi}}_{1})\text{Var}({\bm{\psi}}_{1})^{-1}{\bm{\psi}}_{1}\frac{\pi_{\infty}-I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}\right\}.$
Recall that when $\pi(\cdot)$ is known, the $i$-th influence function of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is:
$\textbf{IF}_{i,\text{PS-PPI}}=-\dot{{\bm{\psi}}}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}+\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}_{\infty}(\bm{Y}_{i})}-\frac{I({\mathcal{C}}_{i}=k)}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\right){\bm{\psi}}_{k}.$
Under the missing outcome setting and the MCAR assumption,
$\textbf{IF}_{i,\text{PS-PPI}}=-\dot{{\bm{\psi}}}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}+\mathbf{W}_{1}\dot{{\bm{\psi}}}^{-1}_{1}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}-\frac{I({\mathcal{C}}_{i}=1)}{\pi_{1}}\right){\bm{\psi}}_{1}.$
From Corollary 3.2, we know that the optimal weight matrix:
$\mathbf{W}^{\ast}=\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,1}}\left(\Sigma_{{\bm{\gamma}}_{1,1}}+\Sigma_{{\bm{\gamma}}_{2,1}}\right)^{-1},$
where
$\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,1}}=\dot{{\bm{\psi}}}^{-1}\mathbb{E}\left[\frac{{\bm{\psi}}{\bm{\psi}}^{\top}_{1}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}^{-1\top}_{1}=\dot{{\bm{\psi}}}^{-1}\text{Cov}({\bm{\psi}},{\bm{\psi}}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\pi^{-1}_{\infty},$
$\Sigma_{{\bm{\gamma}}_{1,1}}=\dot{{\bm{\psi}}}^{-1}_{1}\mathbb{E}\left[\frac{{\bm{\psi}}_{1}{\bm{\psi}}^{\top}_{1}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}^{-1\top}_{1}=\dot{{\bm{\psi}}}^{-1}_{1}\text{Var}({\bm{\psi}}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\pi^{-1}_{\infty},$
$\Sigma_{{\bm{\gamma}}_{2,1}}=\dot{{\bm{\psi}}}^{-1}_{1}\mathbb{E}\left[\frac{{\bm{\psi}}_{1}{\bm{\psi}}^{\top}_{1}}{\pi_{1}(G_{1}(\bm{Y}_{i}))}\right]\dot{{\bm{\psi}}}^{-1\top}_{1}=\dot{{\bm{\psi}}}^{-1}_{1}\text{Var}({\bm{\psi}}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}(1-\pi_{\infty})^{-1}.$
Therefore,
$\mathbf{W}^{\ast}$ $=\left(\dot{{\bm{\psi}}}^{-1}\text{Cov}({\bm{\psi}},{\bm{\psi}}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\pi^{-1}_{\infty}\right)\times\left(\dot{{\bm{\psi}}}^{-1}_{1}\text{Var}({\bm{\psi}}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\right)^{-1}\pi_{\infty}\pi_{1}$
$=\left(\dot{{\bm{\psi}}}^{-1}\text{Cov}({\bm{\psi}},{\bm{\psi}}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\pi^{-1}_{\infty}\right)\times\left(\dot{{\bm{\psi}}}^{\top}_{1}\text{Var}({\bm{\psi}}_{1})^{-1}\dot{{\bm{\psi}}}_{1}\right)\pi_{\infty}\pi_{1}$
$=\dot{{\bm{\psi}}}^{-1}\text{Cov}({\bm{\psi}},{\bm{\psi}}_{1})\text{Var}({\bm{\psi}}_{1})^{-1}\dot{{\bm{\psi}}}_{1}\pi_{1}.$
Plugging $\mathbf{W}^{\ast}$ back to $\textbf{IF}_{i,\text{PS-PPI}}$:
$\textbf{IF}_{i,\text{PS-PPI}}$ $=-\dot{{\bm{\psi}}}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}{\bm{\psi}}+\dot{{\bm{\psi}}}^{-1}\text{Cov}({\bm{\psi}},{\bm{\psi}}_{1})\text{Var}({\bm{\psi}}_{1})^{-1}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}-\frac{I({\mathcal{C}}_{i}=1)}{\pi_{1}}\right){\bm{\psi}}_{1}\pi_{1}$
$=-\dot{{\bm{\psi}}}^{-1}\left\{{\bm{\psi}}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}+\text{Cov}({\bm{\psi}},{\bm{\psi}}_{1})\text{Var}({\bm{\psi}}_{1})^{-1}{\bm{\psi}}_{1}\frac{\pi_{\infty}-I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}\right\}=\textbf{IF}_{i,\text{Chen}}.$
Therefore, the two estimators are asymptotically equivalent. ∎
S1.8 Proof of Proposition 3
Proof.
From Kluger et al., (2025), we know that the asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{PTD}}$ is
$\Sigma_{\text{PTD}}=\Sigma_{{\bm{\theta}}}-\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1}}\left(\Sigma_{{\bm{\gamma}}_{1}}+\Sigma_{{\bm{\gamma}}_{2}}\right)^{-1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1}}^{\top},$
where $\Sigma_{{\bm{\theta}}}$ is the asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{WCC}}$; $\Sigma_{{\bm{\gamma}}_{1}}$ is the asymptotic variance of $\widehat{{\bm{\gamma}}_{1}}$ (the second subscript is removed as there is only two missingness pattern, and one of them is the fully-observed pattern); $\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1}}$ is the asymptotic covariance between $\widehat{{\bm{\theta}}}_{\text{WCC}}$ and $\widehat{{\bm{\gamma}}_{1}}$; $\Sigma_{{\bm{\gamma}}_{2}}$ is the asymptotic variance of $\widehat{{\bm{\gamma}}_{2}}$.
On the other hand, under the MAR assumption, with only two missingness patterns, Equation (8) becomes:
$\Sigma_{\text{PS-PPI}}=\Sigma_{{\bm{\theta}}}-\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1}}\left(\Sigma_{{\bm{\gamma}}_{1}}+\Sigma_{{\bm{\gamma}}_{2}}\right)^{-1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1}}^{\top}.$
This suggests that both $\widehat{{\bm{\theta}}}_{\text{PTD}}$ and $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ have the same asymptotic variance. As they also have the same asymptotic mean, they are asymptotically equivalent. ∎
Appendix S2 Additional Experimental Results
S2.1 Synthetic data simulation
In this subsection we present the additional experimental results of the synthetic data simulation under different simulation settings. Specifically, instead of fixing the prediction bias level, we fix the prediction noise level at $\sigma_{\text{pred}}=0$ and vary the bias level and present the results in Figure S1. Based on the simulation settings adopted in Figure 2 and S1, we correspondingly substitute the estimated propensity score with the real propensity score and present the simulation results in Figure S2 and S3. Similarly, we substitute the estimated propensity score with the one obtained from the misspecified model and present the simulation resutls in Figure S4 and S5.
Figure S1: The same as Figure 2, except that the noise level is fixed at $\sigma_{\text{pred}}=0$, and X-axis in all panels represents the prediction bias level used in the simulations.
Figure S2: The same as Figure 2, except that the propensity scores are assumed to be known.
Figure S3: The same as Figure 2, except that the propensity scores are assumed to be known, and the noise level is fixed at $\sigma_{\text{pred}}=0$, and X-axis in all panels represents the prediction bias level used in the simulations.
Figure S4: The same as Figure 2, except that the propensity scores are assumed to be estimated from a misspecified model.
Figure S5: The same as Figure 2, except that the propensity scores are assumed to be estimated from a misspecified model, and the noise level is fixed at $\sigma_{\text{pred}}=0$, and X-axis in all panels represents the prediction bias level used in the simulations.