### v2

# A Unified Framework for Inference with General Missingness Patterns and Machine Learning Imputation

Abstract
Pre-trained machine learning (ML) predictions have been increasingly used to complement incomplete data to enable downstream scientific inquiries, but their naive integration risks biased inferences. Recently, multiple methods have been developed to provide valid inference with ML imputations regardless of prediction quality and to enhance efficiency relative to complete-case analyses. However, existing approaches are often limited to missing outcomes under a missing-completely-at-random (MCAR) assumption, failing to handle general missingness patterns (missing in both the outcome and exposures) under the more realistic missing-at-random (MAR) assumption. This paper develops a novel method that delivers a valid statistical inference framework for general Z-estimation problems using ML imputations under the MAR assumption and for general missingness patterns. The core technical idea is to stratify observations by distinct missingness patterns and construct an estimator by appropriately weighting and aggregating pattern-specific information through a masking-and-imputation procedure on the complete cases. We provide theoretical guarantees of asymptotic normality of the proposed estimator and efficiency dominance over weighted complete-case analyses. Practically, the method affords simple implementations by leveraging existing weighted complete-case analysis software. Extensive simulations are carried out to validate theoretical results. A real data example is provided to further illustrate the practical utility of the proposed method. The paper concludes with a brief discussion on practical implications, limitations, and potential future directions.
1 Introduction
With the rapid development of machine learning (ML) and artificial intelligence (AI) techniques, ML models, particularly those trained on large-scale datasets, have become increasingly capable of making highly accurate predictions. In statistical analysis, these predictions have the potential to be treated as imputations for missing data (in both dependent $Y$ and independent variables $\bm{X}$) due to nonresponse, expensive data collection costs, or other factors, and then the imputed datasets can be seamlessly applied in downstream statistical analyses. However, such an appealing impute-then-analyze procedure is a double-edged sword. On the one hand, the ease of obtaining ML predictions for imputation catalyzes wide adoption across various disciplines, including finance (bryzgalova2025missing), social science (egami2023using), clinical psychology (little2024missing), biomedicine (barrabes2024advances), and public health (perkins2018principled) studies, where missing data appear commonly during the data collection process. On the other hand, imputing missing values with ML predictions of arbitrary quality may produce datasets that have different conditional and marginal distributions from the underlying true distribution. Proceeding with analysis without careful consideration may result in biased estimation and invalid inference.
To address this issue, recent research has explored prediction-based methods to incorporate ML predictions for single imputation while guaranteeing valid statistical inference on datasets with partially missing outcomes (see, for example, wang2020methods; angelopoulos2023prediction; egami2023using; salerno2025moment; salerno2025ipd). Specifically, prediction-powered inference (PPI, angelopoulos2023prediction) is a framework where inference is conducted using ML-imputed outcomes, with model parameter estimates calibrated, or “rectified”, using a complete-case subset. Since this approach does not impose assumptions on the imputation model itself, it allows for a broad selection of ML-based imputation methods, the quality of which impacts the efficiency but not the validity of inference (e.g., consistency, achieving nominal coverage rate). Several extensions of this framework have adapted the idea of PPI to important areas, including causal inference (demirel2024prediction), Bayesian inference (hofer2024bayesian; rovckova2025ai), and genome-wide association studies (GWAS, miao2024valid). Furthermore, recognizing that the PPI approach does not guarantee a smaller variance relative to complete-case analysis (CCA), a few authors have proposed weighting the rectifier with tuning parameters to achieve better efficiency than CCA (angelopoulos2023PPI++; miao2023assumption; gan2024prediction; gronsbell2024another; ji2025predictions; xu2025unified).
Despite the tremendous progress and development in prediction-based inference, two key challenges remain unaddressed. First, the existing literature considers only the assumption of missing completely at random (MCAR), leaving an open question of the applicability of prediction-based inference under the more common and complicated assumption, that is, the assumption missing at random (MAR) (rubin1976inference; little2019statistical). Except for several specific cases such as linear regression analysis and odds ratio estimation under certain missing probability assumptions (Example 3.3 and 3.4, little2019statistical), failure to account for the MAR assumption can result in biased estimates. Second, the current literature focuses exclusively on partially missing outcomes or specific patterns in covariate missingness (e.g., missing multiple covariates kluger2025prediction), but does not account for general patterns of non-monotone missingness, where missing values can occur in any cell within the tabular (“rectangular”) dataset. For example, in regression analysis, missing values are common in both outcomes and covariates, and we may observe different patterns of missingness in different subjects, e.g., missing $Y$ and $X_{1}$ for one subject and missing $Y$ and $X_{2}$ for another.
In this paper, we address these challenges by proposing a unified framework called Pattern-Stratified PPI (PS-PPI) [^1] By Corollary 3.2, the proposed procedure enjoys efficiency dominance over WCCA (under MAR, general missing patterns). This resembles the property of PPI++ being no less efficient than CCA (under MCAR, outcome missing only). We chose to term our method “PS-PPI” rather than “PS-PPI++” to avoid redundancy. to account for both general missingness patterns and missing-at-random mechanism under the Z-estimation framework. We make a few theoretical and methodological contributions with useful practical implications. First, we develop an inferential procedure that integrates the concept of missingness patterns that are used in semiparametric inference to deal with non-monotone missing data (tsiatis2006semiparametric). We will show that the proposed estimator belongs to the class of Augmented Inverse Probability Weighted (AIPW) estimators, building on a prior work under MCAR with only outcomes subject to missingness (gronsbell2024another). Second, we provide a theoretical guarantee that the resulting PS-PPI estimator is at least as efficient as the weighted complete-case analysis (WCCA) even when the imputation models are of low quality. Third, our proposed PS-PPI procedure is straightforward to implement. It works by breaking the analysis into several weighted complete-case (WCC) Z-estimation problems, each of which can be solved with existing statistical packages designed for WCC analysis.
The remainder of the paper is organized as follows. In Section 2, we introduce necessary notation, the concepts of missing data mechanisms and missingness patterns, and the inferential target under the Z-estimation framework; we review the core idea used in the existing prediction-based inference literature that deals with missing outcome or missing multiple covariates, along with their limitations. In Section 3, we introduce the details of the proposed PS-PPI procedure and present theoretical results on its asymptotic properties; we provide examples to show that our proposed framework subsumes previous works as special cases. In Section 4, we perform data simulation to verify the validity and superior performance of our proposed PS-PPI relative to alternatives that have limited ability to extract and integrate information from incomplete datasets with multiple different missingness patterns. We provide a real data example in Section 5 to illustrate the practical utility of the PS-PPI method. Finally, we discuss the limitations of this work and potential future directions in Section 6.
2 Setup
2.1 Notation
Let $\mathcal{D}=\{\bm{Y}_{1},\ldots,\bm{Y}_{N}\}$ denote the full data for $N$ observations, subject to potential missingness, where $\bm{Y}_{i}=(Y_{i1},\ldots,Y_{ip})^{\top}\in{\mathcal{Y}}={\mathcal{Y}}_{1}\times\cdots\times{\mathcal{Y}}_{p}$ represents the $p$-dimensional data for subject $i$, ${\mathcal{Y}}_{j}\subseteq\mathbb{R}$ is the marginal sample space of the $j$-th component in $\bm{Y}_{i}$. Each $\bm{Y}_{i}$ is assumed to be independently and identically sampled from a population distribution $\mathbb{P}_{\bm{Y}}$. In practice, three types of variables may be included in the vector $\bm{Y}_{i}$: (a) a variable treated as the scalar outcome in a specific scientific inquiry; (b) q1q-1 variables treated as covariates (e.g., main exposure and potential confounders); (c) pqp-q auxiliary variables not included in the scientific analysis but potentially useful to predict variables in type (a) and (b). In a real data analysis, one or more variables of type (b) may not be subject to missingness, which can be used along with type (c) variables as inputs to ML prediction models for imputations. We partition $\bm{Y}_{i}$ into three types of variables for ease of presentation; this partition does not impact the validity of our theoretical results. For readers who require this distinction in their scientific inquiries (e.g., via a regression analysis), one may assume that $Y_{i1}$ and $Y_{i2},\ldots,Y_{iq}$} represent the outcome and the covariates, respectively.
We assume that the analyst may not observe the full data $\mathcal{D}$}. Let $\bm{R}_{i}=(R_{i1},\ldots,R_{ip})^{\top}\\ \in\{0,1\}^{p}$ be a binary vector of $p$ dimension, where $R_{ij}=1$ indicates that $Y_{ij}$ is observed and $R_{ij}=0$ otherwise. Consequently, we denote the observed components $\bm{Y}_{i,\text{obs}}=\{Y_{ij}|R_{ij}=1,j=1,\ldots,p\}\in\prod_{\{j:R_{ij}=1\}}{\mathcal{Y}}_{j}$, and similarly the missing components $\bm{Y}_{i,\text{mis}}=\{Y_{ij}|R_{ij}=0,j=1,\ldots,p\}\in\prod_{\{j:R_{ij}=0\}}{\mathcal{Y}}_{j}$, for subject $i=1,\ldots,N$. Let ${\mathcal{D}}_{\text{obs}}=\{\bm{Y}_{1,\text{obs}},...,\bm{Y}_{N,\text{obs}}\}$ denote the observed data.
2.1.1 Missingness patterns and mechanisms
In the existing literature on prediction-based inference, typically only the outcome variable is subject to potential missingness (e.g., angelopoulos2023prediction; angelopoulos2023PPI++; miao2023assumption; gan2023prediction; gronsbell2024another), with the exception of kluger2025prediction, which focused on the missing-multiple-covariate pattern. However, in practice, missingness can occur anywhere within the dataset, e.g., for both outcomes and covariates under the regression setting.
To address a more general missingness structure, we use the concept of missingness patterns and formalize the general missingness pattern problem by treating each missingness pattern as a special case of coarsening (heitjan1991ignorability), following tsiatis2006semiparametric. A missingness pattern is defined as an element in the sample space $\{0,1\}^{p}$ of the missing indicator vector $\bm{R}_{i}$. Excluding the case where all variables are missing, there are at most $2^{p}-1$ different missingness patterns appearing in the data. In the context of prediction-based inference, it is generally assumed that there exist fully observed, low-cost variables that serve as inputs for predicting the missing values (i.e., type (c) variables). Under this assumption, the number of distinct missingness patterns is reduced to at most $2^{q}$. Each of the missingness patterns can be indexed with $k\in\{1,...,K\}$; we define $k=\infty$ to indicate that the subject is fully observed. The $k$-th missingness pattern is denoted as $\bm{R}^{(k)}=(R^{(k)}_{1},\ldots,R^{(k)}_{p})\in\{0,1\}^{p}$. We call the $i$-th subject $\bm{Y}_{i}$ to be in the $k$-th stratum of missingness pattern if $\bm{R}_{i}=\bm{R}^{(k)}$. Following tsiatis2006semiparametric, for the $k$-th missingness pattern, we introduce a deterministic and many-to-one coarsening function $G_{k}:{\mathcal{Y}}\mapsto\prod_{\{j:R^{(k)}_{j}=1\}}{\mathcal{Y}}_{j}$, such that $G_{k}(\bm{Y}_{i})=\bm{Y}_{i,\text{obs}}$ if $\bm{Y}_{i}$ is under the $k$-th missingness pattern. For fully observed data, $G_{\infty}(\bm{Y}_{i})$ is an identity mapping. Note that the coarsening notation is borrowed only to emphasize the constructive nature of our proposed method, which combines pattern-stratified estimates (see Equation (2)); the prediction-based inference for general coarsened data where $G_{k}$ is broader than entry subset selection is beyond the scope of this work. With the notation of the coarsening functions and missingness patterns, we can represent the observed dataset as ${\mathcal{D}}_{\text{obs}}=\{(G_{{\mathcal{C}}_{i}}(\bm{Y}_{i}),{\mathcal{C}}_{i})\}^{N}_{i=1}$, where ${\mathcal{C}}_{i}\in\{1,...,K,\infty\}$ is the coarsening index for subject $i$.
Missing data mechanisms play an important role in dictating how researchers should deal with the missingness in data analyses (rubin1976inference). Here we provide a brief overview for completeness and refer readers to little2019statistical for a comprehensive, book-length treatment of this topic. Based on how the conditional distribution of the coarsening index ${\mathcal{C}}_{i}$ given $\bm{Y}_{i}$ can be simplified, a few general categories with distinct analytic implications can be defined (tsiatis2006semiparametric). In the simplest case with the most restrictive assumption of missing completely at random (MCAR), the probability distribution of the coarsening index ${\mathcal{C}}_{i}$ does not depend on any observed or unobserved variables, namely, $\pi({\mathcal{C}}_{i}=k|\bm{Y}_{i};{\bm{\omega}})=\pi({\mathcal{C}}_{i};{\bm{\omega}})$ where $\bm{\omega}$} is a vector of parameters of the conditional distribution. A more common and realistic case compared to MCAR is that the probability of a missingness pattern only depends on observed variables: $\pi({\mathcal{C}}_{i}|\bm{Y}_{i};{\bm{\omega}})=\pi({\mathcal{C}}_{i}|\bm{Y}_{i,\text{obs}};{\bm{\omega}})$, referred to as missing at random (MAR). If the missingness depends on the missing values themselves: $\pi({\mathcal{C}}_{i}|\bm{Y}_{i};{\bm{\omega}})=\pi({\mathcal{C}}_{i}|\bm{Y}_{i,\text{obs}},\bm{Y}_{i,\text{mis}};{\bm{\omega}}),$ it is referred to as missing not at random (MNAR). In practice, the MAR assumption is more commonly considered because it does not suffer from the identifiability problems that exist in the MNAR assumption, yet it is weaker than the MCAR assumption (tsiatis2006semiparametric). In this paper, we will introduce and provide theoretical analyses of our methods under the MAR assumption.
2.1.2 Prediction of missing or masked values
For each missingness pattern $k=1,\ldots,K$, suppose we have access to a prediction function $h_{k}$ trained on data external to ${\mathcal{D}}_{\text{obs}}$ where $h_{k}:\prod_{\{j:\bm{R}^{(k)}_{j}=1\}}{\mathcal{Y}}_{j}\mapsto\prod_{\{j:\bm{R}^{(k)}_{j}=0\}}{\mathcal{Y}}_{j}$. A key ingredient in our proposed method is the following “mask-and-impute” operation: in addition to filling the missing values for the subjects in the stratum of the $k$-th missingness pattern using $h_{k}$, we can also mask the entries for any fully-observed subject according to the missingness pattern $k$ (i.e., $\{Y_{ij}:j\in\{j^{\prime}:R_{j^{\prime}}^{(k)}=0\}\}$ for subject $i$ with ${\mathcal{C}}_{i}=\infty$), and then use the same $h_{k}$ to obtain their predictions to impute these “pseudo” missing entries.
In fact, to study prediction-based inference, we will show in our theoretical analyses that it is central to generalize this notion of data masking-and-imputation by missingness pattern $k$ from the fully observed subjects (${\mathcal{C}}_{i}=\infty$) to the entire population. For each missingness pattern $k=1,\ldots,K$, let the random vector $\bm{Y}^{(k)}_{\text{imp}}=\left\{G_{k}(\bm{Y}),h_{k}\left(G_{k}(\bm{Y})\right)\right\}$ be comprised of observed components $G_{k}(\bm{Y})$ and masked-and-imputed components $h_{k}\left(G_{k}(\bm{Y})\right)$, where $\bm{Y}$ follows the population full-data distribution $\mathbb{P}_{\bm{Y}}$. We assume that $\bm{Y}^{(k)}_{\text{imp}}$ is distributed according to a population-level distribution $\mathbb{P}^{(k)}_{\bm{Y}_{\text{imp}}}$. We note two key points. First, $h_{k}$ can be deterministic or stochastic as long as there exists a fixed population distribution $\mathbb{P}^{(k)}_{\bm{Y}_{\text{imp}}}$ which we will assume in theoretical analyses (Section 3). This assumption is less tenable if the prediction mappings $\{h_{k}\}$ shift over time, e.g., due to model updates that are continually pushed into online repositories to progressively impute different subsets of data. Second, the current setup does not require distinct prediction models to be trained for each missingness pattern. In practice, the predictions can be accomplished by using at most $q$ univariate base prediction models for type (a) and (b) variables, whenever a missing value is encountered. In this sense, $h_{k}$ can be viewed as invoking a subset of $\sum_{j}(1-R^{(k)}_{j})$ of the $q$ base prediction models for missingness pattern $k$. For example, the model for predicting diabetes status using BMI only can be used for any missingness pattern $k$ with observed BMI and missing diabetes status. As a result, although the number of missingness patterns may grow exponentially with the number of variables $p$, the number of base prediction models required increases only linearly.
2.2 Inferential target
In this paper, we focus on obtaining valid statistical inference on a $d$-dimensional parameter $\bm{\theta}$} in a Z-estimation framework based on a dataset with general missingness patterns under the MAR assumption. Z-estimation is a popular framework for defining estimands and studying properties of estimators in statistical literature. Several most commonly used statistical methods, such as generalized linear model and generalized estimating equations, can be understood via the Z-estimation framework. Formally, define the true population parameter $\bm{\theta}$$^{\ast}$ as the unique solution to
$\mathbb{E}_{\mathbb{P}_{\bm{Y}}}[{\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}})]=\bm{0},$
where $\bm{\psi}$} are known vector-valued maps to a typically $d$-dimensional space, as defined in van2000asymptotic and ${\bm{\theta}}\in\mathbb{R}^{d}$ is the parameter attached to the population distribution $\mathbb{P}_{\bm{Y}}$. For independent and identically distributed, fully observed data $\mathcal{D}$} of size $N$, a Z-estimator $\widehat{{\bm{\theta}}}_{\textrm{Full}}$ is obtained by solving the following empirical estimating equations:
$\frac{1}{N}\sum^{N}_{i=1}{\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}})=\bm{0}.$ (1)
Under certain regularity conditions introduced in van2000asymptotic and detailed in Section S1.2 in the Supplementary Materials, $\widehat{{\bm{\theta}}}_{\text{Full}}$ is asymptotically normal with mean $\bm{\theta}$$^{\ast}$ and variance $\dot{{\bm{\psi}}}^{-1}_{\textrm{Full}}\mathbb{E}[{\bm{\psi}}^{*}{\bm{\psi}}^{*\top}]\left(\dot{{\bm{\psi}}}^{-1}_{\textrm{Full}}\right)^{\top}$, where $\dot{{\bm{\psi}}}_{\text{Full}}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}}{\partial{\bm{\theta}}}\Bigr|_{{\bm{\theta}}={\bm{\theta}}^{\ast}}\right]$ is an invertible $d\times d$ matrix and ${\bm{\psi}}^{*}={\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}}^{\ast})$ with $\mathbb{E}[{\bm{\psi}}^{*}{\bm{\psi}}^{*\top}]<\infty$.
2.3 Rationale of prediction-based inference
In the presence of missing values, we will propose a general approach and provide the theoretical guarantee of valid inference in Section 3. The proposed approach couples the insights from the original outcome-based prediction-powered inference (PPI) framework (angelopoulos2023prediction; angelopoulos2023PPI++) and the classical augmented inverse probability weighted (AIPW) estimators, of which outcome-based PPI is a special case (gronsbell2024another). It is useful to review the rationale behind the original outcome-based PPI approach, which is designed to incorporate additional information from observations with missing outcomes and the predicted outcomes. Technically, they can be written as a difference between a base estimate (CCA estimate) $\widehat{{\bm{\theta}}}_{\text{CC}}$ and an estimate of zero $\widehat{\mathbf{\Delta}}$: [^2] The notation of $\widehat{\mathbf{\Delta}}$ in this paper differs from the one used in angelopoulos2023prediction which denotes a rectifier to correct the bias introduced by ML predictions. The current form, in our view, is more helpful for motivating the construction of more efficient estimators and for theoretical analyses.
$\widehat{{\bm{\theta}}}_{\text{PP}}=\widehat{{\bm{\theta}}}_{\text{CC}}-\widehat{\mathbf{\Delta}}$,
[…]
Example 1 (PPI for mean outcome estimation).
We illustrate with a simple example based on population mean estimation of a scalar outcome $Y$ in the presence of outcome missingness introduced in angelopoulos2023prediction. Suppose the data are independently and identically distributed according to a population distribution comprised of two partitions of size $m$ and $M$ with marginal mean outcome $\mu=\mathbb{E}[Y]$. For the first partition, we observe the outcome $Y^{\tt lab}_{i}$ and their corresponding ML-predicted alternatives $\widehat{Y}^{\tt lab}_{i}$: ${\mathcal{D}}_{\tt lab}=\{(Y^{\tt lab}_{i},\widehat{Y}^{\tt lab}_{i})\}^{m}_{i=1}$, which is referred to as the “labeled subset”. For the second, often larger, partition of data, we do not get to observe the outcome but can obtain their ML-predicted surrogates, i.e., ${\mathcal{D}}_{\tt unlab}=\{\widehat{Y}^{\tt unlab}_{i}\}^{M}_{i=1}$, referred to as the “unlabeled subset”. We assume the outcome is MCAR, i.e., the full data distribution of $Y$ between the two partitions is identical. For both partitions, the predictions are obtained from other observed covariates and prediction models that have been externally trained to predict $Y$ using these covariates. As such, the pair (Y,Y^)($\widehat{Y}$) has the same distribution between the labeled and unlabeled partitions, although $Y$ is missing for the unlabeled subset.
In PPI, the CCA estimator for $\mu$ is the empirical mean based on the complete subset of data: $\widehat{\theta}_{\text{CC}}=\bar{Y}^{\tt lab}$, and the corresponding estimation of zero is computed as $\widehat{\Delta}=\bar{\widehat{Y}}^{\tt lab}-\bar{\widehat{Y}}^{\tt unlab}$ because the predicted outcomes in the labeled and unlabeled partitions share the same distribution under MCAR. By rewriting $\widehat{\theta}_{\text{PP}}=\widehat{\theta}_{\text{CC}}-\widehat{\Delta}=\bar{Y}^{\tt lab}-(\bar{\widehat{Y}}^{\tt lab}-\bar{\widehat{Y}}^{\tt unlab})=\bar{\widehat{Y}}^{\tt unlab}-\underbrace{(\bar{\widehat{Y}}^{\tt lab}-\bar{Y}^{\tt lab})}_{\tt``rectifier"}$, we have
$\mathrm{Var}(\widehat{\theta}_{\text{PP}})=\frac{1}{M}\mathrm{Var}(\widehat{Y}^{\tt unlab}_{i})+\frac{1}{m}\mathrm{Var}(Y^{\tt lab}_{i}-\widehat{Y}^{\tt lab}_{i}),$
where the denominator of the first term is determined by the sample size of the unlabeled subset $M$, while the numerator of the second term is determined by the accuracy of the outcome prediction. $\widehat{\theta}_{\text{PP}}$ will have a small variance when $M\gg m$ to let the second term dominate, and the machine learning predictions are precise enough such that the variance in the second term is small. To see this, when $M\gg m$, we will have $\mathrm{Var}(\widehat{\theta}_{\text{PP}})\approx\frac{1}{m}\mathrm{Var}(Y^{\tt lab}_{i}-\widehat{Y}^{\tt lab}_{i})<\mathrm{Var}(\widehat{\theta}_{\text{CC}})=\frac{1}{m}\mathrm{Var}(Y_{i}^{\tt lab})$ when the predictions are of good quality, or more precisely when the correlation satisfies $\text{Corr}(Y,\widehat{Y})>\frac{1}{2}\sqrt{\mathrm{Var}(Y)/\mathrm{Var}(\widehat{Y})}$.
However, it is critical to understand when $\widehat{{\bm{\theta}}}_{\text{PP}}$ is more efficient than the complete-case analysis in settings other than the mean estimation problem. gronsbell2024another analyzed the efficiency of the $\widehat{{\bm{\theta}}}_{\text{PP}}$ by deriving its influence function under the linear regression setting, and connecting it to the class of AIPW estimators. Specifically, gronsbell2024another shows that $\widehat{{\bm{\theta}}}_{\text{PP}}$ is an AIPW estimator and is not the most efficient in its class. Although the optimal estimator is difficult to obtain, a weighted version of the PPI estimator was proposed to guarantee the efficiency improvement compared to the complete-case analysis (gronsbell2024another). The estimator is of the form $\widehat{{\bm{\theta}}}_{\text{Chen}}=\widehat{{\bm{\theta}}}_{\text{CC}}-\widehat{\mathbf{W}}_{\text{Chen}}\widehat{\mathbf{\Delta}}$, where $\widehat{\mathbf{W}}_{\text{Chen}}$ is a tuning parameter estimated from the data, originally proposed in chen2000unified under a different double-sampling application context but the same setup. Furthermore, recent modifications to the $\widehat{{\bm{\theta}}}_{\text{PP}}$ all consider multiplying $\widehat{\mathbf{\Delta}}$ with a data-driven weight to seek efficiency improvement over $\widehat{{\bm{\theta}}}_{\text{CC}}$ (angelopoulos2023PPI++; miao2023assumption; gan2024prediction).
2.4 Limitations of existing prediction-based inference methods
We now discuss the limitations of the existing methods, which we will address. First, existing prediction-based inference methods do not account for the missing-at-random mechanism. For example, in mean estimation, suppose $Y$ is the annual income response in a questionnaire. An auxiliary variable, the level of education, affects the missing probability of $Y$, such that respondents with a lower level of education tend to skip the response to the income question, leaving their $Y$ missing. In such a case, the CCA estimator $\widehat{\theta}_{\text{CC}}$ itself is biased, because it can only reflect the income levels of the population with higher education levels. Second, existing methods only consider missing outcome or missing-multiple-covariate scenarios, although in practice, general missingness patterns commonly appear. For example, in responding to a questionnaire, respondents may randomly skip multiple questions, leading to non-monotone missingness in the resulting dataset; in a longitudinal study, participants may skip a visit and return in the next visit (sun2018inverse). In such cases, existing prediction-based inference methods can only be carried out by removing the incomplete covariates, or truncating a longitudinal trajectory up to when the first missingness occurred.
3 Pattern-Stratified Prediction-Powered Inference
3.1 Method
Figure 1: Proposed statistical procedure for ensuring valid prediction-based inference in datasets with general missingness patterns under the missing at random (MAR) assumption. (a) We observe a set of samples with missing values. We assign each missingness pattern with an index; (b) For the $k$-th missingness pattern, we collect the samples with ${\mathcal{C}}_{i}=k$, and impute the missing values with ML predictions. Weighted complete-case analysis is then conducted on the imputed dataset to obtain $\widehat{{\bm{\gamma}}}_{2,k}$; (c) For the completely observed samples, we obtain the weighted analysis estimates $\widehat{{\bm{\theta}}}_{\text{WCC}}$. For each $k$-th missingness pattern, we mask the observed values from the fully observed samples to construct each missingness pattern. We then impute the masked values with machine learning predictions and conduct weighted complete-case analysis to obtain $\widehat{{\bm{\gamma}}}_{1,k}$. $\hat{\Delta}_{k}=\hat{\gamma}_{1,k}-\hat{\gamma}_{2,k}$ converges to zero in probability (implication of Proposition 1).
The PS-PPI estimator is defined as the difference between the WCCA estimator $\widehat{{\bm{\theta}}}_{\text{WCC}}$, computed from the fully observed subset, and a weighted summation of $K$ estimators of zero $\widehat{\mathbf{\Delta}}_{k}$ derived from each missingness pattern:
$\widehat{{\bm{\theta}}}_{\text{PS-PPI}}=\widehat{{\bm{\theta}}}_{\text{WCC}}-\sum^{K}_{k=1}\widehat{\mathbf{W}}_{k}\widehat{\mathbf{\Delta}}_{k}.$ (2)
We introduce each term in (2) in turn. First, $\widehat{{\bm{\theta}}}_{\text{WCC}}$ is the WCCA estimator obtained from the following weighted estimating equations:
$\sum_{i=1}^{N}{\bm{\psi}}^{(i)}_{\text{WCC}}=\sum_{i=1}^{N}\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}_{\infty}(\bm{Y}_{i})}{\bm{\psi}}\left(\bm{Y}_{i};{\bm{\theta}}\right)=\bm{0},$ (3)
[…]
$\sum^{N}_{i=1}{\bm{m}}^{(i)}=\sum^{N}_{i=1}{\bm{m}}(G_{{\mathcal{C}}_{i}}(\bm{Y}_{i});{\bm{\omega}})=\bm{0},$ (4)
where $\bm{m}$} is a vector-valued function that has the same dimension as $\bm{\omega}$} and satisfies regularity conditions detailed in Section S1.2 of the Supplementary Materials.
Second, $\widehat{\mathbf{\Delta}}_{k}$ is an estimate of zero obtained from the $k$-th missingness pattern. Specifically, we define $\widehat{\mathbf{\Delta}}_{k}=\widehat{{\bm{\gamma}}}_{1,k}-\widehat{{\bm{\gamma}}}_{2,k}$, where $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{2,k}$ are the solutions to the weighted estimating equations:
$\sum_{i=1}^{N}{\bm{\psi}}^{(i)}_{1,k}=\sum^{N}_{i=1}\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}_{\infty}(\bm{Y}_{i})}{\bm{\psi}}\left(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}}_{1,k}\right)=\bm{0},$ (5)
and
$\sum_{i=1}^{N}{\bm{\psi}}^{(i)}_{2,k}=\sum^{N}_{i=1}\frac{I({\mathcal{C}}_{i}=k)}{\widehat{\pi}_{k}(G_{k}(\bm{Y}_{i}))}{\bm{\psi}}\left(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}}_{2,k}\right)=\bm{0},$ (6)
respectively. That is, we first compute $\widehat{{\bm{\gamma}}}_{1,k}$ by substituting the values in the missing entries with predictions generated from $h_{k}$, and solve the weighted estimating equations based on complete-case data. Second, we compute $\widehat{{\bm{\gamma}}}_{2,k}$ by imputing the actual missing entries with the predictions from $h_{k}$ using data with the $k$-th missingness pattern. We will show that this construction leads to a consistent and asymptotically normal estimator of zero; as shown in Proposition 1, the population parameter of the estimating equations ${\bm{\psi}}_{1,k}$ and ${\bm{\psi}}_{2,k}$ are both ${\bm{\gamma}}^{\ast}_{k}$. We need the following assumption to establish the proposition.
Assumption 1 (Z-estimation target parameter under population-level mask-and-impute distributions).
There exists a ${\bm{\gamma}}^{\ast}_{k}$ such that it is the unique solution of
$\mathbb{E}_{\bm{Y}}[{\bm{\psi}}(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}})]=\bm{0}.$
Proposition 1 (Estimation of zero).
Under Assumption 1, $({\bm{\omega}}^{\ast},{\bm{\gamma}}^{\ast}_{k})$ will lead to both $\mathbb{E}{\bm{\psi}}_{1,k}=\bm{0}$ and $\mathbb{E}{\bm{\psi}}_{2,k}=\bm{0}$, where the expectations are taken over the full data distribution $\mathbb{P}_{\bm{Y}}$.
The proof is detailed in Section S1.1 of the Supplementary Materials. The intuition behind Proposition 1 is that, after accounting for the missing data mechanism, the “pseudo” imputed dataset derived from the complete-case data in Equation (5) is distributed as $\mathbb{P}^{(k)}_{\bm{Y}_{\text{imp}}}$, which is identical to the distribution of the imputed dataset based on the subjects with the $k$-th missingness pattern in Equation (6).
Finally, $\{\widehat{\mathbf{W}}_{k}\}$ is a set of tuning parameters in the form of weight matrices with dimension $d\times d$. $\{\widehat{\mathbf{W}}_{k}\}$ may depend on the observed data. In Corollary 3.2, we provide a specific construction that guarantees efficiency gains over WCCA.
Asymptotic properties of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$.
We show that $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is asymptotically normal.
Theorem 1 (Asymptotic normality of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$).
[…]
The proof is detailed in Section S1.2 in the Supplementary Materials. The asymptotic variance-covariance matrix accounts for the uncertainty of the estimated propensity score models. Following sun2018inverse and robins1994estimation, we provide a convenient robust sandwich estimator that ignores the uncertainty contributed by the estimated propensity score model. Such an estimator results in a slightly more conservative yet valid and implementable sandwich variance for $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ when the propensity scores are estimated; the form is shown in Equation (7). Our numerical experiments in Section 4 validate the efficiency advantage of the PS-PPI estimator based on this strategy.
Corollary 3.1 (Asymptotic variance-covariance matrix of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$).
[…]
$\Sigma_{\text{PS-PPI}}=\Sigma_{{\bm{\theta}}}-\sum^{K}_{k=1}\mathbf{W}_{k}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}-\sum^{K}_{k=1}(\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}})^{\top}\mathbf{W}_{k}^{\top}+\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\mathbf{W}_{k}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\mathbf{W}_{k^{\prime}}^{\top}+\sum^{K}_{k=1}\mathbf{W}_{k}\Sigma_{{\bm{\gamma}}_{2,k}}\mathbf{W}_{k}^{\top}.$ (7)
Here, $\Sigma_{{\bm{\theta}}}$ is the asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{WCC}}$. $\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$ is the asymptotic covariance between $\widehat{{\bm{\theta}}}_{\text{WCC}}$ and $\widehat{{\bm{\gamma}}}_{1,k}$. $\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$ is the asymptotic covariance between $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{1,k^{\prime}}$; when $k=k^{\prime}$, it is the asymptotic variance of $\widehat{{\bm{\gamma}}}_{1,k}$. $\Sigma_{{\bm{\gamma}}_{2,k}}$ is the asymptotic variance of $\widehat{{\bm{\gamma}}}_{2,k}$.
From Equation (7), $\Sigma_{\text{PS-PPI}}$ is a function of $\{\mathbf{W}_{k}\}=(\mathbf{W}_{1},...,\mathbf{W}_{K})$, based on which we may seek the optimal $\{\mathbf{W}^{\ast}_{k}\}$ such that $\Sigma_{\text{PS-PPI}}(\{\mathbf{W}^{\ast}_{k}\})\preceq\Sigma_{\text{PS-PPI}}(\{\mathbf{W}_{k}\})$ for any $\{\mathbf{W}_{k}\}$, where, for two square matrices $\mathbf{A}$ and $\mathbf{B}$, $\mathbf{A}\preceq\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive semidefinite; $\mathbf{A}\prec\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive definite. Such an optimal set of $\{\mathbf{W}_{k}\}$ has no closed-form and may be computationally expensive to solve. However, when assuming identical $\mathbf{W}_{k}$ across $k$, we have a closed-form solution of $\{\mathbf{W}^{\ast}_{k}\}$, and the resulting $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is guaranteed to be more efficient than $\widehat{{\bm{\theta}}}_{\text{WCC}}$, as we state in Corollary 3.2 and prove in Section S1.3 of the Supplementary Materials.
Corollary 3.2.
Suppose $\mathbf{W}_{k}=\mathbf{W}$ across $k$, then
$\mathbf{W}^{\ast}=\left(\sum^{K}_{j=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,j}}\right)\times\left[\left(\sum^{K}_{j=1}\sum^{K}_{j^{\prime}=1}\Sigma_{{\bm{\gamma}}_{1,j},{\bm{\gamma}}_{1,j^{\prime}}}\right)+\left(\sum^{K}_{j=1}\Sigma_{{\bm{\gamma}}_{2,j}}\right)\right]^{-1}$
leads to an optimal $\Sigma_{\text{PS-PPI}}$. Accordingly, the asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is
$\Sigma_{\text{PS-PPI}}=\Sigma_{{\bm{\theta}}}-\left(\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}\right)\times\left[\left(\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\right)+\left(\sum^{K}_{k=1}\Sigma_{{\bm{\gamma}}_{2,k}}\right)\right]^{-1}\times\left(\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}\right)^{\top}.$ (8)
Furthermore, we have $\Sigma_{\text{PS-PPI}}\preceq\Sigma_{{\bm{\theta}}}$, where $\Sigma_{{\bm{\theta}}}$ is the asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{WCC}}$. $\Sigma_{\text{PS-PPI}}=\Sigma_{{\bm{\theta}}}$ if and only if $\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}=\bm{0}_{d\times d}$.
The $\mathbf{W}^{\ast}$ defined in Corollary 3.2 is controlled by three components: 1) $\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$: the asymptotic covariance between $\widehat{{\bm{\theta}}}_{\text{WCC}}$ and each of the $\widehat{{\bm{\gamma}}}_{1,k}$; 2) $\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$: the asymptotic covariance between any two of the $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{1,k^{\prime}}$; 3) $\Sigma_{{\bm{\gamma}}_{1,k}}$ and $\Sigma_{{\bm{\gamma}}_{2,k}}$: the asymptotic variance of $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{2,k}$ under each missingness pattern. Specifically, a large weight matrix is assigned to $\widehat{\mathbf{\Delta}}_{k}$ when: 1) we have a large covariance between $\widehat{{\bm{\theta}}}_{\text{WCC}}$ and $\widehat{{\bm{\gamma}}}_{1,k}$, which occurs when the imputations faithfully reflect the missing true values; 2) we have small variances of $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{2,k}$, which means that we can construct a precise estimation of zero with small variances; 3) we have small covariances between any pairs of the $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{1,k^{\prime}}$, which means that we can efficiently aggregate across different missingness patterns without introducing too much extra uncertainty.
Because the optimal $\mathbf{W}^{*}$ depends on the variance-covariance matrices, it can be estimated by letting $\widehat{\mathbf{W}}=\left(\sum^{K}_{j=1}\widehat{\Sigma}_{{\bm{\theta}},{\bm{\gamma}}_{1,j}}\right)\times\left[\left(\sum^{K}_{j=1}\sum^{K}_{j^{\prime}=1}\widehat{\Sigma}_{{\bm{\gamma}}_{1,j},{\bm{\gamma}}_{1,j^{\prime}}}\right)+\left(\sum^{K}_{j=1}\widehat{\Sigma}_{{\bm{\gamma}}_{2,j}}\right)\right]^{-1}$.
We present the full algorithmic details in Algorithm 1 to obtain $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ and the corresponding uncertainty estimates.
The theoretical results from Theorem 1 and Corollary 3.2 allow us to compute $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ and quantify its uncertainty in a Z-estimation framework. In particular, $\widehat{{\bm{\theta}}}_{WCC}$, $\widehat{{\bm{\gamma}}}_{1,k}$, $\widehat{{\bm{\gamma}}}_{2,k}$ can be obtained by solving the estimating equations in Equation (3), (5), (6). For uncertainty quantification, $\widehat{\Sigma}_{{\bm{\theta}}}$, $\widehat{\Sigma}_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$, $\widehat{\Sigma}_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$, and $\widehat{\Sigma}_{{\bm{\gamma}}_{2,k}}$ may be calculated by deriving closed-form expressions in a Z-estimation framework based on results in Equation (S1.3), (S1.4), (S1.5), (S1.6), and compute their empirical estimates. We provide more detailed calculations for linear regression and logistic regression as special cases in Section S1.4 in the Supplementary Materials. For general Z-estimation problems without readily computable empirical estimates of the variance-covariance matrices, resampling techniques such as Bootstrap and delete-d Jackknife (efron1982jackknife) can be used. Under certain technical conditions, both Bootstrap and delete-$d$ Jackknife can produce a consistent variance-covariance estimation (shao1989general; shao1989efficiency; cheng2010bootstrap).
It is worth noting that the PS-PPI method can be implemented provided there exists an off-the-shelf package for the target Z-estimation problem with options for weighted analyses based on complete data. Specifically, Equation (3), (5), (6) can be viewed as separate weighted complete-case analyses on different data using the same Z-estimation framework. Hence, no significant additional efforts need to be made to compute $\widehat{{\bm{\theta}}}_{\text{WCC}}$, $\widehat{{\bm{\gamma}}}_{1,k}$, $\widehat{{\bm{\gamma}}}_{2,k}$, as well as variance-covariance matrices $\widehat{\Sigma}_{{\bm{\theta}}}$ and $\widehat{\Sigma}_{{\bm{\gamma}}_{2,k}}$ once a weighted estimation software is available. The major effort for the implementation will be, therefore, to compute $\widehat{\Sigma}_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$ and $\widehat{\Sigma}_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$. As discussed above, this can be implemented by either deriving the specific closed-form based on Equation (S1.4) and (S1.5) in the Supplementary Materials, or applying resampling techniques. Since both $\widehat{\Sigma}_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$ and $\widehat{\Sigma}_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$ are variance-covariance matrices only relevant to the fully-observed data, the resampling will only be conducted among the fully-observed data. In this paper, we use the delete-$1$ Jackknife to estimate the covariances.
3.2 Connections to the existing literature
Connections to the setting when only the outcome is missing.
We show that the PS-PPI procedure subsumes the outcome-based prediction-based inference framework in the existing literature. Specifically, assume that in addition to the fully observed data subset, there is only one missingness pattern with only the outcome, say $Y_{i1}$, missing. Furthermore, assume that the underlying missing mechanism is MCAR, i.e., the propensity score model $\bm{\pi}(\cdot)$ is a constant 2-dimensional vector $(\pi_{1},\pi_{\infty})$ across all the samples, where $\pi_{1}$ is the probability of missing outcome, and $\pi_{\infty}$ is the probability of being fully observed. The PS-PPI estimator under this setting become $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}=\widehat{{\bm{\theta}}}_{\text{CC}}-\widehat{\mathbf{W}}(\widehat{{\bm{\gamma}}}_{1,1}-\widehat{{\bm{\gamma}}}_{2,1})$. Note that here we have the CCA estimate $\widehat{{\bm{\theta}}}_{\text{CC}}$ instead of $\widehat{{\bm{\theta}}}_{\text{WCC}}$, because the weights are the same across all samples. In gronsbell2024another and chen2000unified, their proposed estimator is $\widehat{{\bm{\theta}}}_{\text{Chen}}=\widehat{{\bm{\theta}}}_{\text{CC}}-\widehat{\mathbf{W}}_{\text{Chen}}(\widehat{{\bm{\gamma}}}_{\text{lab}}-\widehat{{\bm{\gamma}}}_{\text{all}})$. Here, $\widehat{{\bm{\theta}}}_{\text{CC}}$ follows the same definition in the present paper and $\widehat{{\bm{\gamma}}}_{\text{lab}}=\widehat{{\bm{\gamma}}}_{1,1}$ with $K=1$ missingness pattern (missing outcome only); $\widehat{{\bm{\gamma}}}_{\text{all}}$ is a variant of $\widehat{{\bm{\gamma}}}_{2,1}$ in this paper, which is estimated by all the data in ${\mathcal{D}}_{\text{obs}}$, instead of only data with ${\mathcal{C}}_{i}=1$; $\widehat{\mathbf{W}}_{\text{Chen}}$ is a tuning parameter matrix estimated from the data. We show the asymptotic equivalence between $\widehat{{\bm{\theta}}}_{\text{Chen}}$ and $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ in Proposition 2.
Proposition 2.
Assume that only the outcome, say $Y_{i1}$, is possibly missing, we have that $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ and $\widehat{{\bm{\theta}}}_{\text{Chen}}$ are asymptotically equivalent in the linear regression (squared error loss) and MCAR setting considered in gronsbell2024another.
Proposition 2 is proved by showing that their influence functions are the same. Note that in $\widehat{{\bm{\theta}}}_{\text{Chen}}$, $\widehat{{\bm{\gamma}}}_{\text{all}}$ is estimated using both partially and fully-observed data. This is different from $\widehat{{\bm{\gamma}}}_{2,1}$ in $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$, which is estimated using only the missing-outcome subset of data. gronsbell2024another argue that this modification guarantees better efficiency than the original PPI procedure (angelopoulos2023prediction). However, when an optimal $\mathbf{W}^{\ast}$ is selected, Proposition 2 indicates that such efficiency improvement is absorbed into the $\mathbf{W}^{\ast}$, and thereby the two estimators become asymptotically equivalent. We elaborate this equivalence using the example of perfect predictions, i.e., the predictions are exactly the true values. Specifically, we have the following remark:
Remark 3.1.
Suppose perfect predictions are available, then for $\widehat{{\bm{\theta}}}_{\text{Chen}}$, $\widehat{\mathbf{W}}_{\text{Chen}}={\bm{I}}$; for $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$, $\widehat{\mathbf{W}}=\widehat{\pi}_{1}{\bm{I}}$.
Remark 3.1 can be proved by deriving the closed-form of the variance-covariance matrices under linear regression and obtaining the closed form of $\Sigma_{\text{PS-PPI}}$. From Remark 3.1, when perfect predictions are available, $\widehat{{\bm{\theta}}}_{\text{Chen}}=\widehat{{\bm{\gamma}}}_{\text{all}}$, which is equivalent to the results from the underlying full data since the predictions are perfect. On the other hand, $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}=\widehat{{\bm{\theta}}}_{\text{CC}}-\widehat{\pi}_{1}(\widehat{{\bm{\gamma}}}_{1,1}-\widehat{{\bm{\gamma}}}_{2,1})=\widehat{\pi}_{\infty}\widehat{{\bm{\theta}}}^{(\infty)}_{\text{CC}}+\widehat{\pi}_{1}\widehat{{\bm{\theta}}}^{(1)}_{\text{CC}}$. Here $\widehat{{\bm{\theta}}}^{(\infty)}_{\text{CC}}$ is the CCA estimate obtained from the complete-case subset; $\widehat{{\bm{\theta}}}^{(1)}_{\text{CC}}$ is the CCA estimate obtained from the underlying full data of the missing-outcome subset of data. We are able to replace $\widehat{{\bm{\gamma}}}_{2,1}$ with $\widehat{{\bm{\theta}}}^{(1)}_{\text{CC}}$ because the perfect predictions make the imputed dataset exactly the same as the underlying full data. As a result, $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ can be viewed as a weighted sum of the full data estimator from the complete-case and the missing-outcome subset of data, respectively; the weights are the proportions of each subset: $\widehat{\pi}_{\infty}$ and $\widehat{\pi}_{1}$. Therefore, $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ can be viewed as an inverse variance weighted version of $\widehat{{\bm{\theta}}}_{\text{Chen}}$.
$\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ also shares core insights with other existing outcome-based prediction-based inference literature (angelopoulos2023PPI++; miao2023assumption; gan2024prediction), although they may not be asymptotically equivalent. In particular, angelopoulos2023PPI++; miao2023assumption; gan2024prediction introduce tunable weights within the estimating equations, and derive similar estimators with guarantees of being more efficient than CCA. All of these methods are based on the outcome-missing setting.
Connections to missing multiple covariates setting.
In addition, the PS-PPI procedure also subsumes the missing covariates scenario studied in kluger2025prediction, which considered two missingness patterns: 1) complete cases (all $p$ variables in $\bm{Y}_{i}$); 2) $s$ different variables in $\bm{Y}_{i}$ are missing simultaneously. Furthermore, the missing mechanism was assumed to be MAR with known probabilities for each missingness pattern. In the above missing covariates setting, we can establish the following asymptotic equivalence between $\widehat{{\bm{\theta}}}_{\text{PTD}}$ proposed in kluger2025prediction and $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$:
Proposition 3.
$\widehat{{\bm{\theta}}}_{\text{PTD}}$ and $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ are asymptotically equivalent.
Connections to semiparametric inference literature.
In the context of semiparametric inference theory, $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ belongs to the class of AIPW estimators, which can be shown by its influence functions derived from the proof of Theorem 1 (detailed in Section S1.5 in the Supplementary Materials). In the context of semiparametric theory, one may be interested in searching for the most efficient estimators within the AIPW class of estimators under specific settings (robins1994estimation; robins1995analysis; chaudhuri2016gmm). However, under general missingness patterns and with an assumption of MAR, the optimal AIPW Z-estimators, despite being the most efficient estimator in integrating the predicted surrogates, suffer from difficulties in implementation as they generally do not have a closed-form and would require iterative computation to solve. We refer readers to robins1994estimation, tsiatis2006semiparametric, and chaudhuri2016gmm for a more in-depth discussion. The proposed $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ in this paper does not aim to guarantee optimal influence functions within the AIPW class estimators. Rather, the major advantage of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is its ability to be easily applied in analysis without significant additional efforts in the implementation. With the rapid development of deep learning and machine learning techniques, highly precise ML predictions become increasingly available. Enabling analysts to easily combine these ML-based prediction surrogates with their statistical models to perform the analysis without additional significant implementation efforts is necessary. Our proposed PS-PPI procedure circumvents the implementation challenges by allowing researchers to apply existing statistical software designed for weighted complete-case analysis to obtain the PS-PPI estimator: each component of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ (i.e., $\widehat{{\bm{\theta}}}_{\text{WCC}},\{\widehat{{\bm{\gamma}}}_{1,k}\},\{\widehat{{\bm{\gamma}}}_{2,k}\}$) can be obtained from weighted complete-case analysis algorithms and the weights can be obtained by resampling techniques (see Algorithm 1).
Connections to multiple imputation.
Multiple imputation based on Rubin’s rule is a well-established methodology that has been widely applied to account for the uncertainty introduced by imputation models to guarantee valid downstream statistical inference (little2019statistical). Multiple imputation often assumes a scenario where the observed data are the only resource, and imputation models are fitted based on these observed data to generate repeated imputations. To guarantee valid inference from these multiply-imputed datasets, multiple imputation imposes assumptions on the imputation models (rubin1987multiple). Such a scenario is different from the scenario assumed in PS-PPI, as well as the broader prediction-based inference literature. In addition to the observed data, current prediction-based inference literature assumes that analysts also have access to some potentially precise predictions generated from models that are trained by external data. Commonly, these external models are ML or even deep learning (DL)-based. Therefore, they are either challenging to manipulate to satisfy any assumptions or beyond analysts’ control. This is the core reason why prediction-based inference literature, including this work, does not impose any assumptions on the prediction models. However, under the MAR missing mechanism, this weak assumption on the quality of the prediction models comes at the expense of requiring the correct specification of the propensity score model. We present a numerical experiment to investigate PS-PPI’s robustness against propensity score misspecification in Section 4 and leave a detailed discussion on propensity score model specification in Section 6.
Algorithm 1 Pseudocode for obtaining $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ and uncertainty estimates.
1:An observed dataset ${\mathcal{D}}_{\text{obs}}=\{(G_{{\mathcal{C}}_{i}}(\bm{Y}_{i}),{\mathcal{C}}_{i})\}^{N}_{i=1}$ with $K$ missingness patterns (${\mathcal{C}}=1,\ldots,K$) and a subset of complete data ($\mathcal{C}$$\infty$); a set of external machine learning prediction models $h_{k}$ imputing missing entries under each missingness pattern; a propensity score models (either known or estimated from the data) $\widehat{\bm{\pi}}(\cdot)\in\mathbb{R}^{K+1}$, with each dimension $\pi_{k}(\cdot)$ representing the probability of ${\mathcal{C}}=k$. A Z-estimation procedure $f({\tt complete\penalty 10000\ data};{\tt weighting\penalty 10000\ function})$ that supports weighted complete-case analysis and returns the estimates and the variance estimates for the parameter of interest $\bm{\theta}$.
2:for k=1,,Kk=1,...,K do
3:  For subjects with ${\mathcal{C}}_{i}=k$, impute the missing values with ML predictions and obtain the ML-imputed dataset ${\mathcal{D}}_{2,k}=\{\bm{Y}^{(k)}_{\text{imp},i}:{\mathcal{C}}_{i}=k\}$.
4:  Compute
$\widehat{{\bm{\gamma}}}_{2,k},\widehat{\Sigma}_{{\bm{\gamma}}_{2,k}}\leftarrow f\left({\mathcal{D}}_{2,k},\pi^{-1}_{k}(\cdot)\right).$
end for
5:Denote the set of completely observed data ${\mathcal{D}}_{\text{CC}}=\{\bm{Y}_{i}:{\mathcal{C}}_{i}=\infty\}$.
6:for $j=1,...,|{\mathcal{D}}_{\text{CC}}|$ do $\triangleright$ Jackknife
7:  Drop the $j$-th completely observed data and obtain ${\mathcal{D}}^{(j)}_{\text{CC}}$.
8:  Compute
$\widehat{{\bm{\theta}}}^{(j)}_{\text{WCC}}\leftarrow f\left({\mathcal{D}}^{(j)}_{\text{CC}},\pi^{-1}_{\infty}(\cdot)\right).$
9:  for k=1,,Kk=1,...,K do
10:   Mask observed values based on $k$-th missingness patterns and impute predictions to obtain the pseudo ML-imputed dataset ${\mathcal{D}}^{(j)}_{1,k}=\{\bm{Y}^{(k)}_{\text{imp},i}:\bm{Y}_{i}\in{\mathcal{D}}^{(j)}_{\text{CC}}\}$.
11:   Compute
$\widehat{{\bm{\gamma}}}^{(j)}_{1,k}\leftarrow f\left({\mathcal{D}}^{(j)}_{1,k},\pi^{-1}_{\infty}(\cdot)\right).$
  end for end for
12:Compute the jackknife estimates of $\widehat{{\bm{\theta}}}_{\text{WCC}},\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{\Sigma}_{{\bm{\theta}}},\widehat{\Sigma}_{{\bm{\gamma}}_{1,k}},\widehat{\Sigma}_{{\bm{\theta}},{\bm{\gamma}}_{1,k}},\widehat{\Sigma}_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$.
[…]
15:Compute $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ from Equation (2): $\widehat{\bm{\theta}}\leftarrow\widehat{{\bm{\theta}}}_{\text{WCC}}-\sum_{k=1}^{K}\widehat{\mathbf{W}}\left(\widehat{{\bm{\gamma}}}_{1,k}-\widehat{{\bm{\gamma}}}_{2,k}\right).$
16:Compute the estimated variance $\widehat{\Sigma}_{\text{PS-PPI}}$ based on Equation (8) by plugging in estimated quantities.
[…]
4 Numerical Experiments
In this section, we conduct simulation studies to show the operating characteristics of the proposed method to corroborate the theoretical predictions of estimation consistency, inferential validity, and efficiency gains. We compare our proposed approach with the following alternatives: 1) unweighted complete-case analysis (“CCA”); 2) inverse probability weighted (IPW) complete-case analysis (“WCCA”), with the weights set at true or estimated propensity scores; 3) multiple imputation (MI) (rubin1976inference), a classical and popular method to deal with datasets with a non-monotone missingness under the MAR assumption. In this work, we use the implementation of MI provided by the MICE 3.16.0333https://github.com/amices/mice package in R (van2011mice); 4) PPI++ (angelopoulos2023PPI++), a recent extension of the prediction-powered inference framework that guarantees better estimation efficiency compared to the CCA approach. For PPI++, fully observed cases and the cases with missing outcomes are used for parameter estimation; the cases with other missingness patterns will be dropped.
4.1 Experimental setup
To demonstrate the operating characteristics of the proposed method under various ML prediction qualities, here we focus on a typical analytic model specified by a multivariate linear regression as follows:
$Y_{i}=\beta_{0}+\beta_{1}X_{i1}+\beta_{2}X_{i2}+\epsilon_{i},$
where the covariates are generated from $X_{i1}=0.1\exp(Z_{i1})+\tau_{i}$, $X_{i2}=\sin(Z_{i2})+\nu_{i}$, with $(Z_{i1},Z_{i2})$ randomly and independently drawn from a bivariate normal distribution with mean zero and a homogeneous marginal variance $\sigma_{z}^{2}$ and Pearson correlation $\rho$. The error terms are independently distributed as $\epsilon_{i}\sim{\mathcal{N}}(0,\sigma^{2})$, $\tau_{i}\sim{\mathcal{N}}(0,\sigma_{\tau}^{2})$, $\nu_{i}\sim{\tt Exponential}(\lambda)$, respectively. To connect with the notation $\bm{Y}_{i}$ introduced in Section 3, we have $\bm{Y}_{i}=(Y_{i},X_{i1},X_{i2},Z_{i1},Z_{i2})$ in this simulation setting. $Z_{i1}$ and $Z_{i2}$ are assumed to be fully observed and not used in the main analysis model. They can be used as inputs to produce ML predictions of possibly missing $(Y_{i},X_{i1},X_{i2})$ in the analysis model of scientific interest. We set ${\mathcal{C}}_{i}=\infty$ to indicate that all $Y_{i},X_{i1},X_{i2}$ are observed. We introduce three distinct missingness patterns in addition to the fully observed data as follows: ${\mathcal{C}}_{i}=1$ indicates that $Y_{i}$ is missing; ${\mathcal{C}}_{i}=2$ indicates that only $X_{i2}$ is missing; ${\mathcal{C}}_{i}=3$ indicates that both $Y_{i}$ and $X_{i1}$ are missing. We assume an MAR mechanism. We define a propensity score model in the form proposed by sun2018inverse as follows:
$\textrm{logit}\{\pi({\mathcal{C}}_{i}|Z_{i},Y_{i,\text{obs}})\}=\begin{cases}-1+0.1X_{i2}+0.1Z_{i1}+0.1X_{i1}X_{i2},&{\mathcal{C}}_{i}=1\\ -1.8-0.2Y_{i}+0.1X_{i1}+0.1Z_{i1}+0.3X_{i1}Y_{i},&{\mathcal{C}}_{i}=2\\ -1.0+0.1X_{i2}+0.2Z_{i1},&{\mathcal{C}}_{i}=3\\ \end{cases},$ (9)
$\pi({\mathcal{C}}_{i}=\infty|Z_{i},Y_{i,\text{obs}})=1-\sum^{3}_{k=1}\pi({\mathcal{C}}_{i}=k|Z_{i},Y_{i,\text{obs}}).$ (10)
To generate simulation data, we first sample $(Z_{i1},Z_{i2})$, then generate $(Y_{i},X_{i1},X_{i2})$ accordingly to obtain the full data. For each subject, we then sample ${\mathcal{C}}_{i}$ from the propensity score model and mask the variables based on the missingness pattern assigned to construct the observed dataset. To enable nimble control of bias and variance of the ML predictions for revealing the performance of the different methods under distinct qualities of ML predictions, instead of an explicit ML prediction, we design the ML predictions by adding additive noise and bias at a pre-specified level onto the unknown real values. More specifically,
$\widehat{X}_{i1}=X_{i1}+b_{i1}+\epsilon_{i1},\widehat{X}_{i2}=X_{i2}+b_{i2}+\epsilon_{i2},\widehat{Y}_{i}=Y_{i}+b_{i3}+\epsilon_{i3}.$
Here both $\epsilon_{i1}$, $\epsilon_{i2}$, and $\epsilon_{i3}$ follow Normal distribution ${\mathcal{N}}(0,\sigma^{2}_{\text{pred}})$, while $b_{i1}$, $b_{i2}$, and $b_{i3}$ follow Exponential distribution ${\tt Exponential}(\lambda_{\text{pred}})$. The two types of error terms represent the noise and bias imposed on the ML predictions, respectively. When the noise and bias are set to 0, perfect ML predictions result.
In our simulation, we fix $\sigma=0.5$, $\sigma_{\tau}=0.3$, $\lambda=0.02$, $\sigma_{z}=0.2$, and $\rho=0.4$. To investigate the impact of the quality of the ML predictions on the performance of our proposed methods, we iterate both $\sigma^{2}_{\text{pred}}$ and $\lambda_{\text{pred}}$ from $\{0,0.2,0.4,0.6,0.8,1.0,1.5,2.0\}$. For propensity scores, we adopt three settings. In the first setting, we assume that the propensity score model is unknown yet correctly specified in terms of both the functional forms and covariates. Under this assumption, we estimate the propensity score by maximizing the likelihood function proposed in sun2018inverse. In the second setting, we again assume an unknown propensity score model in the form proposed by sun2018inverse, while we only include $Z_{1}$ in the propensity score model to make it misspecified. This setting is used to investigate the robustness of the PS-PPI procedure against propensity score model misspecification. In the final setting, we assume that the true propensity score is known and can be directly applied to estimate the parameters of interest $\bm{\beta}$}. We set the sample size $N=5,000$ and repeat the simulation $500$ times. The average number of subjects under each missingness pattern in the 500 simulations is $1,469.9$ for complete cases, and $1,350.7$, $829.5$, $1,349.9$ for ${\mathcal{C}}=1,2,3$, respectively.
4.2 Evaluation metrics
To evaluate the performance of the proposed and the four alternative methods in simulations, we report three key metrics: 1) estimation bias; 2) the coverage rates of the 95% confidence intervals (CIs); 3) the average width of these intervals. Specifically, suppose that $M$ replicates are generated in data simulation, and for the $j$-th replicate, let $\widehat{\beta}^{(j)}$ be the point estimate for a scalar parameter, and $(L^{(j)},U^{(j)})$ its 95% CI. We compute the estimation bias by $\texttt{Bias}=\frac{1}{M}\sum^{M}_{j=1}\left(\widehat{\beta}^{(j)}-\beta_{\text{true}}\right).$ We compute the coverage rate by $\texttt{Coverage Rate}=\frac{1}{M}\sum^{M}_{j=1}I\left(\beta_{\text{true}}\in(L^{(j)},U^{(j)})\right).$ Finally, we compute the average width of the 95% CIs by $\texttt{Avg. 95\% CI Width}=\frac{1}{M}\sum^{M}_{j=1}\left(U^{(j)}-L^{(j)}\right).$
4.3 Results
Figure 2: Performance comparison among the proposed (PS-PPI) and four alternative methods (CCA, WCCA, MI, PPI++). (a): Widths of 95% confidence intervals ($\log_{10}$ transformed); (b): coverage rates of the 95% confidence intervals (red dashed lines indicate 95%); (c): estimation biases for different methods with estimated propensity scores based on correctly specified propensity score models. The X-axis in all the panels represents the prediction noise level used in the simulations. The prediction bias level is fixed at $\lambda_{\text{pred}}=0$.
In the main text, we present the results under varying noise levels with a fixed prediction bias ($\lambda_{\text{pred}}=0$), and estimated propensity scores in Figure 2. Additional experimental results obtained under varying prediction biases, as well as different settings of the propensity score models are provided in Figures S1, S2, S3, S4, and S5, Section S2.1 of the Supplementary Materials. Note that because CCA, WCCA, and MI do not use ML predictions, the coverage rate, the width of the 95% CIs, and the bias remain constant with respect to the quality of the ML predictions.
Coverage rate
Figure 2(b) shows that the CCA approach achieves coverage rates of 51.0% and 81.6% for $\beta_{0}$ and $\beta_{1}$, respectively, which are below the nominal level of 95%. This is expected because CCA does not use propensity scores to adjust for the missing-at-random mechanism that renders the complete cases non-representative of the entire population. WCCA achieves conservative coverage rates of 98.2%, 98.0%, and 96.2% on the three coefficients, respectively. The conservativeness is due to the use of the robust sandwich variance that does not account for the uncertainty from the estimated propensity score model. From Figure S2 in the Supplementary Materials, when the propensity score model is known, WCCA exhibits valid coverages around 95% for the three coefficients. MI achieves coverage rates of 93.2%, 91.6% and 56.6%, respectively for $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$, falling short of the nominal 95% level. Under perfect predictions, PPI++ achieves coverage rates of 71.6% and 89.2% for $\beta_{0}$ and $\beta_{1}$ due to the fact that it does not use the propensity score, similar to the CCA approach. When the quality of the ML predictions degrades, the coverage rates for $\beta_{0}$ and $\beta_{1}$ decrease and converge to levels similar to the CCA approach. PS-PPI, on the other hand, achieves a coverage rate of around 95% when the perfect predictions are available and tends to provide conservative confidence intervals when the quality of the ML predictions deteriorates, similar to the behavior of its base WCCA estimator.
Widths
Based on Figure 2(a), the widths of the 95% CIs for all three coefficients obtained from CCA are 0.027, 0.085, and 0.129, respectively; the widths for WCCA are similar (0.028, 0.087, and 0.130, respectively). MI has 95% CIs of width 0.025, 0.077, and 0.130 for the three coefficients; however, the empirical coverage rates for $\beta_{1}$ and $\beta_{2}$ are below nominal levels. PS-PPI yields uniformly narrower 95% CIs compared to WCCA, the only alternative method with valid coverage rates for all three coefficients. Furthermore, the widths of 95% CIs obtained from PS-PPI are strongly impacted by the quality of the ML predictions. Specifically, PS-PPI tends to produce narrower 95% CIs when ML predictions have better quality. The PPI++ method shows a similar trend, but has wider 95% CIs than PS-PPI, except for the 95% CI of $\beta_{2}$ when the quality of ML predictions becomes mediocre.
Bias
Figure 2(c) shows that CCA, PPI++, and MI fail to achieve 95% nominal coverage (see Figure 2 (b)) primarily due to their biased estimation. For $\beta_{0}$ and $\beta_{1}$, WCCA results in negligible biases of 0.0005 and 0.0006, respectively. On the other hand, PS-PPI results in negligible biases of 0.0007 and 0.0008, even with the worst quality of ML predictions. For the parameters that CCA, PPI++, and MI fail to achieve 95% nominal coverage, they exhibit greater biases. Specifically, MI shows a bias of 0.115 for the estimation of $\beta_{2}$, while CCA results in a bias of -0.043 for $\beta_{1}$. In terms of PPI++, the bias for $\beta_{1}$ increases from -0.023 to -0.043 when the quality of ML predictions degrades.
Similar findings are observed from Figure S1, S2, S3 in the Supplementary Materials. The simulation results demonstrate that under non-monotone missingness patterns with the MAR assumption for the missing mechanism, the proposed PS-PPI approach can produce consistent estimates and valid inference with efficiency dominance over WCCA when the propensity score model is correctly estimated or known. Such consistency and validity in inference hold regardless of the quality of the ML predictions. Furthermore, when precise ML predictions are available, PS-PPI is much more efficient than WCCA. This corroborates our theoretical results.
Robustness against propensity score model misspecification
The simulation results under misspecified propensity score models are shown in Figure S4 and S5 in the Supplementary Materials. Note that WCCA and PS-PPI are the only two methods that are impacted by the propensity score model specification. When the propensity score model is misspecified in terms of the covariates, all the methods show a degree of performance in the width of the 95% CIs similar to the one shown in Figure 2. In terms of the coverage rate, Figure S4 and S5 show that WCCA fails to achieve the nominal 95% coverage rate for $\beta_{0}$ and $\beta_{1}$. PS-PPI achieves the 95% coverage rate when the ML predictions are sufficiently precise, while these coverage rates decline when the quality of ML predictions degrades. However, it can be seen that the coverage rates from PS-PPI still consistently outperform the WCCA baseline. Furthermore, PS-PPI also maintains a higher coverage rate and narrower intervals compared to the PPI++ method, which only utilizes the fully-observed and missing-outcome subsets of data.
5 Application to the All of Us Data
Background
Education attainment is a key social determinant of cardiovascular health, shaping lifestyle behaviors, access to healthcare, and long-term disease outcomes (wang2024association). Among the biological pathways linking social conditions to cardiovascular risk, systemic inflammation (often measured through C-reactive protein (CRP)) plays a central role. CRP has been widely associated with the development of cardiovascular disease (CVD)(ridker2003clinical). In this section, we use data from the All of Us (AoU) program (all2019all) to examine how education levels relate to CRP concentrations, accounting for demographic and clinical covariates, to explore whether disparities in inflammation help explain the socioeconomic gradient in CVD risk (magnani2024educational). AoU was launched in 2018 to establish a large-scale database of electronic health records (EHRs) and biospecimens for hypothesis-free population research. In this study, we analyze data from 314,625 participants who self-reported as male or female and have both EHR and genotype data available in the Curated Data Repository (CDR) version 8.
Analysis model with ML predictions
To investigate the association, we fit the following regression model with the primary covariate (education), adjusting for age, sex, race, smoking status, and the first 5 genetic principal components (PCs):
$\text{CRP}\sim$ $\text{Intercept}+\text{Education}+\text{Age}+\text{Sex}+\text{Race}+\text{Smoking}+\text{PC1}+\cdots+\text{PC5}.$ (11)
To address genetic relatedness, we keep one individual from each kinship group for analysis and use the rest to train ML models (mccaw2024synthetic); this results in the total $314,625$ split into i) 20,278 participants for ML model building, and ii) 294,347 participants for the analytic sample. In the analytical sample, 47,317 subjects are fully observed; Missingness occurs in race, smoking status, education, and CRP (15 patterns with at least one missing entry). Besides the complete cases, only patterns with more than 10,000 samples are kept: 1) missing CRP only (172,685); 2) missing race and CRP (50,040); 3) missing race only (10,934). We drop the remaining patterns with a total of 13,371 samples (4.54% of the analytical sample). This analysis decision simplifies modeling of the probabilities of missingness for 4 instead of 16 patterns with a significant portion of the original analytic sample retained for analysis. Each missing variable is imputed using either gradient boosting decision trees (XGBoost, chen2016xgboost) or random forest, whichever performs better on a validation set. Predictors for imputation include age, sex, genetically inferred ancestry, PCs, and diagnostic events coded by Systematized Medical Nomenclature for Medicine–Clinical Terminology (SNOMED) (stearns2001snomed). We apply the PS models in the form proposed by sun2018inverse (similar to Eq. (9) and (10)) and include all the available variables in each pattern for fitting. We compare against WCCA, or using only the missing outcome pattern for PS-PPI. We conduct sensitivity analysis of PS-PPI without IPW and compare against CCA and its outcome-based variant. For each method, we report coefficient estimates, standard error, 95% CIs, and variance ratio (relative to WCCA or CCA).
Table 1: Summary statistics for different subsets of the All of Us analytic sample used in this study: 1) complete cases ($C$); 2) $C$ + missing outcome only data; 3) $C$ + data with missingness patterns of size $\geqslant$ 10K; 4) all data. For C-reactive protein (CRP), mean, standard deviation, and range are reported. For other variables, the values represent percentages; counts are shown in parentheses. “Missing” rows show counts only. Missingness in smoking status and education levels are due to responses of “Skip”, “Don’t know”, or “Prefer not to answer”; missingness in race is due to responses of “Skip”, “None of these”, or “Prefer not to answer”. Note that age is treated as a continuous variable in the regression model.
$C$ CC+missing CRP CC+patterns of size $\geqslant$ 10K Total
Sample Size 47,317 220,002 280,976 294,347
C-reactive protein (CRP; mg/L)
Mean (SD) 3.9 (5.0) 3.9 (5.0) 4.1 (5.1) 4.1 (5.1)
Range 0.0–26.1 0.0–26.1 0.0–26.1 0.0–26.1
College graduate or higher 50.1 (23,729) 51.6 (113,617) 46.5 (130,688) 46.3 (133,523)
Missing (0) (0) (0) (6,072)
Age group
18–24 0.4 (170) 0.7 (1,464) 0.8 (2,312) 0.8 (2,489)
25–34 4.9 (2,332) 9.4 (20,575) 10.9 (30,497) 10.9 (31,998)
35–44 9.7 (4,583) 13.8 (30,386) 14.9 (41,829) 14.9 (43,882)
45–54 12.1 (5,734) 13.3 (29,193) 14.3 (40,168) 14.3 (42,150)
55–64 20.0 (9,468) 19.4 (42,609) 19.4 (54,384) 19.5 (57,277)
65+ 52.9 (25,030) 43.5 (95,775) 39.8 (111,786) 39.6 (116,551)
Sex at birth
Female 61.5 (29,102) 59.6 (131,070) 60.9 (171,010) 60.6 (178,463)
Male 38.5 (18,215) 40.4 (88,932) 39.1 (109,966) 39.4 (115,884)
Race
Asian 2.3 (1,083) 3.9 (8,666) 3.9 (8,666) 3.9 (8,995)
White 78.8 (37,273) 72.7 (159,848) 72.7 (159,848) 71.8 (164,883)
Black 16.6 (7,862) 20.9 (45,993) 20.9 (45,993) 21.7 (49,708)
Others 2.3 (1,099) 2.5 (5,495) 2.5 (5,495) 2.6 (5,968)
Missing (0) (0) (60,974) (64,793)
Have ever smoked $\geqslant$100 cigarettes 45.4 (21,463) 42.1 (92,627) 40.3 (113,103) 40.3 (115,628)
Missing (0) (0) (0) (7,731)
Depression 42.1 (19,932) 26.5 (58,217) 26.2 (73,746) 26.2 (77,196)
Hypertension 63.2 (29,909) 42.4 (93,384) 41.3 (115,996) 41.2 (121,409)
Diabetes 30.6 (14,467) 18.5 (40,739) 19.5 (54,732) 19.5 (57,469)
Anxiety 45.6 (21,596) 29.0 (63,872) 28.6 (80,450) 28.6 (84,152)
Obesity 45.0 (21,289) 28.8 (63,268) 29.5 (82,849) 29.4 (86,515)
Table 2: All of Us based coefficient estimates and 95% CIs for education level in the regression model in Eq. (11). Smaller variance ratios represent larger efficiency gains.
Method $\widehat{\beta}_{\textrm{edu}}$ Standard Error 95% CIs Variance Ratio
without IPW CCA (baseline) -1.187 0.048 [1.281,1.093][-1.281,-1.093] 1
PS-PPI (outcome missing only) -0.946 0.045 [1.035,0.857][-1.035,-0.857] 0.888
PS-PPI -1.091 0.042 [1.174,1.008][-1.174,-1.008] 0.778
with IPW WCCA (baseline) -1.192 0.053 [1.295,1.088][-1.295,-1.088] 1
PS-PPI (outcome missing only) -0.928 0.050 [1.027,0.830][-1.027,-0.830] 0.903
PS-PPI -1.110 0.047 [1.202,1.017][-1.202,-1.017] 0.796
Results
Table 1 shows the summary statistics of the included variables and several key conditions across different data subsets. Specifically, we compute the summary statistics for 1) the complete-case subset; 2) the subset consisting of complete cases and outcome-missing-only cases; 3) the subset of complete cases and all missingness patterns with more than 10,000 samples; and 4) the full dataset. The characteristics of several variables, including both the outcome and the covariate of interest, differ across these subsets. Specifically, in the subset with only complete cases and outcome-missing cases, the sample mean of the CRP level (in mg/L) is 3.9, compared with 4.1 in the full data. Similarly, the proportion of participants with a college degree or higher is 50.1% and 51.6% for the complete-case and complete-case plus outcome-missing subsets, respectively, both higher than that in the full dataset (46.3%). These discrepancies indicate that CCA and outcome-based prediction-based inference methods rely on subsets with distributions shifted from the overall population, thereby making the result less generalizable to the target population. Even if a method that uses the cases with complete or outcome-missing-only data can be modified to account for such distributional differences using propensity scores for outcome missingness, participants with other missingness patterns and the associated information are not used. In contrast, for the subset including complete cases and all patterns with more than 10,000 samples, the sample mean of CRP (4.1) matches that of the full data, and the proportion of college graduates (46.5%) closely approximates the full-sample value (46.3%). This demonstrates that the proposed PS-PPI method enables inference based on data subsets that are more representative of the full data distribution, thereby improving the generalizability of the inferential results under data missingness.
In terms of the statistical analysis, from Table 2, all methods show that a higher education level is significantly associated with a lower CRP level, adjusting for age, sex, race, smoking status, and the first 5 genetic PCs. Specifically, the WCCA estimate for ${\beta}_{\textrm{edu}}$ is 1.192-1.192 (95% CI: [-1.295, -1.088]). After integrating additional data from the three retained missingness patterns (each of size $\geqslant$ 10K) and ML imputations, the PS-PPI achieves a 20.4% variance reduction relative to WCCA, resulting in an estimate of 1.110-1.110 (95 % CI: [-1.202, -1.017]). We observe a smaller magnitude of variance reduction (9.7%) when integrating additional data but with only CRP missing and predicted. As a sensitivity analysis, without IPW, the CCA estimate is 1.187-1.187 (95% CI: [-1.281, -1.093]), similar to the WCCA estimate, indicating a small degree of selection bias of having CRP test results. The percent variance reduction of PS-PPI with only complete plus missing CRP data and PS-PPI with data from the three retained missingness patterns is 11.2% and 22.2%, respectively, favoring PS-PPI, which integrates more data and ML predictions.
6 Conclusion and Discussion
This paper introduces a unified framework for conducting valid statistical inference using machine learning imputations in datasets characterized by general, multiple different missing data patterns under the MAR assumption. By leveraging the concept of missingness patterns and building upon the theoretical foundations of AIPW estimators, our work significantly extends the applicability of prediction-based inference beyond the more restrictive MCAR assumption and simple missing data structures, such as missing outcomes or missing multiple covariates. A key practical advantage of our proposed methods is their adaptability to any Z-estimation framework, allowing researchers to utilize existing statistical software for weighted complete-case analysis with minimal additional implementation effort. The method introduced here can integrate individuals across diverse missingness patterns, hence maximizing data utility and inclusiveness while ensuring accuracy.
The significance of this development lies in the theoretical foundation it provides to potentially unlock more robust and reliable insights from the vast and often messy datasets prevalent in modern scientific inquiry. As machine learning models become increasingly sophisticated in their predictive capabilities, their use for imputation is a natural and appealing step. However, without rigorous inferential correction, analyses based on such imputed data risk biased conclusions. PS-PPI provides a principled approach to harness the power of ML predictions while maintaining statistical validity, which is crucial in fields ranging from biomedicine and public health to economics and social sciences, where missing data are ubiquitous and decisions based on data analysis can have profound consequences. We provide a theoretical guarantee of improved efficiency over weighted complete-case analysis, making it an attractive option especially when ML predictions are of reasonable quality.
Despite these important contributions, our work has several limitations that inform avenues for future research. First, a central assumption of our methodology is the correct specification of the propensity score model, which governs the probability of observing different missingness patterns. This assumption is the price we pay in the challenging scenario where prediction-based inference is most needed: analysts typically have minimal or no control over the prediction models. As such, assumptions on the prediction models need to be avoided, and the correct specification of the propensity score model comes at a price. Although we demonstrate valid inference when propensity scores are known or consistently estimated from a correctly specified model, results from Figure S4 and S5 in the Supplementary Materials show that propensity score model misspecification will lead to biased estimation when the ML predictions are not sufficiently precise. It is worth noting that results from Figure S4 and S5 in the Supplementary Materials show that PS-PPI achieves the 95% nominal coverage rate when the ML predictions are precise enough (or perfect). Future work should focus on mitigating the effect of propensity score model misspecification. Second, another practical limitation of PS-PPI is that it imposes assumptions on the minimum size of data within each missingness pattern to guarantee the learnability of the parameter of interest. For example, to fit a linear regression analysis model with $p$ parameters (including the intercept), at least $p+1$ data points are needed. In this case, PS-PPI requires each missingness pattern within the data to contain at least $p+1$ data points, and if this requirement is not satisfied, then data with that missingness pattern have to be dropped from the dataset. Third, while PS-PPI provides a data-driven approach for selecting tuning parameters ($\mathbf{W}_{k}$) that guarantees efficiency gains, the assumption of identical $\mathbf{W}_{k}$ across strata was made for simplicity and ease of implementation. Future work could explore more adaptive stratum-specific weight selection to potentially achieve even greater efficiency, although this may involve more complex optimization procedures. Fourth, the number of possible missingness patterns can grow exponentially with the number of variables. Although in many practical applications, the number of observed patterns is manageable, scenarios with a very large number of distinct patterns could pose computational challenges. Future work can improve the utility of the framework by developing adaptive procedures to use data from a small number of the most informative missingness patterns. Fifth, while our framework is versatile within Z-estimation problems, its extension to other inferential paradigms or specific complex data types, such as longitudinal or spatial data where the temporal and spatial dependency of missingness is critical, or high-dimensional settings where variable selection and inference are performed simultaneously, presents fertile ground for future research. The current work also focuses on the MAR assumption. Extending these prediction-powered inference ideas to handle missing not at random (MNAR) scenarios, while notoriously challenging due to identifiability issues, would be a significant theoretical and practical advancement. We leave these directions for future work.
Acknowledgment
We gratefully acknowledge All of Us participants for their contributions, without whom this research would not have been possible. We also thank the National Institutes of Health’s All of Us Research Program for making available the participant data examined in this study. The authors acknowledge the use of Large language models (LLM)-assisted technologies by GPT-5 to refine the language and grammar. All LLM suggestions were evaluated by the first and the last authors prior to their incorporation into the manuscript. LLMs were strictly limited to proposing edits to the initial drafts and were not used to generate any new materials.
Supplementary Materials
Supplementary Materials contain a glossary of notation, proofs, supporting figures, longer derivations, and additional results. Code to reproduce simulations and data analysis will be made available online at https://github.com/chenxran/ps-ppi.
Supplementary Materials
We first provide a glossary for key notation used in the Main Paper.
Symbol Description
$\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ The PS-PPI estimator proposed in this work.
$\widehat{{\bm{\theta}}}_{\text{Full}}$ An estimator obtained from the fully-observed data.
$\widehat{{\bm{\theta}}}_{\text{WCC}}$ An estimator from complete-case analysis with inverse probability weighting (IPW).
$\widehat{{\bm{\theta}}}_{\text{PP}}$ An estimator from outcome-based prediction-powered inference (angelopoulos2023prediction).
$\widehat{{\bm{\theta}}}_{\text{CC}}$ An estimator from unweighted complete-case analysis.
$\widehat{{\bm{\theta}}}_{\text{Chen}}$ An outcome-based PPI++ style estimator from gronsbell2024another; chen2000unified that guarantees to be more efficient than $\widehat{{\bm{\theta}}}_{\text{CC}}$.
$\widehat{{\bm{\theta}}}_{\text{PTD}}$ An estimator from kluger2025prediction for missing multiple covariates with nonuniform sampling weights.
$\widehat{\mathbf{\Delta}}_{k}$ Estimates of zero obtained from the $k$-th missingness patterns; $\widehat{\mathbf{\Delta}}=\widehat{{\bm{\gamma}}}_{1,k}-\widehat{{\bm{\gamma}}}_{2,k}$.
$\widehat{{\bm{\gamma}}}_{1,k}$ The solution of Equation (5), which use complete-case data, with all entries supposed to be missing in the $k$-th missingness patterns being substituted with predictions generated from $h_{k}$.
$\widehat{{\bm{\gamma}}}_{2,k}$ The solution of Equation (6), which use data with $k$-th missingness patterns, with all missing values imputed by predictions generated from $h_{k}$.
$\widehat{\mathbf{W}}_{k}$ Weight matrices of dimension $d\times d$ that may depend on data. In this work, Corollary 3.2 is used to determine $\widehat{\mathbf{W}}_{k}$.
${\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}})$ The estimating equation for $\widehat{{\bm{\theta}}}_{\text{Full}}$ with the $i$-th data plugged in. It is a function of $\bm{\theta}$} and can be abbreviated as ${\bm{\psi}}^{(i)}$. When the true parameter $\bm{\theta}$$^{\ast}$ is plugged in, i.e., ${\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}}^{\ast})$, it can be abbreviated as ${\bm{\psi}}^{\ast(i)}$.
${\bm{\psi}}^{(i)}_{\text{WCC}}$ The estimating equation for $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ with the $i$-th data point plugged in.
${\bm{\psi}}^{(i)}_{1,k}$ The estimating equation for $\widehat{{\bm{\gamma}}}_{1,k}$ with the $i$-th data point plugged in.
${\bm{\psi}}^{(i)}_{2,k}$ The estimating equation for $\widehat{{\bm{\gamma}}}_{2,k}$ with the $i$-th data point plugged in.
${\bm{m}}^{(i)}$ The estimating equation of the propensity score model with the $i$-th data point plugged in.
$\Sigma_{{\bm{\theta}}}$ The asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{WCC}}$.
$\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$ The asymptotic covariance between $\widehat{{\bm{\theta}}}_{\text{WCC}}$ and $\widehat{{\bm{\gamma}}}_{1,k}$.
$\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$ The asymptotic covariance between $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{1,k^{\prime}}$, $k,k^{\prime}\in\{1,...,K\}$. When $k=k^{\prime}$, it becomes the asymptotic variance of $\widehat{{\bm{\gamma}}}_{1,k}$.
$\Sigma_{{\bm{\gamma}}_{2,k}}$ The asymptotic variance of $\widehat{{\bm{\gamma}}}_{2,k}$.
$\Sigma_{\text{PS-PPI}}$ The asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$.
Appendix S1 Proofs
S1.1 Proof of Proposition 1
Proof.
Under the $k$-th missingness pattern, from Equation (5) and Assumption 1 we have:
$\;\mathbb{E}_{(\bm{Y},{\mathcal{C}})}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}\left(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}}^{\ast}_{k}\right)\right]$
$=\mathbb{E}_{\bm{Y}}\left[{\bm{\psi}}\left(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}}^{\ast}_{k}\right)\mathbb{E}_{({\mathcal{C}}|\bm{Y})}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}\right]\right]$
$=\mathbb{E}_{\bm{Y}}\left[{\bm{\psi}}\left(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}}^{\ast}_{k}\right)\right]=\bm{0}.$
Similarly, from Equation (6) we have:
$\quad\;\mathbb{E}_{(\bm{Y},{\mathcal{C}})}\left[\frac{I({\mathcal{C}}_{i}=k)}{\pi_{j}(G_{k}(\bm{Y}_{i}))}{\bm{\psi}}\left(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}}^{\ast}_{k}\right)\right]$
$=\mathbb{E}_{\bm{Y}}\left[{\bm{\psi}}\left(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}}^{\ast}_{k}\right)\mathbb{E}_{({\mathcal{C}}|\bm{Y})}\left[\frac{I({\mathcal{C}}_{i}=k)}{\pi_{j}(G_{k}(\bm{Y}_{i}))}\right]\right]$
$=\mathbb{E}_{\bm{Y}}\left[{\bm{\psi}}\left(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}}^{\ast}_{k}\right)\right]=\bm{0}.$
S1.2 Proof of Theorem 1
To prove Theorem 1, we first stack the estimating equations and obtain:
$\bm{0}=\sum^{N}_{i=1}\bm{\Psi}^{(i)}=\begin{bmatrix}\sum^{N}_{i=1}{\bm{\psi}}_{\text{WCC}}^{(i)}\\[6.0pt] \sum^{N}_{i=1}{\bm{\psi}}_{1,1}^{(i)}\\[2.0pt] \vdots\\[2.0pt] \sum^{N}_{i=1}{\bm{\psi}}^{(i)}_{1,K}\\[6.0pt] \sum^{N}_{i=1}{\bm{\psi}}^{(i)}_{2,1}\\[2.0pt] \vdots\\[2.0pt] \sum^{N}_{i=1}{\bm{\psi}}^{(i)}_{2,K}\\[6.0pt] \sum^{N}_{i=1}{\bm{m}}^{(i)}\end{bmatrix}$.
[…]
Assumption 2.
We assume that $\bm{\Psi}$ is measurable vector-valued function and for every $\bm{\vartheta}_{1}$ and $\bm{\vartheta}_{2}$ in a neighborhood of $\bm{\vartheta}^{\ast}$,
[…]
The proof sketch of the Theorem 1 is that: 1) we derive the asymptotic linear expansion of the stacked estimator $\widehat{\bm{\vartheta}}$; 2) we show that $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is a linear transformation of the elements in $\widehat{\bm{\vartheta}}$, thereby obtaining the asymptotic linear expansion of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$.
Proof.
By working on the stacked estimating equations, we have the following asymptotic linear expansion:
$\sqrt{N}(\widehat{\bm{\vartheta}}-\bm{\vartheta}^{\ast})=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}-\dot{\bm{\Psi}}^{-1}\bm{\Psi}^{\ast(i)}+o_{p}(1),$
where
$\dot{\bm{\Psi}}=\mathbb{E}\Bigl[\tfrac{\partial\bm{\Psi}^{(i)}}{\partial\bm{\vartheta}}\Bigr|_{\bm{\vartheta}=\bm{\vartheta}^{\ast}}\Bigr]=\mathbb{E}\!\left[\,\begin{bmatrix}\frac{\partial{\bm{\psi}}^{(i)}_{\text{WCC}}}{\partial{\bm{\theta}}}&\frac{\partial{\bm{\psi}}^{(i)}_{\text{WCC}}}{\partial{\bm{\gamma}}_{1,1}}&\cdots&\frac{\partial{\bm{\psi}}^{(i)}_{\text{WCC}}}{\partial{\bm{\gamma}}_{1,K}}&\frac{\partial{\bm{\psi}}^{(i)}_{\text{WCC}}}{\partial{\bm{\gamma}}_{2,1}}&\cdots&\frac{\partial{\bm{\psi}}^{(i)}_{\text{WCC}}}{\partial{\bm{\gamma}}_{2,K}}&\frac{\partial{\bm{\psi}}^{(i)}_{\text{WCC}}}{\partial{\bm{\omega}}}\\[6.0pt] \frac{\partial{\bm{\psi}}^{(i)}_{1,1}}{\partial{\bm{\theta}}}&\frac{\partial{\bm{\psi}}^{(i)}_{1,1}}{\partial{\bm{\gamma}}_{1,1}}&\cdots&\frac{\partial{\bm{\psi}}^{(i)}_{1,1}}{\partial{\bm{\gamma}}_{1,K}}&\frac{\partial{\bm{\psi}}^{(i)}_{1,1}}{\partial{\bm{\gamma}}_{2,1}}&\cdots&\frac{\partial{\bm{\psi}}^{(i)}_{1,1}}{\partial{\bm{\gamma}}_{2,K}}&\frac{\partial{\bm{\psi}}^{(i)}_{1,1}}{\partial{\bm{\omega}}}\\[2.0pt] \vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots\\[2.0pt] \frac{\partial{\bm{m}}^{(i)}}{\partial{\bm{\theta}}}&\frac{\partial{\bm{m}}^{(i)}}{\partial{\bm{\gamma}}_{1,1}}&\cdots&\frac{\partial{\bm{m}}^{(i)}}{\partial{\bm{\gamma}}_{1,K}}&\frac{\partial{\bm{m}}^{(i)}}{\partial{\bm{\gamma}}_{2,1}}&\cdots&\frac{\partial{\bm{m}}^{(i)}}{\partial{\bm{\gamma}}_{2,K}}&\frac{\partial{\bm{m}}^{(i)}}{\partial{\bm{\omega}}}\end{bmatrix}_{\bm{\vartheta}=\bm{\vartheta}^{\ast}}\,\right].$
To obtain the inverse of $\dot{\bm{\Psi}}$, we first introduce the following notation:
$\dot{{\bm{\psi}}}_{\infty}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{\text{WCC}}}{\partial{\bm{\theta}}}\Bigr|_{({\bm{\theta}},{\bm{\omega}})=({\bm{\theta}}^{\ast},{\bm{\omega}}^{\ast})}\right],\penalty 10000\ \dot{{\bm{\psi}}}_{k}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{1,k}}{\partial{\bm{\gamma}}_{1,k}}\Bigr|_{({\bm{\gamma}}_{1,k},{\bm{\omega}})=({\bm{\gamma}}^{\ast}_{k},{\bm{\omega}}^{\ast})}\right]=\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{2,k}}{\partial{\bm{\gamma}}_{2,k}}\Bigr|_{({\bm{\gamma}}_{2,k},{\bm{\omega}})=({\bm{\gamma}}^{\ast}_{k},{\bm{\omega}}^{\ast})}\right],$
$\dot{{\bm{\psi}}}_{\infty}^{\bm{\omega}}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{\text{WCC}}}{\partial{\bm{\omega}}}\Bigr|_{({\bm{\theta}},{\bm{\omega}})=({\bm{\theta}}^{\ast},{\bm{\omega}}^{\ast})}\right],\penalty 10000\ \dot{{\bm{\psi}}}^{\bm{\omega}}_{1,k}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{1,k}}{\partial{\bm{\omega}}}\Bigr|_{({\bm{\gamma}}_{1,k},{\bm{\omega}})=({\bm{\gamma}}^{\ast}_{k},{\bm{\omega}}^{\ast})}\right],$
$\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,k}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{2,k}}{\partial{\bm{\omega}}}\Bigr|_{({\bm{\gamma}}_{2,k},{\bm{\omega}})=({\bm{\gamma}}^{\ast}_{k},{\bm{\omega}}^{\ast})}\right],\penalty 10000\ \mathbf{M}=\mathbb{E}\left[\frac{\partial{\bm{m}}^{(i)}}{\partial{\bm{\omega}}}\Bigr|_{{\bm{\omega}}={\bm{\omega}}^{\ast}}\right].$
Note that according to Proposition 1, $\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{1,k}}{\partial{\bm{\gamma}}_{1,k}}\Bigr|_{({\bm{\gamma}}_{1,k},{\bm{\omega}})=({\bm{\gamma}}^{\ast}_{k},{\bm{\omega}}^{\ast})}\right]=\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{2,k}}{\partial{\bm{\gamma}}_{2,k}}\Bigr|_{({\bm{\gamma}}_{2,k},{\bm{\omega}})=({\bm{\gamma}}^{\ast}_{k},{\bm{\omega}}^{\ast})}\right]$, therefore we are able to use a single notation to represent both. However, when taking derivative on $\bm{\omega}$} instead of ${\bm{\gamma}}_{1,k}$ or ${\bm{\gamma}}_{2,k}$, there is no guarantee that $\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{1,k}}{\partial{\bm{\omega}}}\Bigr|_{({\bm{\gamma}}_{1,k},{\bm{\omega}})=({\bm{\gamma}}^{\ast}_{k},{\bm{\omega}}^{\ast})}\right]$ and $\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{2,k}}{\partial{\bm{\omega}}}\Bigr|_{({\bm{\gamma}}_{2,k},{\bm{\omega}})=({\bm{\gamma}}^{\ast}_{k},{\bm{\omega}}^{\ast})}\right]$ are the same. As a result, we use two different notation to represent the two.
Then, by noticing that all the non-diagonal elements except those in the last column are $\bm{0}$, we have:
$\dot{\bm{\Psi}}=\begin{bmatrix}\dot{{\bm{\psi}}}_{\infty}&\bm{0}&\cdots&\dot{{\bm{\psi}}}_{\infty}^{{\bm{\omega}}}\\ \bm{0}&\dot{{\bm{\psi}}}_{1}&\cdots&\dot{{\bm{\psi}}}^{{\bm{\omega}}}_{1,1}\\ \vdots&\vdots&\ddots&\vdots\\ \bm{0}&\bm{0}&\cdots&\mathbf{M}\end{bmatrix}=\begin{bmatrix}\mathbf{J}_{1}&\mathbf{J}_{2}\\ \bm{0}&\mathbf{M}\end{bmatrix}$
where $\mathbf{J}_{1}=\text{diag}(\dot{{\bm{\psi}}}_{\infty},\dot{{\bm{\psi}}}_{1},\cdots,\dot{{\bm{\psi}}}_{K},\dot{{\bm{\psi}}}_{1},\cdots\dot{{\bm{\psi}}}_{K})$, $\mathbf{J}_{2}=(\dot{{\bm{\psi}}}_{\infty}^{\bm{\omega}},\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,1},\cdots,\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,K},\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,1},\cdots\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,K})^{\top}$. From lu2002inverses, the inverse of such a upper triangle $2\times 2$ block matrix $\dot{\bm{\Psi}}$ is:
$\dot{\bm{\Psi}}^{-1}=\begin{bmatrix}\dot{{\bm{\psi}}}^{-1}_{\infty}&\bm{0}&\cdots&-\dot{{\bm{\psi}}}_{\infty}^{-1}\dot{{\bm{\psi}}}_{\infty}^{{\bm{\omega}}}\mathbf{M}^{-1}\\[6.0pt] \bm{0}&\dot{{\bm{\psi}}}_{1}^{-1}&\cdots&-\dot{{\bm{\psi}}}_{1}^{-1}\dot{{\bm{\psi}}}_{1,1}^{{\bm{\omega}}}\mathbf{M}^{-1}\\[2.0pt] \vdots&\vdots&\ddots&\vdots\\[2.0pt] \bm{0}&\bm{0}&\cdots&\mathbf{M}^{-1}\end{bmatrix}$
From van2000asymptotic and Assumption 2, $\sqrt{N}(\widehat{\bm{\vartheta}}-\bm{\vartheta}^{\ast})$ is asymptotically normal. To show that $\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})$ is asymptotically normal, we want to show that it is $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is a linear combination of the elements in $\widehat{\bm{\vartheta}}$. We know that $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}=\widehat{{\bm{\theta}}}_{\text{WCC}}-\sum^{K}_{k=1}\widehat{\mathbf{W}}_{k}(\widehat{{\bm{\gamma}}}_{1,k}-\widehat{{\bm{\gamma}}}_{2,k})=\widehat{{\bm{c}}}^{\top}\widehat{\bm{\vartheta}}$, where $\widehat{{\bm{c}}}=(\mathbf{I},-\widehat{\mathbf{W}}_{1},\cdots,-\widehat{\mathbf{W}}_{K},\widehat{\mathbf{W}}_{1},\cdots,\widehat{\mathbf{W}}_{K},\bm{0})^{\top}$. On the other hand, since the true parameter of interest $\bm{\theta}$$^{\ast}$ is the first element in $\bm{\vartheta}^{\ast}$, we can rewrite it as ${\bm{\theta}}^{\ast}={\bm{\theta}}^{\ast}-\sum^{K}_{k=1}\widehat{\mathbf{W}}_{k}{\bm{\gamma}}^{\ast}_{k}+\sum^{K}_{k=1}\widehat{\mathbf{W}}_{k}{\bm{\gamma}}^{\ast}_{k}=\widehat{{\bm{c}}}^{\top}\bm{\vartheta}^{\ast}$. Define ${\bm{c}}=(\mathbf{I},-\mathbf{W}_{1},\cdots,-\mathbf{W}_{K},\mathbf{W}_{1},\cdots,\\ \mathbf{W}_{K},\bm{0})^{\top}$; since $\widehat{\mathbf{W}}_{k}\overset{p}{\rightarrow}\mathbf{W}_{k}$, we have $\widehat{{\bm{c}}}\overset{p}{\rightarrow}{\bm{c}}$. The asymptotic linear expansion of $\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})$ becomes:
$\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})=\widehat{{\bm{c}}}^{\top}\sqrt{N}(\widehat{\bm{\vartheta}}-\bm{\vartheta}^{\ast})=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}-{\bm{c}}^{\top}\dot{\bm{\Psi}}^{-1}\bm{\Psi}^{\ast(i)}+o_{p}(1).$
Hence, $\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})$ is asymptotically normal. We now derive the form of its asymptotic variance. First,
${\bm{c}}^{\top}\dot{\bm{\Psi}}^{-1}=\begin{bmatrix}\dot{{\bm{\psi}}}_{\infty}^{-1}\\[6.0pt] -\mathbf{W}_{1}\dot{{\bm{\psi}}}^{-1}_{1}\\[6.0pt] \vdots\\ -\mathbf{W}_{1}\dot{{\bm{\psi}}}^{-1}_{K}\\[6.0pt] \mathbf{W}_{1}\dot{{\bm{\psi}}}^{-1}_{1}\\[6.0pt] \cdots\\ \mathbf{W}_{K}\dot{{\bm{\psi}}}^{-1}_{K}\\[6.0pt] \left(\dot{{\bm{\psi}}}_{\infty}^{-1}\dot{{\bm{\psi}}}_{\infty}^{\bm{\omega}}-\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}(\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,k}-\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,k})\right)\mathbf{M}^{-1}\end{bmatrix}^{\top}$
Denote ${\bm{\psi}}^{(i)}={\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}})$, and ${\bm{\psi}}^{(i)}_{k}={\bm{\psi}}(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}})$; similarly we have ${\bm{\psi}}^{\ast(i)}={\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}}^{\ast})$ and ${\bm{\psi}}^{\ast(i)}_{k}={\bm{\psi}}(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}}_{k}^{\ast})$. Since $\widehat{\mathbf{W}}_{k}\overset{p}{\rightarrow}\mathbf{W}_{k}$, we have:
$\;\;\;\;\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})$
$=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}-{\bm{c}}^{\top}\dot{\bm{\Psi}}^{-1}\bm{\Psi}^{\ast(i)}+o_{p}(1)$
$=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}\left[-\dot{{\bm{\psi}}}_{\infty}^{-1}{\bm{\psi}}^{\ast(i)}_{\text{WCC}}+\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}({\bm{\psi}}^{\ast(i)}_{1,k}-{\bm{\psi}}^{\ast(i)}_{2,k})\right$.
$\;\;\;+\left.\left(\dot{{\bm{\psi}}}_{\infty}^{-1}\dot{{\bm{\psi}}}_{\infty}^{\bm{\omega}}-\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}(\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,k}-\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,k})\right)\mathbf{M}^{-1}{\bm{m}}^{\ast(i)}\right]+o_{p}(1).$ (S1.1)
Denote ${\bm{\Gamma}}^{(i)}=-\dot{{\bm{\psi}}}_{\infty}^{-1}{\bm{\psi}}^{(i)}_{\text{WCC}}+\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}({\bm{\psi}}^{(i)}_{1,k}-{\bm{\psi}}^{(i)}_{2,k})$. Note that this is a function of $\bm{\vartheta}$. Then we have
$\dot{{\bm{\Gamma}}}_{\bm{\omega}}:=\mathbb{E}\left[\frac{\partial{\bm{\Gamma}}^{(i)}}{\partial{\bm{\omega}}}\Bigr|_{\bm{\vartheta}=\bm{\vartheta}^{\ast}}\right]=\dot{{\bm{\psi}}}^{-1}\dot{{\bm{\psi}}}^{\bm{\omega}}-\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}(\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,k}-\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,k}).$
As a result, from Eq. (S1.1) we can obtain:
$\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}\left({\bm{\Gamma}}^{\ast(i)}+\dot{{\bm{\Gamma}}}_{\bm{\omega}}\mathbf{M}^{-1}{\bm{m}}^{\ast(i)}\right)+o_{p}(1),$
and therefore the $i$-th influence function of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is
$\varphi_{i}={\bm{\Gamma}}^{\ast(i)}+\dot{{\bm{\Gamma}}}_{\bm{\omega}}\mathbf{M}^{-1}{\bm{m}}^{\ast(i)}.$
Accordingly, $\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}})\overset{d}{\rightarrow}{\mathcal{N}}(0,\text{Var}(\varphi_{i}))$.
$\text{Var}(\varphi_{i})=\mathbb{E}\left[\left({\bm{\Gamma}}^{\ast(i)}+\dot{{\bm{\Gamma}}}_{\bm{\omega}}\mathbf{M}^{-1}{\bm{m}}^{\ast(i)}\right)\left({\bm{\Gamma}}^{\ast(i)}+\dot{{\bm{\Gamma}}}_{\bm{\omega}}\mathbf{M}^{-1}{\bm{m}}^{\ast(i)}\right)^{\top}\right]$ (S1.2)
S1.3 Proof of Corollary 3.1
Proof.
[…]
${\bm{\Gamma}}^{\ast(i)}$ $=-\dot{{\bm{\psi}}}_{\infty}^{-1}{\bm{\psi}}^{\ast(i)}_{\text{WCC}}+\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}({\bm{\psi}}^{\ast(i)}_{1,k}-{\bm{\psi}}^{\ast(i)}_{2,k})$
$=-\underbrace{\dot{{\bm{\psi}}}_{\infty}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}^{\ast(i)}}_{\mathbf{A}^{(i)}}+\underbrace{\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}-\frac{I({\mathcal{C}}_{i}=k)}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\right){\bm{\psi}}^{\ast(i)}_{k}}_{\mathbf{B}^{(i)}}.$
Correspondingly, the asymptotic variance becomes $\text{Var}({\bm{\Gamma}}^{\ast(i)})=\mathbb{E}\left[{\bm{\Gamma}}^{\ast(i)}{\bm{\Gamma}}^{\ast(i)\top}\right]=\mathbb{E}[\mathbf{A}^{(i)}\mathbf{A}^{(i)\top}-\mathbf{A}^{(i)}\mathbf{B}^{(i)\top}-\mathbf{B}^{(i)}\mathbf{A}^{(i)\top}+\mathbf{B}^{(i)}\mathbf{B}^{(i)\top}]$. Here,
$\mathbf{A}^{(i)}\mathbf{A}^{(i)\top}$ $=\dot{{\bm{\psi}}}_{\infty}^{-1}{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}\dot{{\bm{\psi}}}_{\infty}^{-1\top}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi^{2}_{\infty}(\bm{Y}_{i})},$
$\mathbf{A}^{(i)}\mathbf{B}^{(i)\top}$ $=\sum^{K}_{k=1}\dot{{\bm{\psi}}}_{\infty}^{-1}{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}_{k}\dot{{\bm{\psi}}}^{-1\top}_{k}\mathbf{W}^{\top}_{k}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}-\frac{I({\mathcal{C}}_{i}=k)}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\right)$
$=\sum^{K}_{k=1}\dot{{\bm{\psi}}}_{\infty}^{-1}{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}_{k}\dot{{\bm{\psi}}}^{-1\top}_{k}\mathbf{W}^{\top}_{k}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi^{2}_{\infty}(\bm{Y}_{i})},$
$\mathbf{B}^{(i)}\mathbf{B}^{(i)\top}$ $=\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}{\bm{\psi}}^{\ast(i)}_{k}{\bm{\psi}}^{\ast(i)\top}_{k^{\prime}}\dot{{\bm{\psi}}}^{-1\top}_{k^{\prime}}\mathbf{W}^{\top}_{k^{\prime}}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi^{2}_{\infty}(\bm{Y}_{i})}+\frac{I({\mathcal{C}}_{i}=k)I({\mathcal{C}}_{i}=k^{\prime})}{\pi_{k}(G_{k}(\bm{Y}_{i}))\pi_{k^{\prime}}(G_{k^{\prime}}(\bm{Y}_{i}))}\right)$
Therefore,
$\mathbb{E}\left[\mathbf{A}^{(i)}\mathbf{A}^{(i)\top}\right]$ $=\underbrace{\dot{{\bm{\psi}}}_{\infty}^{-1}\mathbb{E}\left[\frac{{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}_{\infty}^{-1\top}}_{\Sigma_{{\bm{\theta}}}},$ (S1.3)
$\mathbb{E}\left[\mathbf{A}^{(i)}\mathbf{B}^{(i)\top}\right]$ $=\sum^{K}_{k=1}\underbrace{\dot{{\bm{\psi}}}_{\infty}^{-1}\mathbb{E}\left[\frac{{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}_{k}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}^{-1\top}_{k}}_{\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}}\mathbf{W}^{\top}_{k},$ (S1.4)
$\mathbb{E}\left[\mathbf{B}^{(i)}\mathbf{B}^{(i)\top}\right]$ $=\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\mathbf{W}_{k}\underbrace{\dot{{\bm{\psi}}}^{-1}_{k}\mathbb{E}\left[\frac{{\bm{\psi}}^{\ast(i)}_{k}{\bm{\psi}}^{\ast(i)\top}_{k^{\prime}}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}^{-1\top}_{k^{\prime}}}_{\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}}\mathbf{W}^{\top}_{k^{\prime}}$ (S1.5)
$+\sum^{K}_{k=1}\mathbf{W}_{k}\underbrace{\dot{{\bm{\psi}}}^{-1}_{k}\mathbb{E}\left[\frac{{\bm{\psi}}^{\ast(i)}_{k}{\bm{\psi}}^{\ast(i)\top}_{k}}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\right]\dot{{\bm{\psi}}}^{-1\top}_{k}}_{\Sigma_{{\bm{\gamma}}_{2,k}}}\mathbf{W}^{\top}_{k}.$ (S1.6)
From the asymptotic linear expansion of $\widehat{{\bm{\theta}}}_{\text{WCC}}$, $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\theta}}}_{2,k}$, we know that $\Sigma_{{\bm{\theta}}}$ is the asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{WCC}}$; $\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$ is the asymptotic covariance between $\widehat{{\bm{\theta}}}_{\text{WCC}}$ and $\widehat{{\bm{\gamma}}}_{1,k}$; $\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$ is the asymptotic covariance between $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{1,k^{\prime}}$; $\Sigma_{{\bm{\gamma}}_{2,k}}$ is the asymptotic variance of $\widehat{{\bm{\theta}}}_{2,k}$. As a results, the asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is:
$\Sigma_{\text{PS-PPI}}=\Sigma_{{\bm{\theta}}}-\sum^{K}_{k=1}\mathbf{W}_{k}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}-\sum^{K}_{k=1}(\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}})^{\top}\mathbf{W}_{k}^{\top}+\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\mathbf{W}_{k}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\mathbf{W}_{k^{\prime}}^{\top}+\sum^{K}_{k=1}\mathbf{W}_{k}\Sigma_{{\bm{\gamma}}_{2,k}}\mathbf{W}^{\top}_{k}.$ (S1.7)
S1.4 Details of estimating $\Sigma_{\text{PS-PPI}}$.
To obtain a consistent estimate of $\Sigma_{\text{PS-PPI}}$, we may want to obtain consistent estimates for both the meat (e.g., $\mathbb{E}\left[\frac{{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}}{\pi_{\infty}(\bm{Y}_{i})}\right]$ for $\Sigma_{\bm{\theta}}$) and the bread (e.g., $\dot{{\bm{\psi}}}_{\infty}^{-1}$ and $\dot{{\bm{\psi}}}_{\infty}^{-1\top}$ for $\Sigma_{\bm{\theta}}$) parts for each of the variance-covariance matrices $\Sigma_{{\bm{\theta}}}$, $\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$, $\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$, and $\Sigma_{{\bm{\gamma}}_{2,k}}$. This can be done once the closed-form expression of ${\bm{\psi}}^{\ast(i)}$, $\dot{{\bm{\psi}}}_{\infty}$, ${\bm{\psi}}^{\ast(i)}_{k}$, and $\dot{{\bm{\psi}}}_{k}$ can be derived. In the following, we provide two examples, linear regression and logistic regression, to illustrate the derivation of the closed-form of $\Sigma_{{\bm{\theta}}}$ from Eq. (S1.3). Similar derivations can be done for $\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$, $\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$, and $\Sigma_{{\bm{\gamma}}_{2,k}}$ when referring to Eq. (S1.4), (S1.5), and (S1.6).
S1.4.1 Linear Regression
The $\bm{\psi}$} of the linear regression for each subject:
${\bm{\psi}}^{(i)}=X_{i}(y_{i}-X_{i}^{\top}{\bm{\theta}}).$
From Eq. (S1.3), the meat of $\Sigma_{{\bm{\theta}}}$ is
$\mathbb{E}\left[\frac{{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}}{\pi_{\infty}(\bm{Y}_{i})}\right]=\mathbb{E}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}^{2}(X_{i},y_{i})}X_{i}(y_{i}-X^{\top}_{i}{\bm{\theta}}^{\ast})^{2}X^{\top}_{i}\right].$
On the other hand, the bread of $\Sigma_{{\bm{\theta}}}$ is:
$\dot{{\bm{\psi}}}_{\infty}=\mathbb{E}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi(X_{i},y_{i})}X_{i}X_{i}^{\top}\right].$
Therefore, we can compute the consistent estimate of $\Sigma_{\bm{\theta}}$ by estimating the bread and the meat using data:
$\widehat{\mathbb{E}}\left[\frac{{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}}{\pi_{\infty}(\bm{Y}_{i})}\right]=\frac{1}{N}\sum^{N}_{i=1}\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}^{2}_{\infty}(X_{i},y_{i})}X_{i}(y_{i}-X_{i}^{\top}\widehat{{\bm{\theta}}}_{\text{WCC}})^{2}X^{\top}_{i}=\frac{1}{N}\sum^{N}_{i=1}\frac{I({\mathcal{C}}_{i}=\infty)\widehat{\epsilon}^{2}_{i}}{\widehat{\pi}^{2}_{\infty}(X_{i},y_{i})}X_{i}X^{\top}_{i},$
$\widehat{\mathbb{E}}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi(X_{i},y_{i})}X_{i}X_{i}^{\top}\right]=\frac{1}{N}\sum^{N}_{i=1}\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}_{\infty}(X_{i},y_{i})}X_{i}X^{\top}_{i},$
where $\widehat{\epsilon}_{i}=y_{i}-X_{i}^{\top}\widehat{{\bm{\theta}}}_{\text{WCC}}$. For the corresponding meat and bread parts in all the other variance-covariance matrices, we could apply a similar procedure to obtain their consistent estimates.
S1.4.2 Logistic Regression
In the logistic regression, the procedure for obtaining the consistent estimates of the variance-covariance matrices are similar; the only difference is the closed-form expression of ${\bm{\psi}}^{\ast(i)}$, $\dot{{\bm{\psi}}}_{\infty}$, ${\bm{\psi}}^{\ast(i)}_{k}$, and $\dot{{\bm{\psi}}}_{k}$. Specifically, under the logistic regression setting, the log-likelihood for each subject is:
$\ell_{i}({\bm{\theta}})=y_{i}X_{i}^{\top}{\bm{\theta}}-\log\left(1+e^{X_{i}^{\top}{\bm{\theta}}}\right).$
Therefore ${\bm{\psi}}^{(i)}$ is in the form of its first-order derivative:
${\bm{\psi}}^{(i)}=X_{i}\left(y_{i}-\frac{e^{X_{i}^{\top}{\bm{\theta}}}}{1+e^{X_{i}^{\top}{\bm{\theta}}}}\right).$
As such, the meat is:
$\widehat{\mathbb{E}}\left[\frac{{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}}{\pi_{\infty}(\bm{Y}_{i})}\right]=\frac{1}{N}\sum^{N}_{i=1}\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}^{2}_{\infty}(X_{i},y_{i})}X_{i}\left(y_{i}-\frac{e^{X_{i}^{\top}{\bm{\theta}}}}{1+e^{X_{i}^{\top}{\bm{\theta}}}}\right)^{2}X^{\top}_{i}.$
Similarly, since $\dot{{\bm{\psi}}}_{\infty}=\mathbb{E}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi(X_{i},y_{i})}X_{i}X^{\top}_{i}\mu_{i}(1-\mu_{i})\right]$, where $\mu_{i}=\frac{e^{X_{i}^{\top}{\bm{\theta}}}}{1+e^{X_{i}^{\top}{\bm{\theta}}}}$, the bread part is:
$\widehat{\mathbb{E}}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi(X_{i},y_{i})}X_{i}X^{\top}_{i}\mu_{i}(1-\mu_{i})\right]=\frac{1}{N}\sum^{N}_{i=1}\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}_{\infty}(X_{i},y_{i})}X_{i}X^{\top}_{i}\mu_{i}(1-\mu_{i}).$
For the corresponding meat and bread parts in all the other variance-covariance matrices, we could apply a similar procedure to obtain their consistent estimates.
S1.5 Comment on the PS-PPI estimator as an AIPW estimator
In this subsection, we make a comment on why $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is an AIPW estimator. Specifically, when the propensity score model is known, the $i$-th influence function of the AIPW estimators, as defined in tsiatis2006semiparametric, is:
$\varphi_{\text{Tsiatis},i}$ $=-\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}}{\partial{\bm{\theta}}}\Bigr|_{{\bm{\theta}}={\bm{\theta}}^{\ast}}\right]^{-1}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}}^{\ast})+\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}\left(\sum^{K}_{k=1}\pi_{k}(G_{k}(\bm{Y}_{i}))L_{2k}(G_{k}(\bm{Y}_{i}))\right)\right$.
$\;\;\;\left.-\sum^{K}_{k=1}I({\mathcal{C}}=k)L_{2k}(G_{k}(\bm{Y}_{i}))\right]$
$=-\dot{{\bm{\psi}}}_{\infty}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}^{\ast(i)}-\sum^{K}_{k=1}\left(\frac{I({\mathcal{C}}_{i}=\infty)\pi_{k}(G_{k}(\bm{Y}_{i}))}{\pi_{\infty}(\bm{Y}_{i})}-I({\mathcal{C}}=k)\right)\dot{{\bm{\psi}}}_{\infty}^{-1}L_{2k}(G_{k}(\bm{Y}_{i})).$
Note that $\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}}{\partial{\bm{\theta}}}\Bigr|_{{\bm{\theta}}={\bm{\theta}}^{\ast}}\right]^{-1}=-\dot{{\bm{\psi}}}_{\infty}^{-1}$ holds when the propensity score is known:
$\dot{{\bm{\psi}}}_{\infty}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{\text{WCC}}}{\partial{\bm{\theta}}}\Bigr|_{{\bm{\theta}}={\bm{\theta}}^{\ast}}\right]=\mathbb{E}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}\frac{\partial{\bm{\psi}}^{(i)}}{\partial{\bm{\theta}}}\Bigr|_{{\bm{\theta}}={\bm{\theta}}^{\ast}}\right]=\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}}{\partial{\bm{\theta}}}\Bigr|_{{\bm{\theta}}={\bm{\theta}}^{\ast}}\right].$
$L_{2k}(G_{k}(\bm{Y}_{i}))$ are a series of arbitrarily chosen functions. By carefully designing $L_{2k}(G_{k}(\bm{Y}_{i}))$ as:
$L_{2k}(G_{k}(\bm{Y}_{i}))=-\frac{1}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\dot{{\bm{\psi}}}_{\infty}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}{\bm{\psi}}^{\ast(i)}_{k}$,
we have $\varphi_{\text{Tsiatis},i}={\bm{\Gamma}}^{(i)}$, which suggests that the $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ estimator is an AIPW estimator when the propensity score model is known. A similar conclusion could be derived when the propensity score model is unknown but correctly specified, as one only needs to further expand $\widehat{{\bm{\omega}}}$ in ${\bm{\Gamma}}^{(i)}$ (and $\varphi_{\text{Tsiatis},i}$) to obtain the $i$-th influence function of the PS-PPI estimator $\varphi_{i}$.
S1.6 Proof of Corollary 3.2
Proof.
If $\mathbf{W}_{k}$ are the same across $k$, then
$\Sigma_{{\bm{\theta}}_{\text{PS-PPI}}}=\Sigma_{{\bm{\theta}}}-\sum^{K}_{k=1}\mathbf{W}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}-\sum^{K}_{k=1}(\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}})^{\top}\mathbf{W}^{\top}+\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\mathbf{W}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\mathbf{W}^{\top}+\sum^{K}_{k=1}\mathbf{W}\Sigma_{{\bm{\gamma}}_{2,k}}\mathbf{W}^{\top}.$
Equating its first-order derivative to $\bm{0}$:
$-\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}-\sum^{K}_{k=1}(\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}})^{\top}+2\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\mathbf{W}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}+2\sum^{K}_{k=1}\mathbf{W}\Sigma_{{\bm{\gamma}}_{2,k}}=\bm{0}.$
Solving the equation, we know that
$\mathbf{W}^{\ast}=\left(\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}\right)\times\left[\left(\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\right)+\left(\sum^{K}_{k=1}\Sigma_{{\bm{\gamma}}_{2,k}}\right)\right]^{-1}$
leads to the optimal efficiency of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$. Plugging in $\mathbf{W}^{\ast}$ into Equation (7), we have
$\Sigma_{\text{PS-PPI}}=\Sigma_{{\bm{\theta}}}-\left(\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}\right)\times\left[\left(\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\right)+\left(\sum^{K}_{k=1}\Sigma_{{\bm{\gamma}}_{2,k}}\right)\right]^{-1}\times\left(\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}\right)^{\top}.$
Furthermore, we know that
$\left(\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\right)+\left(\sum^{K}_{k=1}\Sigma_{{\bm{\gamma}}_{2,k}}\right)=\textrm{Var}\left(\sum^{K}_{k=1}\widehat{{\bm{\gamma}}}_{1,k}\right)+\sum^{K}_{k=1}\textrm{Var}(\widehat{{\bm{\gamma}}}_{2,k}).$
This suggests that $\left(\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\right)+\left(\sum^{K}_{k=1}\Sigma_{{\bm{\gamma}}_{2,k}}\right)$ is a positive definite matrix, therefore its inverse is a positive definite matrix. Therefore,
$\left(\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}\right)\times\left[\left(\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\right)+\left(\sum^{K}_{k=1}\Sigma_{{\bm{\gamma}}_{2,k}}\right)\right]^{-1}\times\left(\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}\right)^{\top}$
is a positive definite matrix. As a result, we have $\Sigma_{\text{PS-PPI}}\preceq\Sigma_{{\bm{\theta}}}$.
Now we prove the if-and-only-if condition. If $\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}=\bm{0}$, then obviously $\Sigma_{\text{PS-PPI}}=\Sigma_{{\bm{\theta}}}$. If $\Sigma_{\text{PS-PPI}}=\Sigma_{{\bm{\theta}}}$, since $\left[\left(\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}\right)+\left(\sum^{K}_{k=1}\Sigma_{{\bm{\gamma}}_{2,k}}\right)\right]^{-1}$ is a positive definite matrix, it cannot be $\bm{0}$. Hence, $\sum^{K}_{k=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}=\bm{0}$. ∎
S1.7 Proof of Proposition 2
Proof.
From gronsbell2024another, we know that
$\textbf{IF}_{i,\text{Chen}}=-\dot{{\bm{\psi}}}_{\infty}^{-1}\left\{{\bm{\psi}}^{\ast(i)}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}+\text{Cov}({\bm{\psi}}^{\ast(i)},{\bm{\psi}}^{\ast(i)}_{1})\text{Var}({\bm{\psi}}^{\ast(i)}_{1})^{-1}{\bm{\psi}}^{\ast(i)}_{1}\frac{\pi_{\infty}-I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}\right\}.$
Note that $\dot{{\bm{\psi}}}_{\infty}^{-1}$ introduced in the above equation is equivalent to the $\mathbf{A}=\mathbb{E}\left[\mathbf{X}\mathbf{X}^{\top}\right]^{-1}$ introduced in gronsbell2024another in the linear regression case they considered. On the other hand, recall that when $\pi(\cdot)$ is known, the $i$-th influence function of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is:
$\textbf{IF}_{i,\text{PS-PPI}}=-\dot{{\bm{\psi}}}_{\infty}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}^{\ast(i)}+\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}_{\infty}(\bm{Y}_{i})}-\frac{I({\mathcal{C}}_{i}=k)}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\right){\bm{\psi}}^{\ast(i)}_{k}$.
Under the missing outcome setting and the MCAR assumption,
$\textbf{IF}_{i,\text{PS-PPI}}=-\dot{{\bm{\psi}}}_{\infty}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}{\bm{\psi}}^{\ast(i)}+\mathbf{W}_{1}\dot{{\bm{\psi}}}^{-1}_{1}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}-\frac{I({\mathcal{C}}_{i}=1)}{\pi_{1}}\right){\bm{\psi}}^{\ast(i)}_{1}$.
From Corollary 3.2, we know that the optimal weight matrix:
$\mathbf{W}^{\ast}=\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,1}}\left(\Sigma_{{\bm{\gamma}}_{1,1}}+\Sigma_{{\bm{\gamma}}_{2,1}}\right)^{-1},$
where
$\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,1}}=\dot{{\bm{\psi}}}_{\infty}^{-1}\mathbb{E}\left[\frac{{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}_{1}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}^{-1\top}_{1}=\dot{{\bm{\psi}}}_{\infty}^{-1}\text{Cov}({\bm{\psi}}^{\ast(i)},{\bm{\psi}}^{\ast(i)}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\pi^{-1}_{\infty},$
$\Sigma_{{\bm{\gamma}}_{1,1}}=\dot{{\bm{\psi}}}^{-1}_{1}\mathbb{E}\left[\frac{{\bm{\psi}}^{\ast(i)}_{1}{\bm{\psi}}^{\ast(i)\top}_{1}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}^{-1\top}_{1}=\dot{{\bm{\psi}}}^{-1}_{1}\text{Var}({\bm{\psi}}^{\ast(i)}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\pi^{-1}_{\infty},$
$\Sigma_{{\bm{\gamma}}_{2,1}}=\dot{{\bm{\psi}}}^{-1}_{1}\mathbb{E}\left[\frac{{\bm{\psi}}^{\ast(i)}_{1}{\bm{\psi}}^{\ast(i)\top}_{1}}{\pi_{1}(G_{1}(\bm{Y}_{i}))}\right]\dot{{\bm{\psi}}}^{-1\top}_{1}=\dot{{\bm{\psi}}}^{-1}_{1}\text{Var}({\bm{\psi}}^{\ast(i)}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}(1-\pi_{\infty})^{-1}.$
Therefore,
$\mathbf{W}^{\ast}$ $=\left(\dot{{\bm{\psi}}}_{\infty}^{-1}\text{Cov}({\bm{\psi}}^{\ast(i)},{\bm{\psi}}^{\ast(i)}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\pi^{-1}_{\infty}\right)\times\left(\dot{{\bm{\psi}}}^{-1}_{1}\text{Var}({\bm{\psi}}^{\ast(i)}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\right)^{-1}\pi_{\infty}\pi_{1}$
$=\left(\dot{{\bm{\psi}}}_{\infty}^{-1}\text{Cov}({\bm{\psi}}^{\ast(i)},{\bm{\psi}}^{\ast(i)}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\pi^{-1}_{\infty}\right)\times\left(\dot{{\bm{\psi}}}^{\top}_{1}\text{Var}({\bm{\psi}}^{\ast(i)}_{1})^{-1}\dot{{\bm{\psi}}}_{1}\right)\pi_{\infty}\pi_{1}$
$=\dot{{\bm{\psi}}}_{\infty}^{-1}\text{Cov}({\bm{\psi}}^{\ast(i)},{\bm{\psi}}^{\ast(i)}_{1})\text{Var}({\bm{\psi}}^{\ast(i)}_{1})^{-1}\dot{{\bm{\psi}}}_{1}\pi_{1}.$
Plugging $\mathbf{W}^{\ast}$ back to $\textbf{IF}_{i,\text{PS-PPI}}$:
$\textbf{IF}_{i,\text{PS-PPI}}$ $=-\dot{{\bm{\psi}}}_{\infty}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}{\bm{\psi}}^{\ast(i)}$
$+\dot{{\bm{\psi}}}^{-1}\text{Cov}({\bm{\psi}}^{\ast(i)},{\bm{\psi}}^{\ast(i)}_{1})\text{Var}({\bm{\psi}}^{\ast(i)}_{1})^{-1}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}-\frac{I({\mathcal{C}}_{i}=1)}{\pi_{1}}\right){\bm{\psi}}^{\ast(i)}_{1}\pi_{1}$
$=-\dot{{\bm{\psi}}}_{\infty}^{-1}\left\{{\bm{\psi}}^{\ast(i)}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}+\text{Cov}({\bm{\psi}}^{\ast(i)},{\bm{\psi}}^{\ast(i)}_{1})\text{Var}({\bm{\psi}}^{\ast(i)}_{1})^{-1}{\bm{\psi}}^{\ast(i)}_{1}\frac{\pi_{\infty}-I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}\right\}$
$=\textbf{IF}_{i,\text{Chen}}.$
Therefore, the two estimators are asymptotically equivalent. ∎
S1.8 Proof of Proposition 3
Proof.
From kluger2025prediction, we know that the asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{PTD}}$ is
$\Sigma_{\text{PTD}}=\Sigma_{{\bm{\theta}}}-\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1}}\left(\Sigma_{{\bm{\gamma}}_{1}}+\Sigma_{{\bm{\gamma}}_{2}}\right)^{-1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1}}^{\top},$
where $\Sigma_{{\bm{\theta}}}$ is the asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{WCC}}$; $\Sigma_{{\bm{\gamma}}_{1}}$ is the asymptotic variance of $\widehat{{\bm{\gamma}}_{1}}$ (the second subscript is removed as there is only two missingness pattern, and one of them is the fully-observed pattern); $\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1}}$ is the asymptotic covariance between $\widehat{{\bm{\theta}}}_{\text{WCC}}$ and $\widehat{{\bm{\gamma}}_{1}}$; $\Sigma_{{\bm{\gamma}}_{2}}$ is the asymptotic variance of $\widehat{{\bm{\gamma}}_{2}}$.
On the other hand, under the MAR assumption, with only two missingness patterns, Equation (8) becomes:
$\Sigma_{\text{PS-PPI}}=\Sigma_{{\bm{\theta}}}-\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1}}\left(\Sigma_{{\bm{\gamma}}_{1}}+\Sigma_{{\bm{\gamma}}_{2}}\right)^{-1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1}}^{\top}.$
This suggests that both $\widehat{{\bm{\theta}}}_{\text{PTD}}$ and $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ have the same asymptotic variance. As they also have the same asymptotic mean, they are asymptotically equivalent. ∎
Appendix S2 Additional Experimental Results
S2.1 Synthetic data simulation
In this subsection, we present the additional experimental results of the synthetic data simulation under different simulation settings. Specifically, instead of fixing the prediction bias level, we fix the prediction noise level at $\sigma_{\text{pred}}=0$ and vary the bias level and present the results in Figure S1. Based on the simulation settings adopted in Figure 2 and S1, we correspondingly substitute the estimated propensity score with the real propensity score and present the simulation results in Figure S2 and S3. Similarly, we substitute the estimated propensity score with the one obtained from the misspecified model and present the simulation results in Figure S4 and S5.
Figure S1: The same as Figure 2, except that the noise level is fixed at $\sigma_{\text{pred}}=0$, and the X-axis in all panels represents the prediction bias level used in the simulations.
Figure S2: The same as Figure 2, except that the propensity scores are assumed to be known.
Figure S3: The same as Figure 2, except that the propensity scores are assumed to be known, and the noise level is fixed at $\sigma_{\text{pred}}=0$, and X-axis in all panels represents the prediction bias level used in the simulations.
Figure S4: The same as Figure 2, except that the propensity scores are assumed to be estimated from a misspecified model.
Figure S5: The same as Figure 2, except that the propensity scores are assumed to be estimated from a misspecified model, and the noise level is fixed at $\sigma_{\text{pred}}=0$, and X-axis in all panels represents the prediction bias level used in the simulations.